### v3

# Spectral gap of a convex combination of a random permutation and a deterministic matrix

Abstract
We study the spectral gap behavior of an operator obtained by summing a random permutation $M$ and a deterministic bistochastic matrix $Q$. We are interested in the asymptotic in terms of dimension. In the case where $(M,Q)$ are asymptotically free with amalgamation over the diagonal, we can compute limit operators $(u,q)$ which give the weak limit spectral distribution. Therefore we introduce free with amalgamation operators that are suitable for computing the spectral gap limit of our operator in high dimensions. We then approximate the spectral radius of the corresponding limit operator and finally give an upper bound for the spectral radius of the finite-dimensional operator. In particular, we show that if the deterministic matrix underlying graph is an expander, then the underlying graph associated to the sum with a random permutation is again an expander.
1 Introduction
1.1 Notations and conventions
[…]
For a bounded operator $b\in\mathcal{B}\left(\mathrm{H}\right)$ where $\mathrm{H}$ is an Hilbert space, we denote by $\|b\|$ its operator norm associated to the inner scalar product norm of $\mathrm{H}$ and we denote by $b^{\ast}$ the adjoint operator of $b$. For $\epsilon\in\{+,-\}$, we set $b^{\epsilon}=b$ if $\epsilon=+$ and $b^{\epsilon}=b^{\ast}$ if $\epsilon=-$. Regarding the spectrum, for $B$ an operator on a Hilbert space of dimension $N$, we consider the eigenvalues with multiplicities. We denote the $k$-th eigenvalue by $\lambda_{k}(B)$ and order them as follows
$\left|\lambda_{1}(B)\right|\geq\left|\lambda_{2}(B)\right|\geq\cdots\left|\lambda_{N}(B)\right|.$
We call spectral gap of $B$ the following positive function of $B$
$\gamma(B):=\left|\lambda_{1}(B)\right|-\left|\lambda_{2}(B)\right|.$
We introduce the following sparsity parameters of $B\in\mathrm{M}_{N}(\mathbb{C})$
$\|B\|_{0\rightarrow 1}$ $:=\underset{x\in[N]}{\max}|\{y\in[N]\leavevmode\nobreak\ ;\leavevmode\nobreak\ B_{xy}\neq 0\}|$
$\|B\|_{0\rightarrow 1}^{\ast}$ $:=\max\{\|B\|_{0\rightarrow 1},\|B^{\ast}\|_{0\rightarrow 1}\}.$
Besides we introduce the norm
$\|B\|_{\infty\to\infty}^{\ast}:=\max\{\|B\|_{\infty\to\infty},\|B^{\ast}\|_{\infty\to\infty}\}$
where $\|\cdot\|_{\infty\to\infty}$ is the subordinate norm to the infinite norm on $\mathbb{C}^{N}$ i.e.
$\|B\|_{\infty\to\infty}:=\max_{\|X\|_{\infty}=1}\|BX\|_{\infty}=\max_{x}\sum_{y}|B_{xy}|.$
For any integer $L\geq 0$, we introduce the additional sparsity parameter
$\kappa_{L}(B):=\underset{\begin{subarray}{c}k\leq p\leq L,\\ (x_{t},y_{t})\in[N]^{2k}\end{subarray}}{\max}\leavevmode\nobreak\ \left\{\max_{\begin{subarray}{c}n_{1}+\cdots+n_{k}=p\\ n_{t}\leq\ell(B)\end{subarray}}\leavevmode\nobreak\ \prod_{t=1}^{k}(B^{n_{t}})_{x_{t}y_{t}}\right\}^{\frac{1}{p+k}},$ (1)
where $\ell(B):=\frac{\ln(N)}{7\ln(\|B\|_{0\rightarrow 1}^{\ast})}$ and we set
$\kappa(B):=\sup_{L}\kappa_{L}(B).$ (2)
We will omit the $B$ in the notation $\kappa_{L}$ and $\kappa$ if there is no confusion.
1.2 Model
In this article we will consider the following model of matrix $Q\in\mathrm{M}_{N}(\mathbb{C})$.
  1. (H:1)
    For all $x,y\in[N]$, $Q_{xy}\geq 0$ and the vector $\boldsymbol{1}:=(1,\cdots,1)^{\top}\in\mathbb{R}^{N}$ is an eigenvector of $Q$ and $Q^{\ast}$, i.e.
    $Q\boldsymbol{1}=\delta\boldsymbol{1}\quad\text{ and }\quad Q^{\ast}\boldsymbol{1}=\delta\boldsymbol{1}$
    for some $\delta>0$.
  2. (H:2)
    There exists $K>0$ independent of $N$ such that $\|Q\|\leq K$.
  3. (H:3)
    There exists $10<\alpha<$ 1 such that we have
    $d:=\underset{x\in[N]}{\max}|\{y\in[N]\leavevmode\nobreak\ ;\leavevmode\nobreak\ Q_{xy}\neq 0\}|\vee\underset{x\in[N]}{\max}|\{y\in[N]\leavevmode\nobreak\ ;\leavevmode\nobreak\ Q^{\ast}_{xy}\neq 0\}|=e^{o\left(\ln(N)^{\alpha}\right)}.$
We see $Q$ as an adjacency matrix for a $\delta$-regular graph. We then take interest in the matrix
$A=M+Q$ (3)
where $M$ is a uniformly distributed permutation. In the case where $\delta=1$, $Q$ is called doubly stochastic. The set of bistochastic matrices being a convex polytope, any convex combination of a permutation matrix and $Q$ remains bistochastic. In this paper we also are interested in the following matrix
$A_{r}:=(1-r)M+rQ$ (4)
where $M$ is a uniformly distributed permutation and $10<r<$ 1. If we consider $Q$ and $M$ as transition matrices of Markov chains on the same vertex set $[N]$, $P$ is the sum of these two graphs, rescaled to obtain a Markov chain. According to Perron-Frobenius Theorem, $1$ (resp. $2$) is then the largest eigenvalue of $A_{r}$ (resp. of $A$). We are interested in the asymptotic behavior in dimension $N$ of the spectral gap $1-|\lambda_{2}(A_{r})|$ (resp. $2-\left|\lambda_{2}(A)\right|$). The main result of this article is a probabilistic upper bounds on $|\lambda_{2}(A)|$ and $\left|\lambda_{2}(A_{r})\right|$ that depends only on parameters of $Q$. More precisely the bound we are interested in is the spectral radius of an operator that depends on $Q$ and $r$ and is independent with the permutation matrix $M$.
1.3 Free probability and spectral analysis of random matrices
One way of expressing the interest of free probability in terms of spectral analysis for large random matrices is that it is a suitable framework for studying the asymptotic behaviour of non-commutative random variables. Roughly speaking, the notion of freeness in non-commutative probability is the non-commutative equivalent of independence for (classical commutative) random variables. When considering independent non-commutative variables, such as random matrices, this independence is indeed translated into freeness asymptotically. Voiculescu [Voi91] introduced free probability with the idea of constructing spaces of operators imitating basic probability theory and we refer to Anderson, Guionnet and Zeitouni [AGZ10] for a comprehensive study of random matrices.
The asymptotic freeness of a couple $(A,B)$ of random matrices allows the construction of an explicit operator $\mathrm{P}_{\mathrm{free}}\in\mathcal{B}(\mathrm{H})$ such that the spectrum (resp. the empirical spectral measure) of $\mathrm{P}$a polynomial in $(A,B)$, converges asymptotically to the spectrum (resp. spectral measure) of $\mathrm{P}_{\mathrm{free}}$ in a sense yet to be defined. For instance, considering independent uniformly distributed random permutations matrices, the asymptotic operators are the left representations of the free group. For any integer $d$ we then denote $\mathbf{u}=(u(s))_{s\in[d]}$ the generators of $\mathbb{F}_{d}$ the free group with $d$ generators. We keep the same notation $(u(s))_{s\in[d]}$ for the left-representations, i.e. the operators defined by
$u(s):\ell^{2}(\mathbb{F}_{d})$ $\rightarrow\ell^{2}(\mathbb{F}_{d})$
$\delta_{g}$ $\mapsto\delta_{u(s)g}$
where $\delta_{g}:=(\mathds{1}(h=g))_{h\in\mathbb{F}_{d}}\in\ell^{2}(\mathbb{F}_{d})$. The linear form $\tau:=\langle\cdot\delta_{e},\delta_{e}\rangle$ has to be seen as the non-commutative analog of the expectation $\mathbb{E}(\cdot)$ in the classic probability setting. Then two non-commutative operators ${a}_{1},{a_{2}}\in\mathcal{B}\left(\ell^{2}\left(\mathbb{F}_{d}\right)\right)$ such that $\langle{a}_{i}\delta_{e},\delta_{e}\rangle=0$ are called free if for any non-trivial word ${a}$ made from the alphabet $({a}_{1},{a}_{2})$ we have $\langle{a}\delta_{e},\delta_{e}\rangle=0$. This is the analog of defining the classic independence of two complex random variable. Seen as a family of operators, $\mathbf{u}$ is a $d$-Haar free family. Then the fact that a family of possibly random unit matrices $\mathbf{M}=(M(s))_{s\in[d]}$ are asymptotically free means that for any non-commutative polynomial $P(\cdot)\in\operatorname{\mathbb{C}}\langle X(s),X(s)^{\ast}|s\in[d]\rangle$ we have:
$\lim_{N\rightarrow\infty}\operatorname{Tr}_{N}[P(\mathbf{M})]=\langle P(\mathbf{u})\delta_{e},\delta_{e}\rangle$ (5)
where $e\in\mathbb{F}_{d}$ is the neutral element. In general, for any family of possibly random matrices $\mathbf{M}$ and any family of operators $\mathbf{u}$, if the convergence (5) occurs for all polynomial, we say that the family $\mathbf{M}$ converges in distribution to the family $\mathbf{u}$. If such a convergence occurs it already gives spectral information over some operators obtained with the family $\mathbf{M}$. If one considers $P(\cdot)$ a symmetric polynomial, $\mathrm{P}=P(\mathbf{M})$ and $\mathrm{b}=P(\mathbf{u})$ then the previous convergence implies:
$\mu_{\mathrm{P}}:=\frac{1}{N}\sum_{k=1}^{N}\delta_{\lambda_{k}(\mathrm{P})}\underset{N\rightarrow\infty}{\rightharpoonup}\mu_{\mathrm{b}}$ (6)
where $\mu_{\mathrm{b}}$ is the spectral measure given by the spectral theorem applied to $\mathrm{b}$ and the convergence is weak. When the operator $\mathrm{b}$ is bounded, its spectral measure is characterized by its moments. If we furthermore assume freeness of the family $\mathbf{u}=(u(s))_{s\in[d]}$ then the moments, and therefore the spectral measure, of the operator $\mathrm{b}$ are computable knowing the spectral measure of $u(s)$ for all $s\in[d]$. For the simple case of the sum of two free non-commutative operators, the operation on measures is called free convolution and one can refer to [AGZ10, Section 5.3.3]. To see the interest of asymptotic freeness in our setting, recall that we take interest in a random matrix, given by either (3) or (4), of the form
$A=U+Q$
where $Q$ is deterministic and $U$ random. We then have the following theorem.
Theorem 1.1 (Voiculescu, 1991).
Let $\mathrm{G}_{N}$ be either the orthogonal group $\mathrm{O}_{N}$ or the unitary group $\mathrm{U}_{N}$. If $Q\in\mathrm{M}_{N}(\operatorname{\mathbb{C}})$ converges in distribution and is invariant under $\mathrm{G}_{N}$ conjugation and $U\in\mathrm{G}_{N}$ is Haar distributed, then the couple $(Q,U)$ is asymptotically free.
For instance it means that if $Q$ is in the setting of the previous theorem and if one considers the operator $P=\frac{1}{4}\left(Q+Q^{\ast}+U+U^{\ast}\right)$, then the asymptotic spectral measure is completely characterized by the asymptotic distribution of $Q$. The fact that we softened the regularity of the distribution for the random matrix $U$ by taking a random permutation instead of a Haar distributed unitary, weakens the freeness of the asymptotic operators. This new freeness is called freeness with amalgamation, introduced in Section 2. This new freeness has a particular interest in random graph theory because the previous theorem holds again replacing freeness with freeness with amalgamation and invariance under $\mathrm{G}_{N}$ conjugation by invariance under $\mathrm{S}_{N}$ conjugation.
Theorem 1.2.
[ACD+18, Theorem 1.3] If $Q\in\mathrm{M}_{N}(\operatorname{\mathbb{C}})$ converges in distribution and is invariant under $\mathrm{S}_{N}$ conjugation and $M\in\mathrm{S}_{N}$ is uniformly distributed, then the couple $(Q,M)$ is asymptotically free with amalgamation.
Once again the convergence in distribution and the freeness with amalgamation of the limit operators allows to compute the distribution of the limit family. An explicit algorithm to compute the joint distribution of free with amalgamation operators is given by Belinschi, Mai, Speicher [BMS17]. The previous setting is better suited for graph theory applications, since we can compute random graphs which joint entry wise law is not invariant under $\mathrm{O}_{N}$ or $\mathrm{U}_{N}$ conjugation but under $\mathrm{S}_{N}$. Indeed one can consider $Q$ the adjacency matrix (resp. the Markov operator of the simple walk) of an Erdős-Rényi graph, then the couple $(Q,M)$ is asymptotically free with amalgamation and the asymptotic spectral distribution of the operator defined by (3) (resp. by (4)) can be computed. The previous theorem is exactly the reason why we will, in Section 4, construct the explicit free with amalgamation operators $(q,u)$, so that when $Q$ is invariant with conjugation under $\mathrm{S}_{N}$, then 2.2 also says that we have no outliers, that is to say eigenvalues outside of the spectral support of $(1-r)u+rq$ or $u+q$.
1.4 Application to expansion of graphs and stronger convergences
The density of edges in relation to the number of vertices of a graph is called the sparsity of the graph. The idea of graph expansion is the fact that we want to construct growing networks in terms of vertices, control the sparsity and maintain a consequent connectivity. We call the study of the asymptotic behavior of the sparsity/connectivity ratio graph expansion.
A more local approach to the same question is to ask whether, for a given graph, there is a not too small subset of the vertex set that is poorly connected to the rest of the vertex set. In computer science this phenomenon is called a bottleneck. We introduce an important tool that quantifies the previous questions, the Cheeger constant of $G=(V,E)$ a finite connected undirected graph. It is defined by the following:
[…]
where for $F\subset V$, $\partial F\subset E$ is the set of edges connecting a vertex of $F$ to a vertex of $V\backslash F$. The Cheeger constant of a graph is a common quantifier of graph expansion. For $G=(V,E)$ a graph and $\epsilon>0$ we say that $G$ is a $\epsilon$-expander if we have
$h(G)\geq\epsilon.$
It is the object considered as respectful of the condition of having good connectivity while being sparse. The graph expansion is also of interest when considering the simple walk on the graph. Recall that for any undirected finite graph $G=(V,E)$ one can define its adjacency operator $\mathrm{A}=\left(\mathds{1}\left(\{x,y\}\in E\right)\right)_{x,y\in V}\in\mathrm{M}_{\left|V\right|}(\operatorname{\mathbb{C}})$ but also the Markov operator $\mathrm{P}\in\mathrm{M}_{\left|V\right|}(\operatorname{\mathbb{C}})$ associated to the simple walk on $G$ defined by
$\mathrm{P}_{xy}:=\frac{\mathds{1}\left(\{x,y\}\in E\right)}{\sum_{z\in V}\mathds{1}\left(\{x,z\}\in E\right)}.$
The spectral gap measures the asymptotic mixing rate to equilibrium. More precisely, for $\mathrm{P}$ any transition matrix to an aperiodic and irreducible Markov chain on the set $[N]$, for any probability measure $\pi_{0}$ on $V$ we have
$\lim_{t\to+\infty}\|\pi_{0}\mathrm{P}^{t}-\pi\|_{TV}^{1/t}=|\lambda_{2}(\mathrm{P})|,$
where $\pi$ is the invariant measure of $\mathrm{P}$ and for $\nu$ a measure on $[N]$, $\|\nu\|_{TV}:=\frac{1}{2}\sum_{x}\nu(x)$ denotes the total variation norm. The link between the Cheeger constant and the spectral analysis of either the adjacency operator or the Markov operator for the simple walk is given by the following inequality.
Theorem 1.3 (Cheeger Inequality).
Let $G=(V,E)$ be a connected, $d$-regular undirected graph. We denote by $\mathrm{P}$ the Markov operator of the simple walk on $G$. We have:
$\frac{1-\lambda_{2}(\mathrm{P})}{2}\leq\frac{h(G)}{d}\leq\sqrt{1-\lambda_{2}(\mathrm{P})}.$
Therefore an alternative way to define expansion of graphs is the following.
Definition 1.4 (Spectral-expansion).
A $d$-regular graph $G=(V,E)$ is a $\epsilon$-expander for eigenvalues if:
$\lambda_{2}(\mathrm{P})\leq 1-\epsilon.$
where $\mathrm{P}$ is the Markov operator of the simple walk on $G$.
To apply the previous convergences to expansion of possibly random graphs one can consider the simple walk defined by the following Markov operator:
$\mathrm{P}=\frac{1}{2d}\sum_{s=1}^{d}M(s)+M(s)^{\ast}$ (8)
where $\mathbf{M}=(M(s))_{s\in[d]}$ is a family of permutations or tensor products of permutation. The asymptotic operator $\mathrm{b}=\frac{1}{2d}\sum_{s}u(s)+u(s)^{\ast}$ is then the adjacency operator of $\mathbb{F}_{d}$ Cayley graph. The corresponding spectral measure is given by Kesten-McKay law $\mu_{\mathrm{b}}(\mathrm{d}x)=f(\cdot)\mathrm{d}x$ where the density is given by:
$f:\mathbb{R}$ $\rightarrow\operatorname{\mathbb{R}}$
$x$ $\mapsto\frac{d}{2\pi}\frac{\sqrt{4(d-1)-x^{2}}}{d^{2}-x^{2}}\mathds{1}(|x|\leq 2\sqrt{d-1}).$
Having asymptotic freeness for the family $\mathbf{M}$, i.e. the convergence given in (5) and therefore the convergence (6), also means:
$|\sigma(\mathrm{P})\cap[-2\frac{\sqrt{d-1}}{d},2\frac{\sqrt{d-1}}{d}]^{\mathrm{c}}|=o(N).$
In particular we have at most $o(N)$ outliers, i.e. eigenvalues greater in modulus than the Alon-Boppana bound $2\sqrt{d-1}/d$. Also we have as expected that the Alon-Boppana bound is sharp since the weak convergence gives also that for all $\epsilon>0$ we have $|\sigma(\mathrm{P})\cap[2\sqrt{d-1}/d-\epsilon,2\sqrt{d-1}/d]|>\delta$ for some $\delta>0$ and $N$ large enough. The asymptotic freeness (5) had been proved for $\mathbf{M}$ a family of Haar distributed unitaries of $\mathrm{U}_{N}$ by Voiculescu [Voi91] in 1991 and for random permutations by Nica [Nic93] in 1993. In the case where we consider $Q$ a deterministic hermitian matrix, invariant under $\mathrm{S}_{N}$ conjugation, for instance the adjacency operator of an Erdős-Rényi, and $M$ a uniformly radom permutation, then considering
$\mathrm{P}:=Q+M+M^{\ast}$
one can apply 1.2 and, denoting $(q,u)$ the limit free with amalgamation operators for the joint distribution of $(Q,M)$ we have again
$|\sigma(\mathrm{P})\cap[-\|b\|,\|b\|]^{\mathrm{c}}|=o(N)$
where $b=q+u+u^{\ast}$.
In any case, the convergence given in (5) is not sufficient to conclude about the expansion of operators. In fact, it is precisely the absence of outliers that is needed to satisfy the expander 1.4. One way to obtain these expansions is to show strong asymptotic freeness. Again sticking to the unitary case, we will say that a family of unitary matrices $\mathbf{M}=(M(s))_{s\in[d]}$ is strongly asymptotically free if for any non-commutative polynomial $P(\cdot)\in\operatorname{\mathbb{C}}\langle X(s),X(s)^{\ast}|s\in[d]\rangle$ we have (5) and furthermore
$\lim_{N\rightarrow\infty}\|P(\mathbf{M})\|=\|P(\mathbf{u})\|,$ (9)
where the norms above are the operator norms associated to the inner-product of the Hermitian spaces $\operatorname{\mathbb{C}}^{N}$ and $\ell^{2}(\mathbb{F}_{d})$. Bordenave and Collins [BC19] showed the following theorem.
Theorem 1.5 (Bordenave, Collins 2018).
Let $d\geq 1$ be an integer and $\mathbf{M}=(M(s))_{s\in[d]}\in\mathrm{S}_{N}$ iid uniformly distributed random permutations. The family $\mathbf{M}$ is strongly asymptotically free in probability.
The previous theorem says that the random $d$-regular graph which simple walk operator is given by (8) is an expander and optimal. Indeed for any undirected $d$-regular graph there exists a constant $c_{d}$ such that we have the following Alon-Boppana bound for its simple walk operator $\mathrm{P}$
$\lambda_{2}(\mathrm{P})\geq\frac{2\sqrt{d-1}}{d}-\frac{c_{d}}{\ln(N)^{2}}=\|b\|-\frac{c_{d}}{\ln(N)^{2}},$
see [Moh10]. If now $\mathrm{P}$ is taken as in (8) with random permutations, the previous theorem gives
$\mathbb{P}(\lambda_{2}(\mathrm{P})\geq\frac{2\sqrt{d-1}}{d}+\epsilon)\underset{N\rightarrow\infty}{\rightarrow}0$
for all $\epsilon>0$. The previous convergence is exactly what we want to show for the operator $P$ given by (4).
Remark 1.1.
In the general setting we consider in this paper, that is to say $Q$ doubly stochastic, non-Hermitian and $P$ defined by (4) with $M$ a random permutation, 1.2 does not imply the convergence (6). Indeed we still have asymptotic freeness with amalgamation, but the fact that we are dealing with non-Hermitian operators invalidates the spectral theorem for the operator $p=(1-r)u+rq$. One can then consider the Brown measure of the operator $p$, and ask if Girko’s method applies to show the convergence (6) (see for instance [BC12]). This convergence would then depends the distribution chosen for $Q$.
Finally, an interest of looking directly at the operator $A=Q+M$ is that its structure and the computations of its moments are closer to the more general case where we study $A^{\prime}=a_{0}\otimes\operatorname{Id}_{N}+a_{1}\otimes Q+a_{2}\otimes M$ with $a_{i}$ fixed matrices of dimension $k$. Indeed, in many cases the strong convergence in the free case is showed using a linearization trick saying that if the spectrum of $A^{\prime}$ for any choices of $(a_{i})_{i}$ such that $A^{\prime}$ is Hermitian converges then we also have convergence of the spectrum of $P(M,Q)$ for any symmetric polynomial $P(\cdot)$ (see for instance [BC19, Section 6]). To adapt this linearization trick to our case is the fact that the matrix $a_{1}\otimes Q$ does not verify (H:1).
2 Freeness with amalgamation and main result
2.1 Definitions
We consider $Q\in\mathrm{M}_{N}(\mathbb{C})$ verifying (H:1), (H:2) and (H:3) and $M$ as before a uniformly distributed random permutation. We search for deterministic operators, denoted $(u,q)$, such that an approximation of the spectral radius of the deterministic operators
$a=u+q\leavevmode\nobreak\ \leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ \leavevmode\nobreak\ p=(1-r)u+rq$ (10)
asymptotically bounds the spectral radius of the matrices defined by (3) and (4). As we explained in Section 1.3 we want to construct the couple $(u,q)$ free with amalgamation and such that $q$ has same $\ast$-distribution than $Q$ (see 2.1 below). These notions can be found in the article by Au, Cébron, Dahlqvist, Gabriel and Male or [ACD+18] or in Voiculescu, Dykema, and Nica [VDN92].
Definition 2.1.
We give first the setting and definition of freeness with amalgamation.
  • An Operator valued probability space denoted $(\mathcal{A},\mathcal{B},\mathrm{E},\ast)$ consists in a unital algebra $\mathcal{A}$, an involution $\ast:\mathcal{A}\rightarrow\mathcal{A}$, a unital sub-algebra $\mathcal{B}\subset\mathcal{A}$ stable under $\ast$, and a conditional expectation $\mathrm{E}:\mathcal{A}\to\mathcal{B}$ which is a unit-preserving linear map such that $\mathrm{E}(b_{1}ab_{2})=b_{1}\mathrm{E}(a)b_{2}$ for any $a\in\mathcal{A}$ and $b_{1},b_{2}\in\mathcal{B}$.
  • We write $\mathcal{B}\langle X,X^{\ast}\rangle$ for the algebra generated by $\mathcal{B}$ and the freely non-commuting indeterminates $(X,X^{\ast})$. For a monomial $P(X,X^{\ast})=b_{0}X^{\epsilon_{1}}b_{1}X^{\epsilon_{2}}\cdots X^{\epsilon_{n}}b_{n}\in\mathcal{B}\langle X,X^{\ast}\rangle$ the degree of $P(X,X^{\ast})$ is $n$, its coefficients are $b_{0},...,b_{n}\in\mathcal{B}$ and for all $1\leq k\leq n$ we have $\epsilon_{k}\in\{+,-\}$. Then the operator valued distribution, or $\mathrm{E}$-distribution, of $a\in\mathcal{A}$ is the map given by:
    $\mathrm{E}_{a}\leavevmode\nobreak\ :\leavevmode\nobreak\ P(\cdot)\in\mathcal{B}\langle X,X^{\ast}\rangle\mapsto\mathrm{E}[P(a,a^{\ast})]\in\mathcal{B}.$
  • Two elements $a_{1},a_{2}\in\mathcal{A}^{2}$ are called free with amalgamation over $\mathcal{B}$ whenever for any $n\geq 2$, any $\ell_{1}\neq\ell_{2}\neq\cdots\neq\ell_{n}\in\{1,2\}$ and any monomials $P_{1},...,P_{n}\in\mathcal{B}\langle X,X^{\ast}\rangle$, one has
    $\mathrm{E}\left[\left(P_{1}(a_{\ell_{1}},a_{\ell_{1}}^{\ast})-\mathrm{E}\left[P_{1}(a_{\ell_{1}},a_{\ell_{1}}^{\ast})\right]\right)\cdots\left(P_{n}(a_{\ell_{n}},a_{\ell_{n}}^{\ast})-\mathrm{E}\left[P_{n}(a_{\ell_{n}},a_{\ell_{n}}^{\ast})\right]\right)\right]=0.$
We denote $(\mathcal{A},\mathcal{B},\mathrm{E},\Phi,\ast)$ when the following is verified.
  • The $4$-tuple $(\mathcal{A},\mathcal{B},\mathrm{E},\ast)$ is an operator-valued probability space.
  • The triple $(\mathcal{A},\Phi,\ast)$ is a non-commutative probability space with a faithfull trace, i.e. we have $\Phi:\mathcal{A}\rightarrow\operatorname{\mathbb{C}}$ linear, verifying $\Phi(1)=1$ and
    $\Phi(aa^{\ast})=0\Rightarrow a=0.$
  • Finally we have $\Phi\circ\mathrm{E}=\Phi$.
The triplet $(\mathcal{A},\Phi,\ast)$ is called a non-commutative probability space and the $\ast$-distribution of $a\in\mathcal{A}$ is the map
$\Phi_{a}:P(\cdot)\in\operatorname{\mathbb{C}}\langle X,X^{\ast}\rangle\mapsto\Phi[P(a,a^{\ast})]\in\operatorname{\mathbb{C}}.$
Note that $(\mathrm{M}_{N}(\mathbb{C}),\mathrm{D}_{N}(\mathbb{C}),\Delta,\operatorname{Tr}_{N},\ast)$ and $(\mathrm{M}_{N}(L^{\infty}),\mathrm{D}_{N}(L^{\infty}),\Delta,\mathbb{E}\circ\operatorname{Tr}_{N},\ast)$ are operator-valued probability spaces. With no loss of generality here, any operator-valued probability space can be viewed as a subalgebra of a $\mathcal{B}(\mathrm{H})$, set of bounded operators on a Hilbert space $\mathrm{H}$. The main result of the present article consists in upper bounding $\rho(A_{\#}|_{\boldsymbol{1}^{\perp}})$ for #$\#\in\{\emptyset,r\}$ defined by (3) and (4) by an approximation of the spectral radius of respectively $a$ and $p\in(\mathcal{A},\mathcal{B},\mathrm{E},\Phi)$, defined by (10) and where $q$ has same $\ast$-distribution as $Q$ and $(q,u)$ are free with amalgamation over $\mathcal{B}$. More precisely, owing to Gelfand’s Theorem we have for such $\mathrm{b}\in\mathcal{B}(\mathrm{H})$
$\rho(\mathrm{b})=\lim_{n\to\infty}\|\mathrm{b}^{n}\|^{1/n}.$
Therefore the approximations of spectral radius we will use are
$\mathrm{N}_{n}(Q)=\|a^{n}\|^{1/n}\leavevmode\nobreak\ \leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ \leavevmode\nobreak\ \mathrm{N}_{n,r}(Q)=\|p^{n}\|^{1/n}$ (11)
for a fixed $n\in\operatorname{\mathbb{N}}$. We will omit the dependence on $Q$ and simply denote by $\mathrm{N}_{n}$ or $\mathrm{N}_{r,n}$ when there is no confusion. The fact that this norm depends only on the matrix $Q$ will be a consequence of the construction of the operators $(q,u)$ in Section 4.
2.2 Main result
We fix $10<r<$ 1 and we set
$\rho_{n}=\max\{\mathrm{N}_{n};2\kappa(Q);\delta\}\leavevmode\nobreak\ \leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ \leavevmode\nobreak\ \rho_{n,r}=\max\{\mathrm{N}_{n,r};2r\kappa(Q);r\delta\}$ (12)
where $\kappa(Q)$ is given by (2) and $\mathrm{N}_{n,\#}$ by (11) for #$\#\in\{\emptyset,r\}$. The main theorem is the following one.
Theorem 2.2.
Let $Q\in\mathrm{M}_{N}(\mathbb{C})$ be a deterministic matrix. There exists $(\mathcal{A},\mathcal{B},\mathrm{E},\Phi)$ an operator-valued probability space and $(q,u)\in\mathcal{A}$ with $u$ Haar unit operator such that
  • $q$ and $Q$ have same $\ast$-distribution,
  • $q$ and $u$ are free with amalgamation over $\mathcal{B}$.
We further assume that $Q$ verifies (H:1), (H:2) and (H:3). Then for any $n$ integer and $10<c<$ 1 there exists $C>0$ such that setting
$\epsilon:=C\frac{\ln(d)}{\ln(N)^{\alpha}},$
we have
$\mathbb{P}(\rho(A_{\#}|_{\boldsymbol{1}^{\perp}})\geq(1+\epsilon)\rho_{n,\#})\leq N^{-c},$
for $M\sim\mathcal{U}(\mathrm{S}_{N})$, #$\#\in\{\emptyset,r\}$ and $N$ large enough.
Acknowledgments
I would like to thank Charles Bordenave for the numerous discussions that led to the main ideas of this article and for his constant supervision of the following work. I would also like to thank Camille Male for his interest in my work and for answering my questions about his work, in particular about the freeness of amalgamation. I thank Guillaume Aubrun for his supervision and feedback, even though the pandemic context made it logistically difficult. Finally, I would like to thank Adam Arras, Alexandre Gaudillière, Nordine Moumeni, and in general all the people who, besides my supervisor, discussed my work or theirs with me and enriched my knowledge, understanding, or confidence in the fields of random matrices and graph theory.
3 Some applications
In this section we give examples where 2.2 holds, and where the upper bound $\rho_{n,\#}$ defined by (12) is equal to the approximation of the spectral radius of the deterministic operator. We also give an example where an improvement of $\rho_{n}$ would be appreciated to conclude.
3.1 Convex combination
In the setting of 2.2 and for any $Q\in\mathrm{M}_{N}(\mathbb{C})$ that verifies (H:1), it is sufficient to consider $10<r<$ 1 small enough to have $\rho_{n,r}=\|[(1-r)u+rq]^{n}\|^{1/n}$ where $\rho_{n}$ is defined by (12). Indeed we have then $Q_{xy}\leq\delta$ for all $x,y\in[N]$ and therefore for any integers $1\leq k\leq p$ and $(n_{t})_{t\leq k}$ such that $\sum_{t=1}^{k}n_{t}=p$ and for any $(x_{t},y_{t})_{t}\in[N]^{2k}$ we have
$\prod_{t=1}^{k}(Q^{n_{t}})_{x_{t}y_{t}}\leq\delta^{p},$ (13)
which implies $\kappa(Q)\leq(1\vee\delta)$ where $\kappa(\cdot)$ is defined by (2). Furthermore owing to 4.7, we have that $\|[(1-r)u+rq]^{n}\|^{1/n}\geq(1-r)$. We deduce that for all $0<r<\frac{1}{3(\delta\vee 1)}$ we have $\rho_{n}=\|[(1-r)u+rq]^{n}\|^{1/n}$ as desired. In particular for $Q$ bistochastic the upper bound $\rho_{n}$ is the approximation $\rho_{n}=\|[(1-r)u+rq]^{n}\|^{1/n}$ for any $n\geq 0$ and for any $0<r<\frac{1}{3}$. To then show that the spectral gap is strictly positive, one would then need to provide an upper bound for the norm $\|[(1-r)u+rq]^{n}\|^{1/n}$ that is strictly less than $1$. In the following section we give an explicit example where we can give such an upper bound.
3.2 The asymptotically free case
We consider $Q$ the Markov operator of a simple walk obtain with a perfect matching. One way to compute $Q$ is to take any $S\in\mathrm{S}_{N}$ permutation matrix and consider $Q=SUS^{\ast}$ where $U=\operatorname{Diag}\left(\begin{pmatrix}0&1\\ 1&0\end{pmatrix}\cdots\begin{pmatrix}0&1\\ 1&0\end{pmatrix}\right)$. The spectral gap for the simple walk defined by $Q$ is trivially equal to 00 since $1$ is an eigenvalue with an eigenspace of dimension at least $2$ for $N\geq 4$. Computing the result of Biane and Lehner [BL99] we obtain that the spectral radius of $a=u+q$ where $q$ has same distribution as $Q$ and $u$ is free with $Q$ is $\rho=\sqrt{3}$ (given by the equation $|\lambda|^{2}+1=|\lambda^{2}-1|^{2}$). For any $\epsilon>0$, owing to Gelfand theorem, there exists $n_{0}>0$ such that for all $n\geq n_{0}$ we have $\mathrm{N}_{n}\leq\left(1+\epsilon\right)\rho$. A realization of the spectrum of $A=M+Q$ encircled by the circle of radius $\rho$ is given in Figure 1.
Figure 1: Plot of the eigenvalues of $A=M+Q$ with $Q$ perfect matching when $N=600$. In red the circle of radius equal to $\rho=\sqrt{3}$ the spectral radius of $u+q$ computed from the formula given by Biane and Lehner in [BL99].
However the 2.2 does not allow to directly conclude about the spectral gap for $A$. Indeed in this case we have $\kappa=1$ and therefore $\rho_{n}=2\kappa=\lambda_{1}(A)$. It shows that an improvement on the factor $2$ of $\kappa$ in the definition of $\rho_{n}$ would be beneficial to our purposes. This factor appears in the proof of 7.5, Equation 33.
However, one can consider again the convex combination case (i.e. the matrix $A_{r}$) one can compute the explicit asymptotical radius applying Biane and Lehner method. We rewrite
$p=(1-r)u+rq=(1-r)\left(u+\frac{r}{1-r}q\right)=(1-r)(u+q^{\prime})$
and the domain of the Brown measure of $u+q^{\prime}$ is contained in the curve of equation
$\frac{1}{\left|\lambda-\varphi\right|^{2}}+\frac{1}{\left|\lambda+\varphi\right|^{2}}=2$
where $\varphi=r/(1-r)$ (see [BL99, Equation (4.5)]). In this case, owing to 4.7 we have $\mathrm{N}_{r,n}\geq(1-r)$ for all $n$ and therefore considering $r\leq 1/3$ we have $\rho_{n,r}=\mathrm{N}_{n,r}$. Besides, for all $\epsilon>0$ we have for $n$ large enough $N_{n,r}\leq\rho(p)+\epsilon$ which proves a strictly positive lower bound for the spectral gap for $N$ large enough.
(a) r=1/3
(b) r=1/4
Figure 2: In blue the eigenvalues of $P=(1-r)M+rQ$ with $Q$ perfect matching and $N=600$. In red the spectral radius of $p=(1-r)u+rq$.
3.3 Application to $d$-regular Ramanujan graphs
Let $G=(V,E)$ be a finite, connected, non-oriented $d$-regular graph with $V=[N]$. Let $B$ be its adjacency matrix and we set $Q:=\frac{1}{d}B$. In this case we have $\delta=1$. For all $n\geq 0$ integer we have
$Q^{n}=\frac{1}{N}1\otimes 1+Q^{n}|_{\boldsymbol{1}^{\perp}}$
where $\boldsymbol{1}\otimes\boldsymbol{1}$ is the $N\times N$ matrix which entries are all equal to $1$. Setting
$\rho:=\max\{|\mu|;\leavevmode\nobreak\ \mu\in\operatorname{Sp}(Q);\leavevmode\nobreak\ |\mu|\neq 1\}$
we have that for all $x,y\in[N]$
$(Q^{n})_{xy}\leq\frac{1}{N}+\rho^{n}.$ (14)
Besides for all $d$, there exists $c_{d}>0$ such that we have the following Alon-Boppana lower bound
$\rho\geq\frac{2\sqrt{d-1}}{d}-\frac{c_{d}}{\ln(N)^{2}}.$
For the previous bound one can refer to [Moh10]. From the previous bound we have that for all integer $d$ and $10<c<$ 1 we have $\rho\geq c\frac{2\sqrt{d-1}}{d}$ for $N$ large enough and therefore for all $n\leq\frac{\ln(N)}{7\ln(d)}$ we have $\rho^{n}\geq N^{-1/38}$. Owing to (14) and (1) we have
$\kappa(Q)\leq\left(1+o_{N}(1)\right)\rho.$
Therefore for any undirected $d$-regular graph verifying that $\rho<\frac{1}{2}$, for $N$ large enough we have $\rho_{n}=\mathrm{N}_{n}$. In particular the spectral radius of either $A$ or $A_{r}$ is upper bounded by the spectral radius of $a$ and $p$ respectively. As an example, we can consider Ramanujan graphs which by definition verify
$\rho\leq\frac{2\sqrt{d-1}}{d}$
and therefore for all $d\geq 15$ we have $\rho<\frac{1}{2}$. An explicit example of such graphs are the Ramanujan graphs $X^{p,q}$ of degree $p+1$ and cardinality given by $\frac{q(q^{2}-1)}{2}$ where $(p,q)$ are prime numbers. Their construction is given in [DSV03, Chapter 4].
4 Intermediate construction
We give here a constructive definition of the operator valued probability space $(\mathcal{A},\mathcal{B},\mathrm{E},\Phi)$ and the operators $(q,u)\in\mathcal{A}$ which verifies the first statement of 2.2, i.e.
  • $q$ has same $\ast$-distribution as $Q$,
  • The operator $u$ is unitary (i.e. $uu^{\ast}=u^{\ast}u=1$) and $q$ and $u$ are free with amalgamation over $\mathcal{B}$.
In a second time, we study the $\ast$-distribution of $a=u+q$ and $p=(1-r)u+rq$ and give an approximation of the spectral radius of $a$ and $p$.
4.1 Construction of $q$ and $u$ as adjacency operators of graphs
The construction given here is inspired by the quasi-tree construction described by Hermon, Sly and Sousi [HSS20]. We will consider $q$ and $u$ as adjacency operators of explicit graphs, meaning that we will give explicit writings of these operators allowing us to then compute $a$ and $p$ moments. One has to consider the construction of these graphs as follows. First, we take infinite and countable copies of the vertex set $V=[N]$ on which the adjacency operator $Q$ is defined. Then we consider each copy as a vertex and we put a tree structure on the set of copies. The tree is regular and the degree depends on the dimension $N$ of the operator $Q$. More precisely, we consider one of the copies as the origin (denoted by $H_{\emptyset}$ in the Figure 3). We consider that the origin has $N$ descendants which are $N$ other copies of the vertex set $[N]$. To each vertex $x$ of the origin we assign uniformly at random one of the $N$ descendant copies. In that descendant, we take uniformly at random one of its vertex $x^{\prime}$ to receive the ascending edge. The edge $(x,x^{\prime})$ is represented in blue in Figure 3 and the adjacency operator that will correspond to the graph defined by the blue edges (after constructing them all) is $u$. We repeat this operation for each descendant, taking at random the vertex set copies among the copies never seen before, and so on. The following Figure 3 shows an example for $N=3$.
Figure 3: Quasi-tree, first step where $N=3$. In blue the first step of construction of the operator $u$.
We give simultaneously the construction $u^{\ast}$, inverse of the unit operator $u$. To each blue edge we connect an inverse red edge, which connects it to its antecedent. We also construct the symmetric graph above the origin, where now the red edges corresponding to $u^{\ast}$ are drawn first and can be seen as descendants, and the blue ones corresponding to $u$ are their inverse (see Figure 4).
Figure 4: Quasi-tree, second step. Construction of $u^{\ast}$ in red as the inverse of $u$ and symmetrizing the construction inverting $u$ and $u^{\ast}$ roles.
Finally it remains to place on each set of vertices the edges corresponding to the graph of $Q$ (in Figure 5 we considered $Q=(1,1,1)^{\top}\cdot(1,1,1)$).
Figure 5: Quasi-tree final step. Adding on each vertex-set of the quasi-tree structure the edges defined by the operator $Q$.
The operators $q$ and $u$ we are looking for are the adjacency operators of the graphs described above. For instance, the operator $a=u+q$ is the adjacency operator of the graph where we considered the blue and green edges of Figure 5.
4.2 Construction of the underlying Hilbert space
To make our Hilbert space explicit, we first consider the following vertex set
$\mathcal{S}:=\{(k_{0},\epsilon_{1},k_{1},\epsilon_{2},k_{2},...,\epsilon_{n-1},k_{n-1},\epsilon_{n})\leavevmode\nobreak\ \leavevmode\nobreak\ ;n\geq 1\leavevmode\nobreak\ \leavevmode\nobreak\ ;\forall t\leq n,\leavevmode\nobreak\ \leavevmode\nobreak\ \epsilon_{t}\in\{+,-\}\text{ and }k_{t}\in[N]\}.$
Let $\{H_{\emptyset}\}\cup\{H_{s}\}_{s\in\mathcal{S}}$ be independent copies of $\mathbb{C}^{N}$. With respect to the previous Section 4.1, $\mathcal{S}$ is the index set for the countable and infinite copies of $[N]$ considered as $Q$ vertex set. One can think of $n$ as the future number of generations from the origin and $\epsilon_{t}$ as either descending from generation $t$ to generation $t+1$ taking $u$ or $u^{\ast}$. For $s:=(k_{0},\epsilon_{1},k_{1},...,k_{n-1},\epsilon_{n})\in\mathcal{S}$, we set the Hilbert space
$\mathrm{H}_{s}:=H_{\emptyset}\otimes H_{(k_{0},\epsilon_{1})}\otimes H_{(k_{0},\epsilon_{1},k_{1},\epsilon_{2})}\otimes\cdots\otimes H_{s},$
and we define
$\mathrm{H}_{0}:=\bigoplus_{s\in\mathcal{S}}\mathrm{H}_{s}.$
We will reduce our Hilbert space by considering only the basis vector $e_{1}\otimes\cdots\otimes e_{n}\in\mathrm{H}_{s}$ such that for all $t\leq n$, $e_{t}$ is the $k_{t}$-th canonical basis vector of $H_{(k_{0},\epsilon_{1},...,k_{t-1},\epsilon_{t})}$. From now on for $(k_{0},\epsilon_{1},k_{1},...,\epsilon_{n-1},k_{n-1},\epsilon_{n})\in\mathcal{S}$, for $t\leq n$ we consider $i_{t}:=(\epsilon_{t},k_{t})$. For any $s=(k_{0},i_{1},...,i_{n},\epsilon_{n})\in\mathcal{S}$, we denote by $e(k_{0})\otimes e(i_{1})\otimes\cdots\otimes e(i_{n})\in\mathrm{H}_{s}$ the tensor product where for all $t\leq n$, $e(i_{t})=e(\epsilon_{t},k_{t})$ is the $k_{t}$-th canonical basis vector of $H_{(k_{0},i_{1},...,i_{t-1},\epsilon_{t})}$. Explicitly we consider $\mathrm{H}_{1}$ the subspace of $\mathrm{H}_{0}$ which is the closure of the span of
$\{e(k_{0})\otimes e(i_{1})\otimes\cdots\otimes e(i_{n})\in\mathrm{H}_{(k_{0},i_{1},...,i_{n-1},\epsilon_{n})}\leavevmode\nobreak\ ;\leavevmode\nobreak\ s=(k_{0},i_{1},...,i_{n},\epsilon_{n})\in\mathcal{S}\}.$
We introduce another vertex set
$\mathcal{V}_{1}:=\{0\}\cup\{(k_{0},i_{1},i_{2},...,i_{n})\leavevmode\nobreak\ ;\leavevmode\nobreak\ n\in\mathbb{N}\leavevmode\nobreak\ ;\leavevmode\nobreak\ \forall t\leq n,\leavevmode\nobreak\ i_{t}:=(\epsilon_{t},k_{t})\in\{+,-\}\times[N]\leavevmode\nobreak\ ;\text{ and }k_{0}\in[N]\}.$
One can consider that $\mathcal{S}$ is a partition of $\mathcal{V}_{1}$, i.e. that for all $v:=(k_{0},i_{1},...,i_{n})\in\mathcal{V}_{1}$, we have $v\in s:=(k_{0},i_{1},i_{2},...,i_{n-1},\epsilon_{n})\in\mathcal{S}$. For $v:=(k_{0},i_{1},...,i_{n})\in\mathcal{V}_{1}$, we also denote:
$e(v)$ $:=e(k_{0})\otimes e(i_{1})\otimes\cdots\otimes e(i_{n})$
$e(0)$ $:=0$
$\epsilon(v)$ $:=\epsilon_{n}$
$v^{-}$ $:=(k_{0},i_{1},...,i_{n-1}).$
In particular, we see that $\mathrm{H}_{1}$ is equal to the closure of ${\operatorname{Span}\{e(v)\leavevmode\nobreak\ ;\leavevmode\nobreak\ v\in\mathcal{V}_{1}\}}$ and this space can be seen as all possible ways to connect any $n$-th generation vertex $v=(k_{0},i_{1},...,i_{n})\in\mathcal{V}_{1}$ to any descendant one $v^{\prime}:=(k_{0},i_{1},...,i_{n},i_{n+1})$ (note that $(v^{\prime})^{-}=v$). So for $e=e(k_{0})\otimes e(i_{1})\otimes\cdots\otimes e(i_{n})\in\mathrm{H}_{1}$ we can also denote:
$v_{e}:=(k_{0},i_{1},...,i_{n}).$
Finally for $v:=(k_{0},i_{1},...,i_{n})\in\mathcal{V}_{1}$ (resp. for $s:=(k_{0},i_{1},...,i_{n-1},\epsilon_{n})\in\mathcal{S}$), we define the length of $v$ (resp. of $s$) by the integer $n$. We then denote it by $|v|=|s|=n$. We now construct the random Hilbert space at stake here. First consider the families of random permutations on $[N]$
$(j_{+}(s))_{s\in\mathcal{S}}\text{ and }(j_{-}(s))_{s\in\mathcal{S}}\text{ iidand }\sim\mathcal{U}(\mathrm{S}_{N}).$
Recall that, with respect to the construction in Section 4.1, the randomness comes from the fact that at each generation, for each vertex $v$ we randomly take a descendant copy of the vertex set $[N]$ and a vertex $v^{\prime}$ in it to which we associate $v$. It will verify $(v^{\prime})^{-}=v$. Note that for $s:=(k_{0},i_{1},...,i_{n-1},\epsilon_{n})\in\mathcal{S}$ and $k\in[N]$, we have that $v:=(s,k)=(k_{0},i_{1},...,\epsilon_{n},k)\in\mathcal{V}_{1}$. So, keeping the same notations, from now on we denote
$j_{\epsilon}(v):=j_{\epsilon}(s)(k)$
where $\epsilon\in\{+,-\}$. We may now define the random vertex set
$\mathcal{V}:=\{v=(k_{0},i_{1},...,i_{n})\in\mathcal{V}_{1}\leavevmode\nobreak\ \leavevmode\nobreak\ ;\leavevmode\nobreak\ \leavevmode\nobreak\ \forall\ell<n:\leavevmode\nobreak\ k_{\ell}\neq j_{-\epsilon_{\ell}}(k_{0},i_{1},...,i_{\ell-1})\text{ or }\epsilon_{\ell+1}=\epsilon_{\ell}\},$
and the random Hilbert subspace of $\mathrm{H}_{1}$ obtained by taking the closure of the vector space spanned by
$\Xi:=\{e=e(k)\otimes e(i_{1})\otimes\cdots\otimes e(i_{n})\leavevmode\nobreak\ \leavevmode\nobreak\ ;\leavevmode\nobreak\ v_{e}\in\mathcal{V}\}.$ (15)
We denote by $\mathrm{H}$ the Hilbert space obtained. We now explicit the Banach and operator-valued probability space structure of $\mathcal{B}(\mathrm{H})\subset\mathcal{B}(L^{2}(\mathrm{H}_{0}))$.
Definition 4.1.
  1. 1.
    The space $(\mathrm{H},(\cdot,\cdot))\subset L^{2}(\mathrm{H}_{0})$ is a Hilbert space where we considered the scalar product defined by
    $\forall h,h^{\prime}\in\mathrm{H},\leavevmode\nobreak\ \leavevmode\nobreak\ (h,h^{\prime}):=\mathbb{E}[\langle h,h^{\prime}\rangle]\leavevmode\nobreak\ \leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ \leavevmode\nobreak\ \|{h}\|_{2}:=\sqrt{(h,h)}.$
[…]
  2. 2.
    $(\mathcal{B}(\mathrm{H}),\Phi,\ast)$ is non-commutative probability space where the trace is given by
    $\Phi(a):=\mathbb{E}\left[\langle ae_{\alpha},e_{\alpha}\rangle\right]$
    for $a\in\mathcal{B}(\mathrm{H})$, where $e_{\alpha}\in H_{\emptyset}$ and $\alpha\sim\mathcal{U}([N])$ is independent from $(j_{\epsilon}(s))_{\epsilon\in\{-,+\},s\in\mathcal{S}}$.
  3. 3.
    The conditional expectation $\mathrm{E}$ on $\mathcal{B}(\mathrm{H})$ is the operator verifying
    $\langle\mathrm{E}(a)e,f\rangle=\mathds{1}(e=f)\langle ae,f\rangle$
    for all $a\in\mathcal{B}(\mathrm{H})$ and for all $e,f\in\Xi$.
We now define the operators $(q,u)$ in 2.2.
Definition 4.2.
We define the set $\mathcal{E}_{+}\subset\mathcal{V}^{2}$ as the set of edges $(v,v^{(+)})\in\mathcal{V}^{2}$ verifying that for all $v=(k_{0},...,i_{n})$, we have $v^{(+)}=(k_{0},i_{1},...,i_{n},(+,j_{+}(v)))$ and $k_{n}\neq j_{-}(k_{0},i_{1},...,i_{n-1})$. The operator $u$ is the adjacency operator of the directed graph $\mathcal{G}_{+}=(\mathcal{V},\mathcal{E}_{+})$. It can also be defined by an explicit formula on the basis vectors spanning $\mathrm{H}$. Let $e:=e(k_{0})\otimes e(i_{1})\otimes\cdots e(i_{n})\in\mathrm{H}$, we have
$u(e)=\left\{\begin{array}[]{ll}e\otimes e(+,j_{+}(x))&\mbox{if }\left\{\begin{array}[]{ll}&\epsilon_{n}=+\\ &\mbox{or}\\ &k_{n}\neq j_{-}(k_{0},i_{1},...,i_{n-1})\end{array}\right.\\ e(k_{0})\otimes e(i_{1})\otimes\cdots\otimes e(i_{n-1})&\mbox{otherwise.}\end{array}\right.$
The operator $u^{\ast}$ is then the adjacency operator of $\mathcal{G}_{-}=(\mathcal{V},\mathcal{E}_{-})$ where, taking the same notations for $v$, we have that $(v,v^{(-)})\in\mathcal{E}_{-}$ if $v^{(-)}=(k_{0},i_{1},...,i_{n},(-,j_{-}(v))$ and $k_{n}\neq j_{-}(k_{0},i_{1},...,i_{n-1})$. Using the same notation for $e\in\mathrm{H}$, $u^{*}$ can be defined symmetrically by the formula
$u^{*}(e)=\left\{\begin{array}[]{ll}e\otimes e(-,j_{-}(x))&\mbox{if }\left\{\begin{array}[]{ll}&\epsilon_{n}=-\\ &\mbox{or}\\ &k_{n}\neq j_{+}(k_{0},i_{1},...,i_{n-1})\end{array}\right.\\ e(k_{0})\otimes e(i_{1})\otimes\cdots\otimes e(i_{n-1})&\mbox{otherwise.}\end{array}\right.$
Finally we consider the isometric injection
$\Pi$ $:\mathrm{M}_{N}(\mathbb{C})\longrightarrow\mathcal{B}(\mathrm{H})$
[…]
and we set $q:=\Pi(Q)$. We now consider the operator valued probability space $(\mathcal{A},\mathcal{B},\mathrm{E},\Phi,\ast)$ where $\mathcal{A}$ is the subalgebra of $\mathcal{B}(\mathrm{H})$ spanned by $u$, $u^{\ast}$, $\mathcal{M}:=\Pi(\mathrm{M}_{N}(\operatorname{\mathbb{C}}))$ and $\mathcal{B}:=\mathrm{E}(\mathcal{A})$. Finally we denote $\mathcal{U}$ the sub algebra of $\mathcal{A}$ spanned by $u$ and $u^{\ast}$.
Remark 4.1.
We could have defined both $u$ and $u^{*}:=u^{-1}$ as follows:
$u^{\epsilon}(e)=\left\{\begin{array}[]{ll}e\otimes e(\epsilon,j_{\epsilon}(x))&\mbox{if }\left\{\begin{array}[]{ll}&\epsilon_{n}=\epsilon\\ &\mbox{or}\\ &k_{n}\neq j_{-\epsilon}(k_{0},i_{1},...,i_{n-1})\end{array}\right.\\ e(k_{0})\otimes e(i_{1})\otimes\cdots\otimes e(i_{n-1})&\mbox{otherwise.}\end{array}\right.$
For all $B\in\mathrm{M}_{N}(\mathbb{C})$, we have
$\Phi(b)$ $=\mathbb{E}[\langle be_{\alpha},e_{\alpha}\rangle]=\mathbb{E}[\langle Be_{\alpha},e_{\alpha}\rangle]=\frac{1}{N}\operatorname{Tr}(B).$
Applying the above to $B=P(Q,Q^{\ast})$ for any $P\in\operatorname{\mathbb{C}}\langle X,X^{\ast}\rangle$ we obtain that $Q$ and $q$ have the same $\ast$-distribution.
4.3 Freeness with amalgamation over $\mathcal{B}$ of $u$ and $\mathcal{M}$
The freeness with amalgamation of $q$ and $u$ is a direct consequence of the freeness with amalgamation of $u$ (i.e. the algebra spanned by $u$ and $u^{\ast}$) and the algebra $\mathcal{M}$. It boils down to checking that for all $n\geq 1$ and $b_{1},...,b_{n}\in\mathcal{M}$ such that $\mathrm{E}(b_{i})=0$, if $\alpha_{1},...,\alpha_{n+1}\in\mathbb{Z}$ are such that for all $i\notin\{1,n+1\}$ we have $\alpha_{i}\neq 0$ then we have
$\mathrm{E}\left[u^{\alpha_{1}}b_{1}\cdots u^{\alpha_{n}}b_{n}u^{\alpha_{n+1}}\right]=0.$
It is equivalent to check that for all $e\in\Xi$ (see (15)) we have
$\langle u^{\alpha_{1}}b_{1}\cdots u^{\alpha_{n}}b_{n}u^{\alpha_{n+1}}e,e\rangle=0.$ (16)
This is the main result of the following lemma.
Lemma 4.3.
Let $N\in\mathbb{N}$:
  1. 1.
    For all $v=(k,i_{1},...,i_{p})\in\mathcal{V}$ we have
    $e(v)=u^{\alpha_{m}}b_{m}u^{\alpha_{m-1}}\cdots u^{\alpha_{1}}b_{1}u^{\alpha_{0}}e(k)$
    for some $\alpha_{0},...,\alpha_{m}\in\mathbb{Z}$ and $b_{1}=\Pi({B}_{1}),...,b_{m}=\Pi({B}_{m})\in\mathcal{M}$ where $B_{i}$ are elementary matrices with zeros on the diagonal.
  2. 2.
    Conversely, for all $k\in[N]$, for all $b_{1}=\Pi({B}_{1}),...,b_{m}=\Pi({B}_{m})\in\mathcal{B}_{1}$ where $B_{i}$ are elementary matrices with zeros on the diagonal and for all $\alpha_{0},...,\alpha_{m}\in\mathbb{Z}$ such that $\alpha_{i}\neq 0$ for $i\notin\{0,m\}$, there exists $v\in\mathcal{V}$ such that
    $u^{\alpha_{m}}b_{m}u^{\alpha_{m-1}}\cdots u^{\alpha_{1}}b_{1}u^{\alpha_{0}}e(k)=\#e(v)\text{ and }\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ |v|=\sum_{i=1}^{m}|\alpha_{i}|$ (17)
    with #$\#\in\{0,1\}$. The vertex $v$ verifies $v=(v^{-},(\epsilon(\alpha_{m}),j_{\epsilon(\alpha_{m})}(v^{-})))$ for some $v^{-}\in\mathcal{V}$.
In particular, $u$ and $\mathcal{M}$ are free with amalgamation over $\mathcal{B}$.
[…]
To show specifically the second part of the previous lemma we will use the following.
Lemma 4.4.
Let $N\in\mathbb{N}$, $v:=(k,i_{1},...,i_{p})\in\mathcal{V}$ and $\alpha\in\mathbb{Z}$. We set $\epsilon:=\operatorname{sign}(\alpha)$ and $i_{p}:=(\epsilon_{p},k_{p})$. We assume that either $p=0$ (i.e. $v=(k)$), or $\epsilon_{p}=\epsilon$ or $k_{p}\neq j_{-\epsilon}(k,i_{1},...,i_{p-1})$. We set the notations
$v_{\epsilon}$ $:=(v,(\epsilon,j_{\epsilon}(v)))\in\mathcal{V}$
[…]
Then $|v_{\alpha}|=|v|+|\alpha|$ and verifies $u^{\alpha}e(v)=e(v_{\alpha})$.
Proof.
We denote $e=e(v)$ and proceed by induction on $|\alpha|$. If $|\alpha|=1$ we have according to the definition of $u^{\epsilon}$ that $u^{\epsilon}(e)=e\otimes e(\epsilon,j_{\epsilon}(v))=e(v_{\epsilon})$. Assuming now $|\alpha|\geq 2$ we have
$u^{\alpha}(e)=u^{\alpha-\epsilon}(u^{\epsilon}(e)).$
We denote $\tilde{\alpha}=\alpha-\epsilon$. Since $u^{\epsilon}(e)=e\otimes e(\epsilon,j_{\epsilon}(v))$, $\operatorname{sign}(\alpha-\epsilon)=\epsilon$ and $|\tilde{\alpha}|=|\alpha|-1<|\alpha|$, one can apply the induction hypothesis on $\tilde{\alpha}$ and conclude. ∎
We can now prove 4.3.
Proof of 4.3.
For the first point we proceed by induction over $|v|\geq 1$. Let $v:=(k,i_{1})=(k,(\epsilon,k_{1}))$. Either $k_{1}=j_{\epsilon}(k)$, then $e(v)=e(k)\otimes e(\epsilon,j_{\epsilon}(k))=u^{\epsilon}(e(k))$, or $j_{\epsilon}(k)\neq k_{1}$ and then denoting $B:=E_{j_{\epsilon}(k),k_{1}}$ and $b=\Pi(B)$ we have that $bu^{\epsilon}(e)=b(e(k)\otimes e(j_{\epsilon}(k),\epsilon))=e(k)\otimes Be(\epsilon,j_{\epsilon}(k))=e(v)$. We now consider $v=(k,i_{1},...,i_{p+1})$ for some $p\geq 1$. For $t<p+1$ we consider $v_{t}:=(k,i_{1},...,i_{t})$ and $e_{t}:=e(v_{t})$. Since $e=e(v)\in\Xi$, we have $v_{t}\in\mathcal{V}$ for all $t<p+1$. Let $\alpha_{0},...,\alpha_{m}\in\mathbb{Z}$ and $b_{1},...,b_{m}$ as described in 4.3 and such that
$e_{p}=u^{\alpha_{m}}b_{m}u^{\alpha_{m-1}}\cdots u^{\alpha_{1}}b_{1}u^{\alpha_{0}}e(k).$
The condition $v\in\mathcal{V}$ imposes that either $k_{p}\neq j_{-\epsilon_{p}}(v_{p-1})$ or $\epsilon_{p}=\epsilon_{p+1}$. Therefore we have
$u^{\epsilon_{p+1}}(e_{p})=e_{p}\otimes e(\epsilon_{p+1},j_{\epsilon_{p+1}}(v_{p})).$
We replace $\alpha_{m}$ by $\alpha_{m}+\epsilon_{p+1}$. If $k_{p+1}=j_{\epsilon_{p+1}}(v_{p})$ we have the result. If $k_{p+1}\neq j_{\epsilon_{p+1}}(v_{p})$ then we consider $B_{m+1}:=E_{k_{p+1},j_{\epsilon_{p+1}}(v_{p})}$ and $b_{m+1}:=\Pi(B_{m+1})$ and we have
$b_{m+1}u^{\alpha_{m}+\epsilon_{p+1}}b_{m}\cdots u^{\alpha_{1}}b_{1}u^{\alpha_{0}}e(k)=e(v).$
We now prove the second assertion and we start with the case where we impose also $\alpha_{m}\neq 0$. We fix $k\in[N]$ and proceed by induction, this time on $p:=\sum_{i=0}^{m}|\alpha_{i}|$. For $p=1$ we necessarily have that $m=1$ and $\alpha_{1}=\alpha\in\{+1,-1\}$. For $B=E_{ij}$ an elementary matrix, we have either $j\neq k$ and then $u^{\alpha}be(k)=be(k)=0$, or $j=k$ and then $u^{\alpha}be(k)=u^{\alpha}e(i)=e(i)\otimes e(\alpha,j_{\alpha}(i))$. We now consider $p\geq 2$. If $m=1$ and $|\alpha_{1}|=m$ then we conclude with 4.4. Otherwise we set
$f:=u^{\alpha_{m-1}}b_{m-1}\cdots u^{\alpha_{1}}b_{1}u^{\alpha_{0}}e(k)$
and $\epsilon_{i}:=\operatorname{sign}(\alpha_{i})$. Since $\alpha_{m}\neq 0$ then by induction we have $v\in\mathcal{V}$ such that we have $f=e(v)=e(v^{-})\otimes e(\epsilon_{m-1},j_{\epsilon_{m-1}}(v^{-}))$. We consider $E_{ij}$ the elementary matrix such that $b_{m}=\Pi(E_{ij})$. If $j_{\epsilon_{m-1}}(v^{-})\neq j$ then $b_{m}f=0$ and therefore $e=0$. Otherwise we have that $b_{m}f=e(v^{-})\otimes e(\epsilon_{m-1},i)$ with $i\neq j_{\epsilon_{m-1}}(v^{-})$. Applying 4.4 to $e=b_{m}f$ and $\alpha=\alpha_{m}$ we obtain the expected result. Now if $\alpha_{m}=0$, knowing that $\alpha_{m-1}\neq 0$ we find $v\in\mathcal{V}$ of length $p=\sum|\alpha_{i}|$ such that we have $u^{\alpha_{m-1}}b_{m-1}\cdots u^{\alpha_{0}}e(k)=\epsilon e(v)$. Applying $b_{m}$ to $e(v)$ we either find 00 or change the last coordinate $i_{p}$ of $v$.
We now prove the freeness with amalgamation of $u$ and $\mathcal{M}$, i.e. that (16) holds true for all $e\in\Xi$ and for all $b_{i}=\Pi(B_{i})$ where $B_{i}\in\mathrm{M}_{N}(\mathbb{C})$ and $\mathrm{E}(B_{i})=0$. Owing to the first assertion of 4.3, we only have to check (16) holds for $e\in H_{\emptyset}$ basis vector. Finally by linearity in $b_{i}$ for all $i\leq n$ it is sufficient to consider that $B_{i}$ is an elementary matrix $E_{k,\ell}$ of $\mathrm{M}_{N}(\mathbb{C})$ with zeros on the diagonal. Now let $k\in[N]$, $m\geq 2$, $\alpha_{0},...,\alpha_{m}\in\mathbb{Z}$ such that $\alpha_{i}\neq 0$ for all $i\notin\{0,m\}$ and $b_{1},...,b_{m}$ such that $b_{i}=\Pi(B_{i})$ where $B_{i}$ elementary matrix. Owing to the second assertion in 4.3, we have $f=u^{\alpha_{m}}b_{m}\cdots u^{\alpha_{1}}b_{1}u^{\alpha_{0}}e(k)=\epsilon e(v)$ for some $v\in\mathcal{V}$ such that $|v|\geq 1$. In particular $f$ is indeed orthogonal to $e$. ∎
4.4 Spectral radius of the limit operator
For $n\in\mathbb{N}$ and $T\in[n]$, we define the set
$\mathcal{L}(T,n):=\left\{\vec{n}=(n_{1},...,n_{T})\in\{0,...,n\}^{T}:\leavevmode\nobreak\ \sum_{t=1}^{T}n_{t}+T-1=n\right\}.$
We first express $\Phi[a^{n_{1}}(a^{*})^{n_{2}}]$ and $\Phi[p^{n_{1}}(p^{*})^{n_{2}}]$ for any integer $n_{1},n_{2}\geq 0$ and $a,p$ defined by (10).
Proposition 4.5.
For all $n_{1},n_{2}\in\mathbb{N}$ we have
$\Phi[a^{n_{1}}(a^{*})^{n_{2}}]=\sum_{T=1}^{n_{1}\land n_{2}}\sum_{\begin{subarray}{c}\vec{n}\in\mathcal{L}(T,n_{1})\\ \leavevmode\nobreak\ \vec{n}^{\prime}\in\mathcal{L}(T,n_{2})\leavevmode\nobreak\ \end{subarray}}\prod_{t=1}^{T}\operatorname{Tr}_{N}\left(Q^{n_{t}}(Q^{*})^{n_{t}^{\prime}}\right)$
and
$\Phi[p^{n_{1}}(p^{*})^{n_{2}}]=\sum_{T=1}^{n_{1}\land n_{2}}\sum_{\begin{subarray}{c}\vec{n}\in\mathcal{L}(T,n_{1})\\ \leavevmode\nobreak\ \vec{n}^{\prime}\in\mathcal{L}(T,n_{2})\leavevmode\nobreak\ \end{subarray}}(1-r)^{2(T-1)}\prod_{t=1}^{T}\operatorname{Tr}_{N}\left(\{rQ\}^{n_{t}}\{rQ^{*}\}^{n_{t}^{\prime}}\right).$
For $t\geq 1$, we consider the sigma algebra
$\mathcal{F}_{t}:=\sigma\left(\{\alpha\},\{j_{\epsilon}(s)\leavevmode\nobreak\ ;s\in\mathcal{S},\leavevmode\nobreak\ |s|\leq t-1,\leavevmode\nobreak\ \epsilon\in\{+,-\}\}\right).$
Then 4.5 is a direct consequence of the following lemma.
Lemma 4.6.
Let $T\geq 1$ be an integer and $b_{1},c_{1},...,b_{T},c_{T}\in\mathcal{M}$ be images of $N\times N$ complex matrices by $\Pi$.
  1. 1.
    If for all $t\leq T$, $b_{i}$ and $c_{i}$ are images of elementary matrices, we set
    $e_{t}:=b_{T}ub_{T-1}\cdots b_{2}ub_{1}e(\alpha)$
    and we have that $e_{t}$ is a random vector such that there exists a unique $w_{t}\in\mathcal{V}$ verifying either $|w_{t}|=t$ and $\epsilon(w_{t})=+$ or $w_{t}=0$. Also $w_{t}$ is $\mathcal{F}_{t}$-measurable.
  2. 2.
    We have
    $\mathbb{E}[\langle b_{1}u^{*}b_{2}\cdots u^{*}b_{T-1}u^{*}b_{T}c_{T}u\cdots uc_{1}e(\alpha),e(\alpha)\rangle]=\prod_{t=1}^{T}\operatorname{Tr}_{N}(b_{t}c_{t})$
    where we identify in notations the matrices $b_{t}$ and $c_{t}$ to theirs images through $\Pi$.
Proof.
We proceed by induction on $T\geq 2$. For $T=2$ we have $be(\alpha)=e(i)\mathds{1}(\{\alpha=j\})$ for $b=E_{ij}$, and $ube(\alpha)=e(w)$ with $w:=\mathds{1}(\{j=\alpha\})(i,+,j_{+}(i))$ and $b^{\prime}ube(\alpha)=e(w^{\prime})$ with $w^{\prime}:=(i,+,k)\mathds{1}(\{j=\alpha,l=j_{+}(i,+,j_{+}(w))\})$ for $b^{\prime}=E_{k,l}$. Let now $T\geq 3$. By induction we have
$e_{T}=b_{T}ue(w_{T-1}).$
If $w_{T-1}=0$, then we set $w_{T}=0$. Otherwise we have $\epsilon(w_{T-1})=+$, Therefore $ue(w_{T-1})=e(w_{T-1})\otimes e(+,j_{+}(w_{T-1}))$ and $b_{T}ue(w_{T-1})=e(w_{T})$ where
$w_{T}:=(w_{T-1},+,i)\mathds{1}_{\{j=j_{+}(w_{T-1})\}}\in\mathcal{F}_{T}$
and $b_{T}=E_{ij}$.
We also proceed by induction on $T\geq 1$ for the second part of the lemma. Since each side of the formula are linear in the variables $b_{t}$ and $c_{t}$ for every $t\leq T$, we only have to check the formula for elementary matrices. The formula is trivial for $T=1$. We consider now $T\geq 2$ and for $t\leq T$, we consider the reel random variable
$I_{t}:=\langle b_{1}u^{*}b_{2}\cdots b_{t-1}u^{*}b_{t}c_{t}uc_{t-1}u\cdots uc_{1}e(\alpha),e(\alpha)\rangle$
and the vectors
$e_{t}$ $:=c_{t}uc_{t-1}u\cdots uc_{1}e(\alpha)$
$e^{\prime}_{t}$ $:=b_{t}ub_{t-1}u\cdots ub_{1}e(\alpha).$
According to the first assertion of the lemma, we find a unique $(w_{T-1},w^{\prime}_{T-1})\in\mathcal{V}^{2}$ of length $T-1$ such that $e_{T-1}=e(w_{T-1})$ and $e^{\prime}_{T-1}=e(w_{T-1}^{\prime})$. Then rewriting $I_{T}$ and supposing that $w_{T-1}$ and $w_{T-1}^{\prime}$ are non-zero and therefore $\epsilon(w_{T-1})=\epsilon(w_{T-1}^{\prime})=+$ we obtain
$I_{T}$ $=\langle u^{*}b_{T}c_{T}ue(w_{T-1}),e(w_{T-1}^{\prime})\rangle$
$=\langle b_{T}c_{T}e(w_{T-1})\otimes e(+,j_{+}(w_{T-1})),e(w_{T-1}^{\prime})\otimes e(j_{+}(w_{T-1}^{\prime}))\rangle$
$=(b_{T}c_{T})_{j_{+}(w_{T-1})}\mathds{1}\left(w_{T-1}=w_{T-1}^{\prime}\right)$
where for $M\in\mathrm{M}_{N}(\mathbb{C})$ we use the notation $(M)_{k}:=(M)_{k,k}$. Noticing that for all $w\in\mathcal{V}$ such that $|w|=T-1$ we have that $j_{+}(w)$ is independent from $\mathcal{F}_{T-1}$ we have
$\mathbb{E}(I_{T}|\mathcal{F}_{T-1})=\sum_{w\in\mathcal{V}\leavevmode\nobreak\ ;|w|=T-1}\mathbb{E}((b_{T}c_{T})_{j_{+}(w)})\mathds{1}(w_{T-1}=w_{T-1}^{\prime}=w)=\operatorname{Tr}_{N}(b_{T}c_{T})\langle e(w_{T-1}),e(w_{T-1}^{\prime})\rangle.$
We conclude taking the expectation and applying the induction hypothesis to $\langle e(w_{T-1}),e(w_{T-1}^{\prime})\rangle$. ∎
Finally we have the following lower bound on the norm of the operator.
Lemma 4.7.
For all $B\in\mathrm{M}_{N}(\mathbb{C})$ and $n\geq 1$ integer we have:
$\|[u+\Pi(B)]^{n}\|\geq 1.$
Proof.
Let $e\in\mathrm{H}$ such that $\|e\|_{2}=1$. We set $b:=\Pi(B)$. We have:
$[u+b]^{n}(e)=u^{n}(e)+\sum_{T=0}^{n}\sum_{\sum_{t=1}^{T}\lambda_{t}=n-T+1}b^{\lambda_{1}}ub^{\lambda_{2}}u\cdots ub^{\lambda_{T}}$
where in the term $b^{\lambda_{1}}ub^{\lambda_{2}}u\cdots ub^{\lambda_{T}}$, the number of occurrences of $u$ is bounded by $n-1$, therefore the sum is orthogonal to $u^{n}(e)$. It remains to take the norm of in the equality above. ∎
5 Main technical statement
Having proved the first statement of 2.2 about the existence of the operator $(q,u)\in\mathcal{B}(\mathrm{H})$, the second statement is proved by applying the following theorem. First note that for any $\ell\geq 0$ integer we have
$\rho(A_{\#}|_{\boldsymbol{1}^{\perp}})=\{\rho(A_{\#}^{\ell}|_{\boldsymbol{1}^{\perp}})\}^{1/\ell}\leq\|A_{\#}|_{\boldsymbol{1}^{\perp}}^{\ell}\|^{1/\ell},$
To simplify further computations, we introduce
$A_{\ell,\#}:=MA_{\#}^{\ell}$
for any $\ell\geq 1$ integer. Since $M$ is a permutation and $Q$ verifies (H:1), $A_{\ell,\#}$ and $A_{\#}^{\ell}$ both leave $\boldsymbol{1}$ and $\boldsymbol{1}^{\perp}$ stable and have the same operator norm. Furthermore for $N$ large enough we have
$1+\epsilon\geq\exp(\frac{\epsilon}{2}),$
Where $\epsilon=\epsilon(N)$ is given by 2.2. Therefore for $N$ large enough
$\mathbb{P}\left(\|A^{\ell}|_{\boldsymbol{1}^{\perp}}\|\geq(1+\epsilon)^{\ell}\rho_{n}^{\ell}\right)=\mathbb{P}\left(\|A_{\ell}|_{\mathbf{1}^{\perp}}\|\geq(1+\epsilon)^{\ell}\rho_{n}^{\ell}\right)\leq\mathbb{P}\left(\|{A_{\ell}|}_{\boldsymbol{1}^{\perp}}\|\geq e^{\frac{\epsilon\ell}{2}}\rho_{n}^{\ell}\right).$
The same reasoning holds replacing $A$ with $A_{r}$ and $\rho_{n}$ with $\rho_{n,r}$. Adjusting the constants, the second statement of 2.2 becomes a corollary of the following result applied to $\ell=\frac{(1-c)\ln(N)}{7K_{\#}^{\prime}\ln(d)}$ where
$K^{\prime}=1\vee(2\ln(1+\delta+\kappa))\leavevmode\nobreak\ \leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ \leavevmode\nobreak\ K^{\prime}_{r}=1\vee\left(2\ln\left(\frac{1+r\delta+r\kappa}{1-r}\right)\right).$
The condition (H:2) implies that $K_{\#}^{\prime}$ is uniformly bounded in $N$.
Theorem 5.1.
Let $n\geq 0$, $(M,Q)$ and $\rho_{n,\#}$ be as in 2.2. For any $10<c<$ 1, there exists $C>0$ depending on $n$, $K$ and $\alpha$, such that for any $1\leq\ell\leq\frac{\ln(N)}{7\ln(d)}$ we have:
$\mathbb{P}\left(\|A_{\#}^{\ell}|_{\boldsymbol{1}^{\perp}}\|\geq e^{C\ln(N)^{1-\alpha}}\rho_{n,\#}^{\ell}\right)\leq N^{-c}+C\frac{\ell^{4}(4d)^{3\ell}}{N}.$
To estimate an upper bound on $\|(A_{\ell,\#})|_{\boldsymbol{1}^{\perp}}\|$, we use the high trace method, i.e. Markov inequality combined with the inequality
$\mathbb{E}\left(\|B\|^{2m}\right)=\mathbb{E}\left(\|BB^{\top}\|^{m}\right)\leq\mathbb{E}\left(\operatorname{Tr}(BB^{\top})^{m}\right)$
for any matrix $B$ and integer $m\geq 1$.
6 Path decomposition and computation on random permutation
6.1 Path decomposition
We will here interpret the coordinates of the matrix $A_{\ell,\#}$ as paths in a graph.
Definition 6.1.
A path of length $k$ is a sequence $\gamma=(x_{1},y_{1},\ell_{1},x_{2},y_{2},...,y_{T},\ell_{T},x_{T+1})$ such that and for all $t\leq T$, we have $x_{t},y_{t}\in[N]$, $\ell_{t}\in\{0,...,k\}$, $(Q^{\ell_{t}})_{y_{t},x_{t+1}}>0$ and
$\sum_{t=1}^{T}\ell_{t}+T-1=k.$
We denote by $\Gamma^{k}$ the set of paths of length $k$. For $x,y\in[N]$ we denote by $\Gamma^{k}_{xy}$ the subset of $\Gamma^{k}$ of paths verifying $x_{1}=x$ and $y_{T}=y$.
By convention we assume that the sum over an empty set is 00 and the product over an empty set is $1$. With the previous definition we have
$(A_{\ell})_{xy}=\sum_{\gamma\in\Gamma_{xy}^{\ell}}\prod_{t=1}^{T_{\gamma}}M_{x_{t}y_{t}}(Q^{\ell_{t}})_{y_{t}x_{t+1}}.$
In the previous formula, we let appear the dependence of $T=T_{\gamma}$ in the choice of path $\gamma\in\Gamma_{xy}^{\ell}$. To make the expressions lighter, from now on we will omit this dependency when there is no confusion. To be able to estimate the previous product we look for paths that are highly probable and better suited for estimation.
Definition 6.2.
Let $\ell\geq 0$ be fixed.
  • We call a coincidence a sequence $c=(x_{1},y_{1},...,x_{T},y_{T},x_{T+1})$ such that $(x_{2},...,x_{T+1})$ are pairwise distinct and such that there exists $(\ell_{1},...,\ell_{T})$ a $T$-tuple of integers verifying $\sum_{t}\ell_{t}+T-1\leq\ell$, $\gamma=(x_{1},y_{1},\ell_{1},x_{2},...,y_{T},\ell_{T},x_{T+1})$ is a path and $(B_{1}^{\ell_{1}}B_{2}^{\ell_{2}}\cdots B^{\ell_{T}}_{T})_{x_{1},x_{T+1}}>0$ where $B_{1},...,B_{T}\in\{Q,Q^{\ast}\}$ are alternate terms (i.e. $B_{t+1}=B_{t}^{\ast}$).
  • We say that a path $\gamma=(x_{1},y_{1},\ell_{1},x_{2},...,x_{T},y_{T},\ell_{T},x_{T+1})$ is tangle-free if the sequence $(x_{1},y_{1},x_{2},...,x_{T},y_{T},x_{T+1})$ contains, as a subsequence, at most one coincidence. We denote by $\mathrm{F}^{k}$ (resp. $\mathrm{F}_{xy}^{k}$) the subset of $\Gamma^{k}$ (resp. $\Gamma_{xy}^{k}$) of path of length $k$ (resp. path of length $k$ from $x$ to $y$) that are tangle-free.
  • The pair $(M,Q)$ is $\ell$-tangle-free if for all $k\in[\ell]$, for all $\gamma=(x_{1},y_{1},\ell_{1},...,y_{T},\ell_{T},x_{T+1})\in\Gamma^{k}\backslash\mathrm{F}^{k}$ we have
    $\prod_{t=1}^{T}M_{x_{t}y_{t}}=0.$
When the pair $(M,Q)$ is $\ell$-tangle-free, we have the following equality
$A_{\ell}=A^{(\ell)}$
where we defined $A^{(\ell)}$ by
$(A^{(\ell)})_{xy}:=\sum_{\gamma\in\mathrm{F}^{\ell}_{xy}}\prod_{t=1}^{T}M_{x_{t}y_{t}}(Q^{\ell_{t}})_{y_{t}x_{t+1}},$ (18)
for $x,y\in[N]$. Similarly we define the matrix $\underline{A}^{(\ell)}$ by its coordinates
$(\underline{A}^{(\ell)})_{xy}:=\sum_{\gamma\in\mathrm{F}^{\ell}_{xy}}\prod_{t=1}^{T}\underline{M}_{x_{t}y_{t}}(Q^{\ell_{t}})_{y_{t}x_{t+1}}$ (19)
where the matrix $\underline{M}$ is the centered variable
$\underline{M}:=M-\frac{1}{N}\boldsymbol{1}\otimes\boldsymbol{1},$
where for all $x,y\in[N]$, $(\boldsymbol{1}\otimes\boldsymbol{1})_{xy}=1$. We want to end up with a similar decomposition than in [BQZ18, Lemma 1]. To do such a decomposition we write the product terms in the expression (19) using the following identity
$\prod_{t=1}^{T}a_{t}=\prod_{t=1}^{T}b_{t}+\sum_{k=1}^{T}\prod_{t=1}^{k-1}b_{t}(a_{k}-b_{k})\cdot\prod_{t=k+1}^{T}a_{t}.$ (20)
However the dependence $T=T(\gamma)$ forces us to further segment the sum (19). Therefore we introduce some notations. For integers $1\leq T\leq\ell$ and a vector of integers $\vec{\ell}\in\mathcal{L}(T,\ell)$, we denote by $\mathrm{F}^{\vec{\ell}}$ (resp. $\mathrm{F}^{\vec{\ell}}_{xy}$) the subset path $\gamma=(x_{1},y_{1},\ell_{1},x_{2},...,x_{T},y_{T},\ell_{T},x_{T+1})\in\mathrm{F}^{\ell}$ (resp $\gamma\in\mathrm{F}^{\ell}_{xy}$) such that $(\ell_{1},...,\ell_{T})=\vec{\ell}$. Then the equality (18) becomes
$(A^{(k)})_{xy}$ $=\sum_{T=1}^{\ell}\sum_{\vec{\ell}\in\mathcal{L}(T,\ell)}\sum_{\gamma\in\mathrm{F}^{\vec{\ell}}_{xy}}\prod_{t=1}^{T}M_{x_{t}y_{t}}(Q^{\ell_{t}})_{y_{t}x_{t+1}}=\sum_{T=1}^{\ell}\sum_{\vec{\ell}\in\mathcal{L}(T,\ell)}(A^{(\vec{\ell})})_{xy}$ (21)
where for $\vec{\ell}\in\mathcal{L}(T,\ell)$ we set
$(A^{(\vec{\ell})})_{xy}=\sum_{\gamma\in\mathrm{F}^{\vec{\ell}}_{xy}}\prod_{t=1}^{T}M_{x_{t}y_{t}}(Q^{\ell_{t}})_{y_{t}x_{t+1}}.$ (22)
We define similarly
$(\underline{A}^{(\vec{\ell})})_{xy}=\sum_{\gamma\in\mathrm{F}^{\vec{\ell}}_{xy}}\prod_{t=1}^{T}\underline{M}_{x_{t}y_{t}}(Q^{\ell_{t}})_{y_{t}x_{t+1}}.$ (23)
We consider the integers $1\leq T\leq\ell$ and $\vec{\ell}\in\mathcal{L}(T,\ell)$ fixed. Owing to (20) we have
$(A^{(\vec{\ell})})_{xy}=(\underline{A}^{(\vec{\ell})})_{xy}+\sum_{\gamma\in\mathrm{F}^{\vec{\ell}}_{xy}}\sum_{k=1}^{T}\left(\prod_{t=1}^{k-1}\underline{M}_{x_{t}y_{t}}(Q^{\ell_{t}})_{y_{t}x_{t+1}}\right)\frac{(Q^{\ell_{k}})_{y_{k}x_{k+1}}}{N}\prod_{t=k+1}^{T}M_{x_{t}y_{t}}(Q^{\ell_{t}})_{y_{t}x_{t+1}}.$
The integer $T$ does not depend on the path $\gamma$ anymore and therefore we have
$(A^{(\vec{\ell})})_{xy}=(\underline{A}^{(\vec{\ell})})_{xy}+\sum_{k=1}^{T}\frac{1}{N}\underbrace{\sum_{\gamma\in\mathrm{F}^{\vec{\ell}}_{xy}}\prod_{t=1}^{k-1}\underline{M}_{x_{t}y_{t}}(Q^{\ell_{t}})_{y_{t}x_{t+1}}(Q^{\ell_{k}})_{y_{k}x_{k+1}}\prod_{v=k+1}^{T}M_{x_{v}y_{v}}(Q^{\ell_{v}})_{y_{v}x_{v+1}}}_{(S^{(\vec{\ell})})_{xy}}.$
We rewrite $(S^{(\vec{\ell})})_{xy}$ so that lower powers of $A^{(\vec{\ell})}$ appear. For all $1\leq k\leq T$ we set $\vec{\ell}(k):=(\ell_{1},...,\ell_{k-1})$ and $\vec{\ell}(k,T):=(\ell_{k+1},...,\ell_{T})$. We denote by $\mathrm{T}^{k,\vec{\ell}}_{xy}$ the set of paths $\gamma:=(x,y_{1},...,y_{T},\ell_{T},y)\in\Gamma_{xy}^{\vec{\ell}}$ that verifies:
  1. 1.
    $\gamma^{\prime}:=(x,y_{1},...,x_{k-1},y_{k-1},\ell_{k-1},x_{k})\in\mathrm{F}^{\vec{\ell}(k)}$,
  2. 2.
    $\gamma^{\prime\prime}:=(x_{k+1},y_{k+1},...,x_{T},y_{T},\ell_{T},x_{T+1})\in\mathrm{F}^{\vec{\ell}(k,T)}$,
  3. 3.
    $\gamma=(\gamma^{\prime},y_{k},\ell_{k},\gamma^{\prime\prime})$ is tangled.
To rewrite $S(\vec{\ell},x,y)$, we decompose the last summand noticing that we have
$\mathrm{F}_{xy}^{\vec{\ell}}\sqcup\mathrm{T}_{xy}^{k,\vec{\ell}}=\bigsqcup_{x_{k}\in[N]}\bigsqcup_{y_{k}\in[N]}\bigsqcup_{x_{k+1}\in[N]}\{\gamma=(\gamma^{\prime},y_{k},\ell_{k},\gamma^{\prime\prime})\leavevmode\nobreak\ \leavevmode\nobreak\ ;\leavevmode\nobreak\ \leavevmode\nobreak\ \gamma^{\prime}\in\mathrm{F}^{\vec{\ell}(k)}_{xx_{k}},\leavevmode\nobreak\ \gamma^{\prime\prime}\in\mathrm{F}^{\vec{\ell}(k,T)}_{x_{k+1}y}\}$
and introducing the rest
$(R_{k}^{(\vec{\ell})})_{xy}:=\sum_{\gamma\in\mathrm{T}^{k,\vec{\ell}}_{xy}}\prod_{t=1}^{k-1}\underline{M}_{x_{t}y_{t}}(Q^{\ell_{t}})_{y_{t}x_{t+1}}(Q^{\ell_{k}})_{y_{k}x_{k+1}}\prod_{v=k+1}^{T}M_{x_{v}y_{v}}(Q^{\ell_{v}})_{y_{v}x_{v+1}}.$ (24)
We then have
$(S^{(\vec{\ell})})_{xy}$ $=\sum_{x_{k},y_{k},x_{k+1}=1}^{N}\sum_{\begin{subarray}{c}\gamma^{\prime}\in\mathrm{F}_{xx_{k}}^{\vec{\ell}(k)}\\ \gamma^{\prime\prime}\in\mathrm{F}^{\vec{\ell}(k,T)}_{x_{k+1}y}\end{subarray}}\prod_{t=1}^{k-1}\underline{M}_{x_{t}y_{t}}(Q^{\ell_{t}})_{y_{t}x_{t+1}}(Q^{\ell_{k}})_{y_{k}x_{k+1}}\prod_{v=k+1}^{T}M_{x_{v}y_{v}}(Q^{\ell_{v}})_{y_{v}x_{v+1}}-(R_{k}^{(\vec{\ell})})_{xy}$
$=\sum_{y_{k}=1}^{N}\sum_{x_{k}=1}^{N}\left(\underline{P}^{(\vec{\ell}(k))}\right)_{xx_{k}}(Q^{\ell_{k}})_{y_{k}x_{k+1}}\sum_{x_{k+1}=1}^{n}\left(A^{(\vec{\ell}(k,T))}\right)_{x_{k+1}y}-(R_{k}^{(\vec{\ell})})_{xy}$
$=\left(A^{(\vec{\ell}(k))}Q^{\ell_{k}}(1\otimes 1)A^{(\vec{\ell}(k,T))}\right)_{xy}-(R_{k}^{(\vec{\ell})})_{xy}$
and therefore we obtain
$A^{(\vec{\ell})}=\underline{A}^{(\vec{\ell})}+\frac{1}{N}\sum_{k=1}^{T-1}\underline{A}^{(\vec{\ell}(k))}Q^{\ell_{k}}(1\otimes 1)A^{(\vec{\ell}(k,T))}-\frac{1}{N}\sum_{k=1}^{T}R_{k}^{(\vec{\ell})}.$
Inserting it into the expression (21) we have
$A^{(\ell)}=\underline{A}^{(\ell)}+\sum_{T=1}^{\ell}\sum_{\vec{\ell}\in\mathcal{L}(T,\ell)}\left[\sum_{k=1}^{T}\underline{A}^{(\vec{\ell}(k))}Q^{\ell_{k}}\boldsymbol{1}\otimes\boldsymbol{1}A^{(\vec{\ell}(k,T))}-\frac{1}{N}\sum_{k=1}^{T}R_{k}^{(\vec{\ell})}\right].$
To estimate the operator norm of $A^{(\ell)}|_{\boldsymbol{1}^{\perp}}$ we first show that the terms with $\boldsymbol{1}\otimes\boldsymbol{1}$ in the previous formula have no contribution. For $1\leq T\leq\ell$ and $\vec{\ell}\in\mathcal{L}(T,\ell)$, we set the notations $\ell(k):=\sum_{t=1}^{k-1}\ell_{t}+(k-1)-1$ and $\ell(k,T):=\sum_{t=k+1}^{T}\ell_{t}+T-k-1$. We then have $\ell(k)+\ell(k,T)+\ell_{k}=\ell-2$ and we can partition $\mathcal{L}(\ell,T)$ as follows
$\mathcal{L}(\ell,T)=\bigsqcup_{\ell_{k}=0}^{\ell-T+1}\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \bigsqcup_{\ell(k)=0}^{\ell-\ell_{k}-2}\mathcal{L}(\ell(k),k-1)\times\{\ell_{k}\}\times\mathcal{L}(\ell-\ell_{k}-2-\ell(k),T-k)$
We consider the matrix
$\phi(\ell)$ $:=\sum_{T=1}^{\ell}\sum_{k=1}^{T}\sum_{\vec{\ell}\in\mathcal{L}(T,\ell)}\underline{A}^{(\vec{\ell}(k))}Q^{\ell_{k}}\boldsymbol{1}\otimes\boldsymbol{1}A^{(\vec{\ell}(k,T))}$
$=\sum_{T=1}^{\ell}\sum_{k=1}^{T}\sum_{\ell_{k}=0}^{\ell}\sum_{\ell(k)=0}^{\ell-\ell_{k}-2}\sum_{\begin{subarray}{c}\vec{\ell}(k)\in\mathcal{L}(\ell(k),k-1)\\ \vec{\ell}(k,T)\in\mathcal{L}(\ell-\ell_{k}-\ell(k)-2,T-k)\end{subarray}}\underline{A}^{(\vec{\ell}(k))}Q^{\ell_{k}}\boldsymbol{1}\otimes\boldsymbol{1}A^{(\vec{\ell}(k,T))}$
$=\sum_{T=1}^{\ell}\sum_{k=1}^{T}\sum_{\ell_{k}=0}^{\ell}\sum_{\ell(k)=0}^{\ell-\ell_{k}-2}\underline{A}^{(\ell(k))}Q^{\ell_{k}}\boldsymbol{1}\otimes\boldsymbol{1}A^{(\ell-\ell_{k}-\ell(k)-2)}.$
[…]
$\boldsymbol{1}\otimes\boldsymbol{1}A^{(\ell^{\prime})}w=\boldsymbol{1}\otimes\boldsymbol{1}A_{\ell^{\prime}}w=0.$
In particular on this event we have $\phi(\ell)|_{\boldsymbol{1}^{\top}}=0$. Besides denoting
$R_{k}^{(\ell)}:=\sum_{T=1}^{\ell}\sum_{\vec{\ell}\in\mathcal{L}(T,\ell)}R_{k}^{(\vec{\ell})},$
we have
$A^{(\ell)}=\underline{A}^{(\ell)}+\phi(\ell)-\frac{1}{N}\sum_{k=1}^{\ell}R^{(\ell)}_{k}.$
To have consistent writings between the operators $R_{k}^{(\ell)}$ and $A^{(\ell)}$, for $1\leq k\leq\ell$ and $x,y\in[N]$ we set
$\mathrm{T}_{xy}^{k,\ell}:=\bigsqcup_{T=1}^{\ell}\bigsqcup_{\vec{\ell}\in\mathcal{L}(\ell,T)}\mathrm{T}_{xy}^{k,\vec{\ell}}$
where we consider $\mathrm{T}_{xy}^{k,\vec{\ell}}=\emptyset$ if $k>T$. With this convention and notations we have
$(R_{k}^{(\ell)})_{xy}=\sum_{\gamma\in\mathrm{T}_{xy}^{k,\ell}}\left(\prod_{t=1}^{k-1}\underline{M}_{x_{t}y_{t}}(Q^{\ell_{t}})_{y_{t}x_{t+1}}\right)(Q^{\ell_{k}})_{y_{k}x_{k+1}}\left(\prod_{t=k+1}^{T}M_{x_{t}y_{t}}(Q^{\ell_{t}})_{y_{t}x_{t+1}}\right).$
When it comes to the case #$\#=r$, all the previous reasoning holds if now one considers $A_{r}=(1-r)M+rQ$ and replace respectively (22) and (23) with
$(A_{r}^{(\vec{\ell})})_{xy}=\sum_{\gamma\in\mathrm{F}^{\vec{\ell}}_{xy}}(1-r)^{T}\prod_{t=1}^{T}M_{x_{t}y_{t}}(\{rQ\}^{\ell_{t}})_{y_{t}x_{t+1}}\leavevmode\nobreak\ \leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ \leavevmode\nobreak\ (\underline{A}_{r}^{(\vec{\ell})})_{xy}=\sum_{\gamma\in\mathrm{F}^{\vec{\ell}}_{xy}}(1-r)^{T}\prod_{t=1}^{T}\underline{M}_{x_{t}y_{t}}(\{rQ\}^{\ell_{t}})_{y_{t}x_{t+1}}.$
We also have a new rest term
$(R_{k,r}^{(\vec{\ell})})_{xy}:=(1-r)^{T-1}\sum_{\gamma\in\mathrm{T}^{k,\vec{\ell}}_{xy}}\prod_{t=1}^{k-1}\underline{M}_{x_{t}y_{t}}\{rQ\}^{\ell_{t}})_{y_{t}x_{t+1}}\{rQ\}^{\ell_{k}})_{y_{k}x_{k+1}}\prod_{v=k+1}^{T}M_{x_{v}y_{v}}\{rQ\}^{\ell_{v}})_{y_{v}x_{v+1}}.$ (25)
We have now proved the following lemma.
Lemma 6.3.
[…]
$\|A_{\#}^{\ell}|_{\boldsymbol{1}^{\perp}}\|\leq\|\underline{A}_{\#}^{(\ell)}\|+\frac{1}{N}\sum_{k=1}^{\ell}\|R^{(\ell)}_{k,\#}\|.$
6.2 Computations on random permutations
In this section we check that for $M\in\mathrm{S}_{N}$ uniformly distributed and $Q$ verifying (H:1), (H:2) and (H:3), the pair $(M,Q)$ is $\ell$-tangle free with high probability, provided that the integer $\ell$ is not too large before the dimension.
Lemma 6.4.
There exists $C>0$ such that, for any integer $\ell\geq 1$ the pair $(M,Q)$ is $\ell$-tangle free with probability at least $1-C\frac{\ell^{4}(4d)^{3\ell}}{N}$.
Proof.
We consider $1\leq\ell\leq\frac{N}{2}$. We say that a path $\gamma:=(x_{1},y_{1},\ell_{1},x_{2}...,x_{T},\ell_{T-1},y_{T},\ell_{T},x_{T+1})$ occurs if for all $t\leq T$, we have $M_{x_{t},y_{t}}=1$. If $(M,Q)$ is $\ell$-tangled then there exists $1\leq T\leq\ell$, $\ell_{1},...,\ell_{T}$ integers such that $\sum_{i=1}^{T}\ell_{i}+T-1\leq\ell$ and a path $\gamma:=(x_{1},y_{1},\ell_{1},x_{2},y_{2},...,x_{T},\ell_{T},y_{T},\ell_{T},x_{T+1})$ that occurs such that there exists $1\leq k,i<k^{\prime}\leq T$ such that either
  1. 1.
    $(x_{1},y_{1},...,x_{k},y_{k},x_{k+1})$ and $(x_{i},y_{i},...,x_{T},y_{T},x_{T+1})$ are two distinct coincidences where $x_{t}$ are pairwise distinct except possibly $x_{1}=x_{k+1}$ and $x_{i}=x_{T+1}$. We denote by $I_{1}(k,i)$ this event,
  2. 2.
    $(x_{1},y_{1},...,x_{T},y_{T},x_{T+1})$ and $(x_{k},y_{k},...,x_{k^{\prime}+1})$ are two coincidences (one inside the other) where all $x_{t}$ are pairwise distinct except possibly $x_{1}=x_{T+1}$. We denote by $I_{2}(k,k^{\prime})$ the corresponding event.
The probability for a path $\gamma=(x_{1},y_{1},\ell_{1},x_{2},...,x_{T+1})$ to occur is bounded by
$\mathbb{P}(\gamma)=\mathbb{P}(\prod_{t=1}^{T}\mathds{1}\{\sigma(x_{t})=y_{t}\})\leq\frac{1}{(N)_{T}},$
where $(N)_{t}:=N(N-1)(N-2)\cdots(N-t+1)$. Let $\{x_{t}\}_{t}$ be pairwise distinct elements of $[N]$, except possibly $x_{1}=x_{k+1}$ and $x_{i}=x_{T+1}$ for some $1\leq k,i<T\leq\ell$. We count the number of possible $\gamma$ verifying $I_{1}(k,i)$. For $\{x_{t}\}_{t\notin\{k,T+1\}}$ we have $(N)_{T-1}$ possibilities in regard of the condition of being pairwise distinct. For each $(\ell^{\prime}_{t})_{t\leq T^{\prime}}$ such that $\sum_{t}\ell^{\prime}_{t}+T^{\prime}-1\leq\ell$, we have at most $2d^{\tiny{\sum_{t}\ell^{\prime}_{t}}}\leq 2d^{\ell}$ possibilities for $x_{k+1}$ to have $(B_{1}^{\ell_{1}^{\prime}}\cdots B_{T^{\prime}}^{\ell_{T^{\prime}}^{\prime}})_{x_{1}x_{k+1}}>0$ with $B_{t}\in\{Q,Q^{\ast}\}$ and $B_{t+1}=B_{t}^{\ast}$. Besides the number of $T^{\prime}\leq\ell$ and $(\ell_{t}^{\prime})_{t\leq T^{\prime}}$ such that $\sum_{t}\ell^{\prime}_{t}+T^{\prime}-1\leq\ell$ is given by:
$|\bigsqcup_{p=1}^{\ell}\bigsqcup_{T^{\prime}=1}^{p}\{(\ell^{\prime}_{t})_{t\leq T^{\prime}}:\leavevmode\nobreak\ \sum_{t=1}^{T^{\prime}}\ell_{t}^{\prime}+T^{\prime}-1=p\}|=\sum_{p=1}^{\ell}\sum_{T^{\prime}=1}^{p}\binom{p-1}{T^{\prime}-1}\leq 2^{\ell}.$
Therefore we have at most $(4d)^{2\ell}$ possibilities for $x_{k+1}$ and $x_{T+1}$. For fixed $1\leq T\leq\ell$ and $\vec{\ell}=(\ell_{t})_{t\leq T}\in\mathcal{L}(\ell,T)$ we have at most $d^{\sum_{t}\ell_{t}}=d^{\ell-T+1}\leq d^{\ell}$ possibilities for the $y_{t}$. To chose $(\ell_{t})_{t\leq T}$ we have again $2^{\ell}$ possibilities. We have
$\mathbb{P}\left(I_{k,i}\right)\leq\frac{(N)_{T-1}(4d)^{2\ell}(2d)^{\ell}}{(N)_{T}}\leq\frac{(4d)^{3\ell}}{N}.$
We proceed similarly to bound the number of $\gamma$ verifying $I_{2}(k,k^{\prime})$ and we obtain the exact same bound. Now summing over all $T$, $k$, $i$ and $k^{\prime}$ we obtain the result.
Let $\mathbf{x}:=(x_{1},...,x_{k})$, $\mathbf{y}:=(y_{1},...,y_{k})\in[N]^{k}$. We are looking for an estimate of
$\mathbb{E}\left(\prod_{t=1}^{k_{0}}\underline{M}_{x_{t}y_{t}}\prod_{t=k_{0}+1}^{k}M_{x_{t}y_{t}}\right)$
where $0\leq k_{0}\leq k$. To do this, we define the set of arcs of $(\mathbf{x},\mathbf{y})$ as:
$\mathrm{A}_{\mathbf{x,y}}:=\{(x_{t},y_{t}):t\in[k]\}.$
The multiplicity of $e\in\mathrm{A}_{\mathbf{x,y}}$ is given by $m(e):=\sum_{t=1}^{k}\mathds{1}((x_{t},y_{t})=e)$. An arc $e=(x,y)$ is said to be consistent if $\{t:(x_{t},y_{t})=(x,y)\}=\{t:x_{t}=x\}=\{t:y_{t}=y\}$. Otherwise, it is inconsistent. The following proposition is proved in [Bor15, Proposition 11].
Proposition 6.5.
There exists a constant $c>0$ such that for any $\mathbf{x}:=(x_{1},...,x_{k}),\mathbf{y}:=(y_{1},...,y_{k})\in[N]^{k}$ with $2k\leq\sqrt{N}$ and any $k_{0}\leq k$, we have:
$\left|\mathbb{E}\large(\prod_{t=1}^{k_{0}}\underline{M}_{x_{t}y_{t}}\prod_{t=k_{0}+1}^{k}M_{x_{t}y_{t}})\right|\leq c2^{b}\left(\frac{1}{N}\right)^{a}\left(\frac{3k}{\sqrt{N}}\right)^{a_{1}},$
where $a:=|\mathrm{A}_{\mathbf{x,y}}|$, b is the number of inconsistent arcs of $(\mathbf{x,y})$ and $a_{1}$ is the number of $1\leq t\leq k_{0}$ such that $(x_{t},y_{t})$ is inconsistent and has multiplicity one.
7 High trace method
7.1 Operator norm of $\underline{A}_{\#}^{(\ell)}$
In this section we prove the following assertion.
Proposition 7.1.
Let $(Q,M),\leavevmode\nobreak\ \rho_{n}$ be as defined in 2.2 and for #$\#\in\{\emptyset,r\}$ we consider $A_{\#}$ defined by (3) and (4). For any $c>0$ and $1\leq\ell\leq\frac{\ln(N)}{7K^{\prime}_{\#}\ln(d)}$, there exists $C=C(c,\alpha,n,K)>0$ such that with probability at least $1-N^{-c}$,
$\|\underline{A}_{\#}^{(\ell)}\|\leq e^{C\ln(N)^{1-\alpha}}\rho_{n,\#}^{\ell}.$
We give the details of computations in the case where #=#\#$\emptyset$# = ∅. For any integer $m\geq 1$ and considering as convention $x_{2m+1}:=x_{1}$ we have
$\|\underline{A}^{(\ell)}\|^{2m}=\|\underline{A}^{(\ell)}{\underline{A}^{(\ell)}}^{\top}\|^{m}\leq\operatorname{Tr}\{\left(\underline{A}^{(\ell)}{\underline{A}^{(\ell)}}^{\top}\right)^{m}\}$
$=\sum_{x_{1},...,x_{2m}}\prod_{i=1}^{m}(\underline{A}^{(\ell)})_{x_{2i-1},x_{2i}}(\underline{A}^{(\ell)})_{x_{2i+1},x_{2i}}$
$=\sum_{x_{1},...,x_{2m}}\prod_{i=1}^{m}\bigg{\{}\sum_{\gamma_{2i-1}\in\mathrm{F}^{\ell}_{x_{2i-1},x_{2i}}}\prod_{t=1}^{T_{2i-1}}\underline{M}_{x_{2i-1,t},y_{2i-1,t}}(Q^{\ell_{2i-1,t}})_{y_{2i-1,t}x_{2i-1,t+1}}\cdot$
$\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \sum_{\gamma_{2i}\in\mathrm{F}^{\ell}_{x_{2i+1},x_{2i}}}\prod_{t=1}^{T_{2i}}\underline{M}_{x_{2i,t},y_{2i,t}}(Q^{\ell_{2i,t}})_{y_{2i,t}x_{2i,t+1}}\bigg{\}}$
$=\sum_{x_{1},...,x_{2m}}\displaystyle\sum_{\begin{subarray}{c}\gamma_{1},...,\gamma_{2m}\\ \gamma_{2i-1}\in\mathrm{F}^{\ell}_{x_{2i-1},x_{2i}}\\ \gamma_{2i}\in\mathrm{F}^{\ell}_{x_{2i+1},x_{2i}}\end{subarray}}\prod_{i=1}^{m}\prod_{t=1}^{T_{2i-1}}\underline{M}_{x_{2i-1,t},y_{2i-1,t}}(Q^{\ell_{2i-1,t}})_{y_{2i-1,t}x_{2i-1,t+1}}\cdot\prod_{t=1}^{T_{2i}}\underline{M}_{x_{2i,t},y_{2i,t}}(Q^{\ell_{2i,t}})_{y_{2i,t}x_{2i,t+1}}$
where we use the notation $\gamma_{i}=(x_{i,1},y_{i,1},\ell_{i,1},x_{i,2}...,y_{i,T_{i}},\ell_{i,T_{i}},x_{i,T_{i}+1})\in\mathrm{F}^{\ell}$. For $\gamma_{2i-1}\in\mathrm{F}^{\ell}_{x_{2i-1},x_{2i}}$ (resp. for $\gamma_{2i}\in\mathrm{F}^{\ell}_{x_{2i+1},x_{2i}}$) we have with the previous notation $x_{2i-1,1}=x_{2i-1}$ (resp. $x_{2i,1}=x_{2i+1}$) and $x_{2i-1,T_{2i-1}+1}=x_{2i}$ (resp. $x_{2i,T_{2i}+1}=x_{2i}$).
For $\vec{T}=(T_{1},...,T_{2m})\in[\ell]^{2m}$ we introduce the set $\mathrm{W}_{\ell,\vec{T}}$ of $\gamma=(\gamma_{1},...,\gamma_{2m})$ such that for all $1\leq i\leq 2m$, $\gamma_{i}=(x_{i,1},y_{i,1},\ell_{i,1},x_{i,2}...,y_{i,T_{i}},\ell_{i,T_{i}},x_{i,T_{i}+1})\in\mathrm{F}^{\ell}$ and
$x_{2i,1}=x_{2i+1,1}\qquad\text{ and }\qquad x_{2i-1,T_{2i-1}+1}=x_{2i,T_{2i}+1},$ (26)
with the convention $x_{2m+1,1}=x_{1,1}$. We then obtain as a first bound
$\|\underline{A}^{(\ell)}\|^{2m}\leq\sum_{\vec{T}\in[\ell]^{2m}}\sum_{\gamma\in\mathrm{W}_{\ell,\vec{T}}}\prod_{i=1}^{m}\prod_{t=1}^{T_{i}}\underline{M}_{x_{i,t}y_{i,t}}(Q^{\ell_{i,t}})_{y_{i,t}x_{i,t+1}}.$ (27)
[…]
$x_{i,t+1}=x,\leavevmode\nobreak\ \leavevmode\nobreak\ x_{i^{\prime},t^{\prime}+1}=x^{\prime}\leavevmode\nobreak\ \leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ \leavevmode\nobreak\ y_{i,t}=y_{i^{\prime},t^{\prime}}.$
The graph $K_{\gamma}$ induces an equivalence relation on the set $X_{\gamma}$ where the equivalence classes are the connected components of the graph. For $x\in X_{\gamma}$ we denote
$\operatorname{cc}(x):=\text{connected component of }x.$
We also set the set of arcs of $\gamma$, denoted $\mathrm{A}_{\gamma}$ as the set of distinct pairs $(x_{i,t},y_{i,t})$. We then consider the partition $\{\mathrm{W}_{\ell,\vec{T}}(s,a,p)\}_{s,a,p}$ where $\mathrm{W}_{\ell,\vec{T}}(s,a,p)$ is the set of $\gamma\in\mathrm{W}_{\ell,\vec{T}}$ such that $|X_{\gamma}|=s$, $|\mathrm{A}_{\gamma}|=a$ and $|X_{\gamma}/\sim|=|\{\operatorname{cc}(x)\leavevmode\nobreak\ ;\leavevmode\nobreak\ x\in X_{\gamma}\}|=s-p$. We denote $G_{\gamma}=(X_{\gamma}/\sim,\mathrm{A}_{\gamma})$ the graph where for all arc $(x,y)=(x_{i,t},y_{i,t})\in\mathrm{A}_{\gamma}$ we put an edge $(\operatorname{cc}(x_{i,t}),\operatorname{cc}(x_{i,t+1}))$ of color $(x,y)$.
Example 7.1.
We consider the path
$\gamma=(2,1,\ell_{1},2,3,\ell_{2},4,6,\ell_{3},7,3,\ell_{4},8).$
We represented the corresponding graphs $G_{\gamma}$ and $K_{\gamma}$ in Figure 6. We have in this case $X_{\gamma}:=\{2,4,8,7\}$, a unique edge in $K_{\gamma}$ given by $\{4,8\}$ (dashed in Figure 6), $(X_{\gamma}/\sim)=\{\operatorname{cc}_{1},\operatorname{cc}_{2},\operatorname{cc}_{3}\}$ and $\mathrm{A}_{\gamma}=\{(2,1),(2,3),(4,6),(7,3)\}$. One can also refer to [BDH22, Figure 4,5] and [Bor15, Figure 3,4] for representation of other examples of this graphs.
Figure 6: On the left hand side a representation of $G_{\gamma}$ and on the right hand side a representation of $K_{\gamma}$.
Proceeding similarly in the case where #$\#=r$ and taking the expectation of (27) we have
$\mathbb{E}\|\underline{A}_{\#}^{(\ell)}\|^{2m}\leq\sum_{s,a,p,\vec{T}}\sum_{\leavevmode\nobreak\ \gamma\in\mathrm{W}_{\ell,\vec{T}}(s,a,p)}\mu(\gamma)q_{\#}(\gamma),$
where for $\gamma\in\mathrm{W}_{\ell,\vec{T}}$ we set
$\mu(\gamma):=\mathbb{E}\prod_{i=1}^{2m}\prod_{t=1}^{T_{i}}\underline{M}_{x_{i,t}y_{i,t}},$ (28)
and
$\leavevmode\nobreak\ q(\gamma):=\prod_{i=1}^{2m}\prod_{t=1}^{T_{i}}(Q^{\ell_{i,t}})_{y_{i,t}x_{i,t+1}}\leavevmode\nobreak\ \leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ \leavevmode\nobreak\ q_{r}(\gamma):=(1-r)^{T}\prod_{i=1}^{2m}\prod_{t=1}^{T_{i}}(\{rQ\}^{\ell_{i,t}})_{y_{i,t}x_{i,t+1}}.$ (29)
To estimate the above sum we need to decompose further the set $\mathrm{W}_{\ell,\vec{T}}(s,a,p)$. For $\gamma,\gamma^{\prime}\in\mathrm{W}_{\ell,\vec{T}}(s,a,p)$, we say that $\gamma\sim\gamma^{\prime}$ if there exists a pair of permutations $(\alpha,\beta)\in\mathrm{S}_{N}$ such that $K_{\gamma^{\prime}}$ is the image by $\alpha$ of $K_{\gamma}$ i.e. for any $(i,t)$ we have $x^{\prime}_{i,t}=\alpha(x_{i,t})$ and $y_{i,t}^{\prime}=\beta(y_{i,t})$ (where we consider $\gamma^{\prime}=(\gamma_{1}^{\prime},...,\gamma_{2m}^{\prime})$ and for all $i\in[2m]$, $\gamma_{i}^{\prime}=(x_{i,1}^{\prime},y_{i,1}^{\prime},\ell_{i,1}^{\prime},x_{i,2}^{\prime},...,x_{i,T_{i}}^{\prime},y_{i,T_{i}}^{\prime},\ell_{i,T_{i}}^{\prime},x_{i,T_{i}+1}^{\prime})$). We denote by $\mathcal{W}_{\ell,\vec{T}}(s,a,p)$ the set of equivalent classes. For all $\gamma^{\prime}\sim\gamma$, we have $\mu(\gamma)=\mu(\gamma^{\prime})$.
where for $\gamma\in\mathrm{W}_{\ell,\vec{T}}$ we set $T=\sum_{i=1}^{2m}T_{i}$. Finally we have
$\mathbb{E}\|\underline{A}_{\#}^{(\ell)}\|^{2m}\leq\sum_{s,a,p,\vec{T}}|\mathcal{W}_{\ell,\vec{T}}(s,a,p)|\underset{\gamma\in\mathrm{W}_{\ell,\vec{T}}(s,a,p)}{\max}\left\{|\mu(\gamma)|\sum_{\begin{subarray}{c}\gamma^{\prime}\in\mathrm{W}_{\ell,\vec{T}}(s,a,p)\\ \gamma^{\prime}\sim\gamma\end{subarray}}q_{\#}(\gamma^{\prime})\right\}.$ (30)
First we bound the cardinality of $\mathcal{W}_{\ell,\vec{T}}(s,a,p)$.
Lemma 7.2.
We have
$|\mathcal{W}_{\ell,\vec{T}}(s,a,p)|\leq(\ell a(s-p)^{2}s^{2})^{2m(g+5)}.$
The following section is dedicated to the proof of 7.2.
7.2 Encoding equivalent classes
In all this section we consider the integers $1\leq T\leq\ell$, $m\geq 1$ and $1\leq s,a,p\leq 2m\ell$ fixed. To bound $|\mathcal{W}_{\ell,\vec{T}}(s,a,p)|$ we need an efficient way to encode in a unique way an equivalent class $[\gamma]\in\mathcal{W}_{\ell,\vec{T}}(s,a,p)$. To proceed we start by introducing the order of apparition of the elements while we explore chronologically $\gamma$. This will allow us to code $[\gamma]$ by coding what we will call its canonical path. To do so we consider $\boldsymbol{T}:=\{(i,t)\leavevmode\nobreak\ ;\leavevmode\nobreak\ i\in[2m],\leavevmode\nobreak\ t\in[T_{i}]\}$ the set of times, that we consider ordered canonically, i.e. for $i<i^{\prime}$ we have for any $t\in[T_{i}]$ and $t^{\prime}\in[T_{i^{\prime}}]$ $(i,t)\prec(i^{\prime},t^{\prime})$ and for $1\leq t<t^{\prime}\leq T_{i}$ we have $(i,t)\prec(i,t^{\prime})$. For $(i,t)\neq(1,1)$ we denote by $(i,t)^{-}$ the greatest element of $\boldsymbol{T}$ strictly smaller than $(i,t)$. For a time $(i,t)\in\boldsymbol{T}$ we set $\gamma_{i,t}:=(x_{i,t},{y}_{i,t},{x}_{i,t+1})$. For all $x\in X_{\gamma}$ (resp. $y\in Y_{\gamma}$) we denote by $\bar{x}$ (resp. $\bar{y}$) the order of apparition of $x$ (resp. of $y$) in $\gamma$. We consider $\bar{\gamma}:=(\bar{x}_{1,1},\bar{y}_{1,1},\ell_{1,1},\bar{x}_{1,2},...,\bar{y}_{i,t},\ell_{i,t},\bar{x}_{i,t+1},...,\bar{y}_{2m,T_{2m}},\ell_{2m,T_{2m}},\bar{x}_{2m,T_{2m}+1})$ the canonical path and knowing $\bar{\gamma}$ is equivalent to know $[\gamma]$. In fact, we do not have necessarily $\bar{\gamma}\in[\gamma]$, for instance $\bar{\gamma}_{i}=(\bar{x}_{i,1},\bar{y}_{i,1},\ell_{i,1},\bar{x}_{i,2},...)$ can be tangled. However, if $K_{\bar{\gamma}}$ is constructed as we did for $K_{\gamma}$ then for all $\gamma^{\prime}\in[\gamma]$, the graph $K_{\gamma^{\prime}}$ is isomorphic to $K_{\bar{\gamma}}$.
We define a graph which genius gives a first condition on the integers $s,a$ and $p$ for $\mathrm{W}_{\ell,\vec{T}}$ not to be empty. For $x\in X_{\gamma}$, we introduce the color of the connected component $\operatorname{cc}(x)$ defined by
$\bar{c}(\bar{x}):=\min\{\bar{x}\leavevmode\nobreak\ ;\leavevmode\nobreak\ x\in\operatorname{cc}(x^{\prime})\}.$
For $(i,t)$ a time we denote by $\vec{x}_{i,t+1}:=(\bar{c}(\bar{x}_{i,t+1}),\bar{x}_{i,t+1})$ the connected component mark of $x_{i,t+1}$. For example $\bar{x}_{1,1}=\bar{y}_{1,1}=1$ and $\vec{x}_{1,1}=(1,1)$. If $x_{1,2}\neq x_{1,1}$ then $\bar{x}_{1,2}=2$. One can notice that $x,x^{\prime}\in X_{\gamma}$ being in the same connected component might be an information posterior to the discovery of both $x$ and $x^{\prime}$. Therefore $x_{1,1}$ and $x_{1,2}$ might be in the same connected component and in that case we have $\vec{x}_{1,2}=(1,2)$. Otherwise we have $\vec{x}_{1,2}=(2,2)$. Finally we set $\bar{\gamma}_{i,t}:=(\vec{x}_{i,t},\bar{y}_{i,t},\vec{x}_{i,t+1})$ and we have that knowing $(\bar{\gamma}_{i,t})_{(i,t)\in\boldsymbol{T}}$ is equivalent to knowing the equivalence class of $\gamma$.
The vertex set of the graph at stake here is $V_{\gamma}:=[s-p]\simeq X_{\gamma}/\sim$ thought as the set of connected components of $K_{\gamma}$ ordered by their order of apparition. For all $i\in[2m]$ and all $t\in[T_{i}]$ we define the edge $e_{i,t}:=(\bar{\operatorname{cc}}(x_{i,t}),\bar{\operatorname{cc}}(x_{i,t+1}))$ where $\bar{\operatorname{cc}}(x)\in[s-p]$ for $x\in X_{\gamma}$ is the order of apparition of $\operatorname{cc}(x)$ in $\gamma$. We also associate the color $(\bar{x}_{i,t},\bar{y}_{i,t})$ to the edge $e_{i,t}$. Finally we denote by $E_{\gamma}$ the set of colored edges $e_{i,t}$ for all $i\in[2m]$ and $t\in[T_{i}]$ and $G_{\gamma}=(V_{\gamma},E_{\gamma})$ the corresponding graph.
Lemma 7.3.
If $\mathrm{W}_{\ell,\vec{T}}(s,a,p)\neq\emptyset$ then $g:=a+p-s+1\geq 0$.
Proof.
In regard of the construction of $G_{\gamma}$ we have $|E_{\gamma}|=|A_{\gamma}|=a$. Besides, owing to (26), $G_{\gamma}$ is weakly connected and therefore if it is not empty then its genius $g$ is non-negative i.e.
$g:=|E_{\gamma}|-|V_{\gamma}|+1=a-s+p+1\geq 0.$
We give a second condition on the fixed integers of the section for $\mathrm{W}_{\ell,\vec{T}}$ to be not empty. This condition is given by following lemma.
Lemma 7.4.
If $\mathrm{W}_{\ell,\vec{T}}(s,a,p)\neq\emptyset$ then $2g+2m\leq p$.
The rough idea behind the proof of the previous lemma can be described as follows. On the one hand we have that $p=\sum_{k}(|\operatorname{cc}_{k}|-1)$ where $\operatorname{cc}_{k}$ is the $k$-th discovered connected component of $K_{\gamma}$. Therefore it can be seen as the number of vertices that are discovered after we discovered the underlying connected component of $K_{\gamma}$. On the other hand, the number of excess edges $g$ and $m$ the number of times where we impose the condition (26) will give control of the number of such vertices. To proceed we define an underlying covering tree $\Gamma\subset G_{\gamma}$ as the result of a chronologically growing forest $(\Gamma_{i,t})_{(i,t)\in\boldsymbol{T}}$ that will allow us to keep track of the times where vertices of $K_{\gamma}$ are discovered.
We consider the sequence $(G_{i,t})_{(i,t)\in\boldsymbol{T}}$ of graphs such that for $(i,t)\in\boldsymbol{T}$, $G_{i,t}$ is the graph spanned by $e_{(j,s)}$ for all $(j,s)\in\boldsymbol{T}$ such that $(j,s)\preceq(i,t)$. They are non-decreasing sub-graphs of $G_{\gamma}$ and we have $G_{2m,T_{2m}}=G_{\gamma}$. We then inductively define the sequence of non-decreasing forests of $G_{i,t}$ as follows. By convention $\Gamma_{1,0}$ is the graph with the vertex set $\{1\}$ and no edges. Then for a time $(i,t)\in\boldsymbol{T}$ we construct $\Gamma_{i,t}$ from $\Gamma_{(i,t)^{-}}$ as follows. We say that $(i,t)$ is a first time if adding $e_{i,t}$ to the graph $\Gamma_{(i,t)^{-}}$ does not create any cycle. In this case $\Gamma_{i,t}$ is defined as the graph spanned by $\Gamma_{(i,t)^{-}}$ and the edge $e_{i,t}$. If $(i,t)$ is not a first time we set $\Gamma_{i,t}=\Gamma_{(i,t)^{-}}$. We set $\Gamma:=\Gamma_{2m,T_{2m}}$, covering tree of $G_{\gamma}$. We call an excess edge all edges $e\in G_{\gamma}\backslash\Gamma$ and we say that $(i,t)$ is an important time if $e_{i,t}$ is an excess edge. We now give a sequence of $3$-tuple indexed on $\boldsymbol{T}_{\mathrm{imp}}$ the set of important time such that, knowing this sequence allows to recover $\bar{\gamma}$.
In regard of the relation (26) $\Gamma_{i,t}$ has at most two connected component. More precisely
  • For $i$ odd $\Gamma_{i,t}$ is a tree.
  • For $i$ even $\Gamma_{i,t}$ has at most two connected component and $\Gamma_{i,T_{i}}$ is connected.
For $i$ even we call merging time the time $(i,t_{i})$ at which $\Gamma_{i,t_{i}}$ becomes connected. Possibly we have $t_{i}=1$ and if we have $t_{i}\geq 2$ then $(i,t_{i})$ is also a first time. With these definitions we can now prove the previous lemma.
Proof of 7.4.
For all $1\leq k\leq s-p$, the first seen element of the component $\operatorname{cc}_{k}\in X_{\gamma}/\sim$ is discovered at a first time with a tree edge and the others are either discovered by an excess edge seen for the first time or at a merging time. Each excess edge or tree edge discovered at a merging time can discover two vertices that had not been seen before. Therefore we have $p$ bounded by twice the number of merging times (at most $m$) plus twice the number of excess edges. ∎
Exploring the path $\gamma$ through time $(i,t)\in\boldsymbol{T}$ is a successive repetition of the pattern (1)(2)(3) given by
  1. (1)
    (possibly empty) a sequence of first time which are not merging times;
  2. (2)
    (necessarily) an important time or a merging time;
  3. (3)
    (possibly empty) a path using the colored edges of the tree defined so far.
In (3) we only need the in-coming and out-going vertices to know what happened in the time in between. We may now build our first encoding by marking the important and merging times with information that will be sufficient to recover $(\bar{\gamma}_{i,t})_{(i,t)\in\boldsymbol{T}}$ entirely. For $(i,t)$ a time, we assume known $(\bar{\gamma}_{j,s})_{(j,s)\prec(i,t)}$ and we assume that so far we have seen $u$ vertices of $X_{\gamma}$ and $v$ of $Y_{\gamma}$. If $(i,t)$ is a first time and not a merging time then we can recover $\bar{\gamma}_{(i,t)}$. Indeed necessarily
  • if $t\geq 2$ then we have $\vec{x}_{i,t}=\vec{x}_{i,(t-1)+1}$, $\vec{x}_{i,t+1}=(u+1,u+1)$ and $\bar{y}_{i,t}=v+1$,
  • if $t=1$ and $i$ odd then we still have $\vec{x}_{i,t+1}=\vec{x}_{i,2}=(u+1,u+1)$ and $\bar{y}_{i,t}=\bar{y}_{i,2}=v+1$ but now we have $\vec{x}_{i,1}=\vec{x}_{i-1,1}$,
  • finally if $t=1$ and $i$ even, $(i,1)$ not being a merging time implies here that both connected component seen at the time $(i,1)$ are new and therefore we have $\vec{x}_{i,1}=(u+1,u+1)$, $\bar{y}_{i,1}=v+1$ and $\bar{x}_{i,2}=(u+2,u+2)$.
Notice that at a first time that is not a merging time, the $x\in X_{\gamma}$ discovered gives the color of its connected component. If $(i,t)$ is an important time, we first mark it with the vector $(\bar{y}_{i,t},\vec{x}_{i,t+1},\vec{x}_{i,\tau})$ where $(i,\tau)$ is the next time we step outside of the tree $\Gamma_{i,t}=\Gamma_{(i,t)^{-}}$. If for the remaining time of the era $i$ the path $\gamma$ stays in $\Gamma_{i,t}$ then we set $\tau=T_{i}+1$. If we have $\tau\neq T_{i}+1$ then $(i,\tau)$ is the next important, merging or first time. If $(i,t)$ is a merging time we also mark it $(\bar{y}_{i,t},\vec{x}_{i,t+1},\vec{x}_{i,\tau})$ where $(i,\tau)$ is once again the next important, merging or first time. For these two marked times, in the time in between $(i,t)$ and $(i,\tau)$ we only go through seen vertices of $K_{\gamma}$ and seen vertices of $\Gamma_{i,t}$ and therefore one can recover $\{\bar{\gamma}_{i,s}:\leavevmode\nobreak\ t\leq s\leq\tau\}$. Finally for $i\in[2m]$ odd, we call first outgoing time $(i,t)$ the first time occurring that is either a first or important time and mark it with $(\vec{x}_{1,1},\bar{y}_{i,t},\vec{x}_{i,t+1},\vec{x}_{i,\tau})$ where $(i,\tau)$ is the next important time. If a first outgoing time is also a first time, in the time in between $(i,t)$ and $(i,\tau)$ we only see first times. If for $t\leq s\leq T_{i}$ we have only first times we set $\tau=T_{i}+1$. Therefore knowing the first outgoing, important and merging times and their marks we can recover all $\bar{\gamma}$.
Remark 7.1.
We can describe the recovering of $\bar{\gamma}$ as follows. One has to segment the recovering with each eras $i$. If the time $(1,1)$ is not marked then it a first time, therefore we can set that $\bar{x}_{1,1}=1$, $\bar{c}(1)=\bar{c}(\bar{x}_{1,1})=1$, $\bar{y}_{1,1}=1$, $\bar{x}_{1,2}=2$ and finally $\bar{c}(2)=\bar{c}(\bar{x}_{1,2})=2$. Therefore we know that $\bar{\gamma}_{1,1}=((1,1),1,(2,2))$. If now $(1,1)$ is an important time, we still have $\bar{x}_{1,1}=1$, $\bar{c}(1)=1$ and $\bar{y}_{1,1}=1$ but then we have $\bar{c}(2)=1$ and we know the value of $\bar{x}_{1,2}$ from the mark on $(1,1)$. Since $\Gamma_{1,1}=\Gamma_{1,0}$ has no edge, necessarily the next time we step outside the tree is the next time, i.e the time $(1,2)$ is the next first or important time. We start by explaining, for an era $j=2i$ how to recover $\bar{\gamma}_{j,s}$ for $1\leq s\leq t$ where $(j,t)$ is the first marked time but not first outgoing time of the era $j$ considering that $\bar{\gamma}_{j^{\prime},s^{\prime}}$ for $(1,1)\prec(j^{\prime},s^{\prime})\prec(j,1)$ are known. Let $(2i,t)$ be the first marked time and let $(\bar{y}_{2i,t},\vec{x}_{2i,t+1},\vec{x}_{2i,\tau})$ be its mark. If $\bar{y}_{2i,t}$ (or $\bar{c}(\bar{x}_{2i,t+1})$) has never been seen during the preceding eras then it means $(2i,t)$ is an important time and in the time in between $(2i,1)$ and $(2i,t)^{-}$ we only have first times. We can recover then $\bar{\gamma}_{2i,s}$ for $1\leq s\leq t$. If now $\bar{y}_{2i,t}$ or $\bar{c}(\bar{x}_{2i,t+1})$ has been seen before, then $(2i,t)$ is the merging time and once again we know that we only have first times between $(2i,1)$ and $(2i,t)^{-}$. Now we consider an era $j=2i+1$, $(2i+1,t)$ its first outgoing time and $(\vec{x}_{2i+1,1},\bar{y}_{2i+1,t},\vec{x}_{2i+1,t},\vec{x}_{2i+1,\tau})$ its mark. Then there exists only one path between $\operatorname{cc}(x_{2i+1,1})$ and $\operatorname{cc}(x_{2i+1,\tau})$ and even though $\tau$ is not indexed in the mark it is given by the length of this path. Now if there exists another marked time $(2i+1,t^{\prime})$ in the same era $12i+$ 1, it is necessarily an important time. If we assume $\tau\neq t^{\prime}$, we have $\bar{\gamma}_{2i+1,s}$ for $1\leq s\leq\tau$ given by the only path between $\operatorname{cc}(x_{2i+1,1})$ and $\operatorname{cc}(x_{2i+1,\tau})$. We then recover $\bar{\gamma}_{2i+1,s}$ for $\tau+1\leq s\leq t-1$ from the fact that $\{(2i+1,s):\leavevmode\nobreak\ \tau+1\leq s\leq t^{\prime}-1\}$ are first times. If $\tau=t^{\prime}$ then $\bar{\gamma}_{2i+1,s}$ for $1\leq s\leq\tau=t^{\prime}$ is given by the only path between $\operatorname{cc}(x_{2i+1,1})$ and $\operatorname{cc}(x_{2i+1,t^{\prime}})$. If there is no other marked time in the era $2i$ then we only have first times (and $\tau=T_{2i+1}+1$). We now consider $(i,t)\prec(i,t^{\prime})$ two consecutive marked times of a same era $i$ and we suppose $(\bar{\gamma}_{j,s})_{(j,s)\prec(i,t)}$ is known. Let $(\bar{y}_{i,t},\vec{x}_{i,t+1},\vec{x}_{i,\tau})$ be the mark on $(i,t)$ and $(\bar{y}_{i,t^{\prime}},\vec{x}_{i,t^{\prime}+1},\vec{x}_{i,\tau^{\prime}})$ the one on $(i,t^{\prime})$. We have only one path (possibly empty) possible between $\operatorname{cc}(x_{i,t+1})$ and $\operatorname{cc}(x_{i,\tau})$. The length of this path and the path itself gives us the value of $\tau$ and $\bar{\gamma}_{i,s}$ for $t\leq s\leq\tau$. If $\tau=t^{\prime}$ then we have recovered $\bar{\gamma}_{i,s}$ for $t\leq s\leq t^{\prime}$. Otherwise all $(i,s)$ for $\tau\leq s\leq t^{\prime}$ are necessarily first times.
With the previous encoding, for each important time or merging time we have at most $a(s-p)^{2}s^{2}$ ways to mark it. We have at most $m$ merging times. Finally we have at most $m$ first outgoing time that have each $a(s-p)^{3}s^{3}$ ways to be marked. We have at most $2m\ell$ ways to position the important times. However with no further details on our path, upper bounding roughly the number of important times $|\boldsymbol{T}_{\mathrm{imp}}|$ by $2m\ell-s+p$ is too big to be implemented in (30). Therefore we are going to take into account the tangle-freeness of our paths $\gamma_{i}$ to make some important times superfluous. Meaning that there will be more specific important times that -with the good modified encoding, will allow us to not take into account the other ones.
To proceed we consider each era $i$ separately. We use the fact that $\gamma_{i}$ being tangle-free, if we explore a cycle during the era $i$ we know it to be the only one. For $i\in[2m]$ let $(i,t_{0})$ be the smallest time such that $\operatorname{cc}(x_{i,t_{0}+1})\in\{\operatorname{cc}(x_{i,1}),\operatorname{cc}(x_{i,2}),...,\operatorname{cc}(x_{i,t_{0}})\}$, if it exists. Let $1\leq\sigma\leq t_{0}$ be such that $\operatorname{cc}(x_{i,t_{0}+1})=\operatorname{cc}(x_{i,\sigma})$. We denote by $\mathrm{C}_{i}:=(\bar{\operatorname{cc}}(x_{i,\sigma}),\bar{\operatorname{cc}}(x_{i,\sigma+1}),...,\bar{\operatorname{cc}}(x_{i,t_{0}+1}))$ and $\mathrm{C}_{i}$ is the only cycle visited by $\gamma_{i}$. We define the short cycling time as the greatest important time $(i,t)$ smaller then $(i,t_{0})$ in $(\boldsymbol{T},\prec)$ and $(i,\hat{t})$ the smallest time $(i,\hat{t})\succeq(i,\sigma)$ verifying that $\operatorname{cc}(x_{i,\hat{t}+1})\notin\mathrm{C}_{i}$. As convention we take $\hat{t}=T_{i}+1$ if $\gamma_{i}$ stays in $\mathrm{C}_{i}$ for the remaining time of the era $i$. If $\hat{t}\geq t_{0}+2$ then the cycle $\mathrm{C}_{i}$ is visited more then once, even though it may not finish the cycle after the first visit. We replace the mark on $(i,t)$ with $(\bar{y}_{i,t},\vec{x}_{i,t+1},\sigma,\hat{t},\vec{x}_{i,\tau})$ where now $(i,\tau)$ is the next important time after $(i,\hat{t})$. We take as convention that $\tau=T_{i}+1$ if $\gamma_{i}$ stays in $\Gamma_{(i,t)}$ for the remaining time of the era $i$. During the era $i$ we have only one short cycling time. The important times $(i,t^{\prime})$ that verify $1\leq t^{\prime}\leq t$ or $\tau\leq t^{\prime}\leq T_{i}$ are called long cycling time. The important times in the time in between $(i,t)$ and $(i,\tau)$ are called superfluous and are not marked anymore. We can now prove 7.2.
Proof of 7.2.
The previous encoding also allows us to recover $\bar{\gamma}$ from the marks on the long cycling and short cycling times (see proof of [BQZ18, Lemma 3]). If an excess edge is seen twice in the era $i$ outside of the time interval $(i,t)^{+}$ and $(i,\tau)$ it would mean that we discover another cycle in $\gamma_{i}$, which is impossible. Also the excess edge seen at the short cycling time $(i,t)$ cannot be seen by a long cycling time either, for the same reason. We have the following for each kind of marked times. We have at most $g$ long cycling times for each era $i\in[2m]$. For each long cycling time we have at most $a(s-p)^{2}s^{2}$ ways to mark it. We also have possibly $m$ merging times, and again at most $a(s-p)^{2}s^{2}$ to mark one of them. Similarly we have at most $m$ first outgoing time. For each first outgoing time we have at most $a(s-p)^{3}s^{3}$ ways to mark it. Finally at each era we have at most one short cycling time with at most $a(s-p)^{2}s^{2}\ell^{2}$ ways to mark it. In each era, for any short cycling, long cycling or first outgoing time, we have at most $T_{i}\leq\ell$ ways of positioning the time in the era. Therefore we have
$|\mathcal{W}_{\ell,\vec{T}}(s,a,p)|$ $\leq[\ell a(s-p)^{2}s^{2}]^{2mg}[\ell a(s-p)^{2}s^{2}\ell^{2}]^{2m}[\ell a(s-p)^{2}s^{2}]^{m}[\ell a(s-p)^{3}s^{3}]^{m}$
$\leq[\ell a(s-p)^{2}s^{2}]^{2m(g+5)}.$
7.3 Contributions of the matrix $Q$ for a fixed path and proof of 7.1
In this section we fix $[\gamma]\in\mathcal{W}_{\ell,\vec{T}}(s,a,p)$ and we denote by $a_{1}$ the number of inconsistent edges $e\in\mathrm{A}_{\gamma}$ of multiplicity $1$ (see 6.5).
Lemma 7.5.
Let $m\geq 1$ and $1\leq\ell\leq\frac{\ln(N)}{6\ln(d)}$ be fixed integers. There exists $c,c(n)\geq 1$ such that
$\sum_{\gamma^{\prime}\sim\gamma}q_{\#}(\gamma^{\prime})\leq c(n)N^{s-p}(cm\ell)^{a_{1}+3(m+g)+p}\rho_{n,\#}^{2m(\ell+1)}.$
To prove the previous lemma we need the following definition and lemmas.
Definition 7.6.
Let $\mathcal{G}=(\mathcal{V},\mathcal{E})$ be a directed graph with $\mathcal{E}$ being a multi-set, and connected if we do not take into account the orientation (i.e. weakly connected). We also consider that $\mathcal{G}$ is bi-partite in the following sense. There exists $\mathcal{X}\bigsqcup\mathcal{Y}$ partition of $\mathcal{V}$ such that for all $e=(u,v)\in\mathcal{E}$, we have $u\in\mathcal{X}$ and $v\in\mathcal{Y}$. Let $\mathcal{T}\subset\mathcal{G}$ be (any) underlying covering tree of $\mathcal{G}$ and $\mathcal{R}=\mathcal{G}\backslash\mathcal{T}$.
  • We say that $(e,e^{\prime})\in\mathcal{E}^{2}$ are a double edge if both $e$ and $e^{\prime}$ goes from $u\in\mathcal{X}$ to $v\in\mathcal{Y}$ and $e\in\mathcal{T}$.
  • For $e,e^{\prime}\in\mathcal{R}$ we denote $e\sim e^{\prime}$ when both $e$ and $e^{\prime}$ goes from $u\in\mathcal{X}$ to $v\in\mathcal{Y}$.
Lemma 7.7.
Let $\mathcal{G}=(\mathcal{X}\sqcup\mathcal{Y},\mathcal{E})$ be a bi-partite connected and directed graph in the sense of 7.6 and $\mathcal{T}$ an underlying tree. If there is a double edge in $\mathcal{G}$ we denote it $\𝒯(e,e^{\prime})\in\mathcal{T}\times(\mathcal{E}\backslash\mathcal{T})$. Let $(\ell(f))_{f\in\mathcal{E}}$ be a family of integers. We define
$Z(\mathcal{G}):=\sum_{\begin{subarray}{c}y_{u}\in[N]\\ u\in\mathcal{\mathcal{Y}}\end{subarray}}\sum_{\begin{subarray}{c}x_{v}\in[N]\\ v\in\mathcal{\mathcal{X}}\end{subarray}}\prod_{\begin{subarray}{c}f\in\mathcal{E}\\ f=(u,v)\end{subarray}}(Q^{\ell(f)})_{y_{u}x_{v}}\leavevmode\nobreak\ \leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ \leavevmode\nobreak\ L_{\mathcal{T}}:=\sum_{f\in\mathcal{T}\backslash\{e\}}\ell(f),$
where $\{e\}=\emptyset$ if there is no double edge. If we have $\mathcal{G}=\mathcal{T}\sqcup\{e^{\prime}\}$ and $(e,e^{\prime})$ a double edge, then we have
$Z(\mathcal{G})\leq N\operatorname{Tr}_{N}\left((Q^{\ell(e)}{Q^{\ast}}^{\ell(e^{\prime})}\right)\delta^{L_{\mathcal{T}}}.$
If $\mathcal{G}=\mathcal{T}$ we have
$Z(\mathcal{T})\leq N\delta^{L_{\mathcal{T}}}.$
Proof of 7.7.
To ease notations we denote by $\mathcal{T}$ the underlying tree but also the set of edges of this same underlying tree. We proceed by induction on $|\mathcal{T}|=|\mathcal{X}|+|\mathcal{Y}|-1$, where the equality is due to the the fact that $\mathcal{T}$ is a covering tree on a set of cardinal $|\mathcal{X}|+|\mathcal{Y}|$. In the case where $|\mathcal{T}|=1$ and $\mathcal{G}$ has a double edge we have
$Z(\mathcal{G})$ $=\sum_{x,y}(Q^{\ell}(e))_{yx}(Q^{\ell(e^{\prime})})_{yx}=N\operatorname{Tr}_{N}\left(Q^{\ell(e)}{Q^{\ast}}^{\ell(e^{\prime})}\right).$
In the case where there is no double edge and $|\mathcal{T}|=1$, then necessarily we have only one edge and therefore
$Z(\mathcal{G})=\sum_{x,y}(Q^{\ell(f)})_{yx}\leq N\delta^{\ell(f)}.$
We now assume $|\mathcal{T}|\geq 2$. We consider $\{e,e^{\prime}\}$ the possibly empty set of double edge and $e_{0}=(u_{0},v_{0})$ a leaf of $\mathcal{T}$ that is not $e$. Such a leaf exists since for any tree with more than two edges there exists at least two different leaves. Since we are considering an oriented graph that is weakly connected, denoting $\mathrm{d}_{\mathcal{T}}^{\mathrm{o}}(\cdot)$ the weak degree in $\mathcal{T}$, one needs to distinguish whether we have $\mathrm{d}_{\mathcal{T}}^{\mathrm{o}}(u_{0})=1$ or $\mathrm{d}_{\mathcal{T}}^{\mathrm{o}}(v_{0})=1$. We consider the first case and we have
$Z(\mathcal{G})$ $=\sum_{\begin{subarray}{c}(y_{u},x_{v})\in[N]^{2}\\ (u,v)\in\mathcal{Y}\backslash\{u_{0}\}\times\mathcal{X}\end{subarray}}\prod_{f\in\mathcal{T}\backslash\{e_{0}\}}(Q^{\ell(f)})_{y_{u}x_{v}}\sum_{y_{u_{0}}\in[N]}(Q^{\ell(e_{0})})_{y_{u_{0}}x_{v_{0}}}.$
We then bound the sum over $y_{u_{0}}\in[N]$ by $\delta^{\ell(e_{0})}$ and apply the induction to $\mathcal{G}^{\prime}=(\mathcal{X}\sqcup(\mathcal{Y}\backslash\{u_{0}\}),\mathcal{E}\backslash\{e_{0}\})$. In the case where we have $\mathrm{d}^{\mathrm{o}}_{\mathcal{T}}(v_{0})=1$ we similarly obtain
$Z(\mathcal{G})$ $=\sum_{\begin{subarray}{c}(y_{u},x_{v})\in[N]^{2}\\ (u,v)\in\mathcal{Y}\times\mathcal{X}\backslash\{v_{0}\}\end{subarray}}\prod_{f\in\mathcal{T}\backslash\{e_{0}\}}(Q^{\ell(f)})_{y_{u}x_{v}}\sum_{x_{v_{0}}\in[N]}(Q^{\ell(e_{0})})_{y_{u_{0}}x_{v_{0}}}.$
We then bound the sum over $x_{v_{0}}\in[N]$ by $\delta^{\ell(e_{0})}$ and apply the induction to $\mathcal{G}^{\prime}=((\mathcal{X}\backslash\{v_{0}\})\sqcup\mathcal{Y},\mathcal{E}\backslash\{e_{0}\})$. ∎
For $\gamma\in\mathrm{W}_{\ell,\vec{T}}(s,a,p)$, we rewrite $q(\gamma)$ as follows
$q(\gamma)=\prod_{i=1}^{2m}\prod_{t=1}^{T_{i}}(Q^{\ell_{i,t}})_{y_{i,t},x_{i,t+1}}=\prod_{\operatorname{cc}\in\chi}\prod_{\begin{subarray}{c}f\in\mathcal{E}_{\operatorname{cc}}\\ f=(u,v)\end{subarray}}(Q^{\ell(f)})_{x_{u},x_{v}}$
where $\chi:=X_{\gamma}/\sim$ are the connected components of $K_{\gamma}$ and for $\operatorname{cc}\in\chi$ we considered the graph $\mathcal{G}_{\operatorname{cc}}:=(\mathcal{V}_{\operatorname{cc}},\mathcal{E}_{\operatorname{cc}})$ defined by $\mathcal{V}_{\operatorname{cc}}:=\{\bar{x};\leavevmode\nobreak\ x\in\operatorname{cc}\}\sqcup\{\bar{y};\leavevmode\nobreak\ \exists(i,t)\in\boldsymbol{T},\leavevmode\nobreak\ x\in\operatorname{cc};\leavevmode\nobreak\ (y_{i,t},x_{i,t+1})=(y,x)\}=:\mathcal{X}_{\operatorname{cc}}\sqcup\mathcal{Y}_{\operatorname{cc}}$ and $\mathcal{E}_{\operatorname{cc}}$ is the multi-set defined by $\mathcal{E}_{\operatorname{cc}}:=\{(u,v)\in\mathcal{V}_{\operatorname{cc}}^{2};\leavevmode\nobreak\ \leavevmode\nobreak\ \exists(i,t)\in\boldsymbol{T};\leavevmode\nobreak\ \leavevmode\nobreak\ (\bar{y}_{i,t},\bar{x}_{i,t+1})=(u,v)\}$. The notations are the one introduced in the proof of 7.2 and the graphs $(\mathcal{G}_{\operatorname{cc}})_{\operatorname{cc}\in\chi}$ are directed, weakly connected and bi-partite in the sense of 7.6. In the writing of $q(\gamma)$ above we have considered that for $u=\bar{x}$ we set $x_{u}=x$. We consider $\mathcal{T}_{\operatorname{cc}}$ the covering tree obtained by constructing the tree while we explore $\gamma$ chronologically and we set $\mathcal{R}_{\operatorname{cc}}:=\mathcal{G}_{\operatorname{cc}}\backslash\mathcal{T}_{\operatorname{cc}}$. We introduce $\chi_{\mathrm{d}}\subset\chi$ the set of connected component $\operatorname{cc}\in\chi$ such that there exists $(e(\mathrm{cc}),e^{\prime}(\mathrm{cc}))$ a double edge in $\mathcal{G}_{\operatorname{cc}}$. We consider the graphs
$\mathcal{G}=(\mathcal{X}\sqcup\mathcal{Y},\mathcal{E}):=\bigsqcup_{\operatorname{cc}\in\chi}\mathcal{G}_{\operatorname{cc}},\leavevmode\nobreak\ \leavevmode\nobreak\ \mathcal{F}=\bigsqcup_{\operatorname{cc}\in\chi}\mathcal{T}_{\operatorname{cc}}\backslash\{e(\operatorname{cc})\}\leavevmode\nobreak\ \leavevmode\nobreak\ \mathcal{R}:=\bigsqcup_{\operatorname{cc}\in\chi}\mathcal{R}_{\operatorname{cc}}\backslash\{e^{\prime}(\operatorname{cc})\},\leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ \mathcal{D}:=\bigsqcup_{\operatorname{cc}\in\chi}\{e(\operatorname{cc}),e^{\prime}(\operatorname{cc})\}$
where for $\operatorname{cc}\notin\chi_{\mathrm{d}}$ we have $\{e(\operatorname{cc})\}=\{e^{\prime}(\operatorname{cc})\}=\{\emptyset\}$. We set
$F:=|\mathcal{F}|\leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ R:=|\mathcal{R}|.$ (31)
In order to prove 7.9, we provide an upper bound for $F$.
Lemma 7.8.
Let $m,\ell\geq 1$, $1\leq s,a,p\leq 2m\ell$ and $\vec{T}\in[\ell]^{2m}$ be fixed. For all $\gamma\in W_{\ell,\vec{T}}(s,a,p)$, considering the corresponding $F$ given by (31) we have
$F\leq 3(m+g)-2+a_{1}+p.$
We use the previous lemmas to prove 7.5.
Proof of 7.5.
For a finite set $\mathcal{S}$ and an integer $I$ we introduce
$\mathcal{L}^{\prime}(\mathcal{S},I):=\left\{(\ell(s))_{s\in\mathcal{S}}\in[\ell]^{\mathcal{S}}:\leavevmode\nobreak\ \sum_{s\in\mathcal{S}}\ell(s)=I\right\}.$
We give the detail of the upper bound for $q$ defined by (29) and we precise along the proof where the reasoning needs to be adapted for $q_{r}$. We bound our sum as follows
$\sum_{\gamma^{\prime}\sim\gamma}q(\gamma^{\prime})$ $=\sum_{\gamma^{\prime}\sim\gamma}\prod_{\operatorname{cc}\in\chi}\prod_{\begin{subarray}{c}f\in\mathcal{E}_{\operatorname{cc}}\\ f=(u,v)\end{subarray}}(Q^{\ell(f)})_{x^{\prime}_{u},x^{\prime}_{v}}\leq\sum_{\begin{subarray}{c}y_{u},x_{v}\in[N]\\ (u,v)\in\mathcal{Y}\times\mathcal{X}\end{subarray}}\sum_{(\ell(f))_{f\in\mathcal{E}}\in\mathcal{L}^{\prime}(\mathcal{E},M)}\prod_{\operatorname{cc}\in\chi}\prod_{f\in\mathcal{E}_{\operatorname{cc}}}(Q^{\ell(f)})_{y_{u}x_{v}}$ (32)
where we set $M=2m(\ell+1)-T$ for
$T:=\sum_{i=1}^{2m}T_{i}.$
For a fixed choice of $\{(y_{u},x_{v}):\leavevmode\nobreak\ (u,v)\in\mathcal{Y}\times\mathcal{X}\}$, we rewrite and bound the second sum above as follows
$\sum_{\ast}\prod_{\operatorname{cc}\in\chi}\prod_{f\in\mathcal{E}_{\operatorname{cc}}}(Q^{\ell(f)})_{y_{u}x_{v}}$ $=\sum_{L=0}^{M}\sum_{\begin{subarray}{c}\ast\ast\end{subarray}}\prod_{\operatorname{cc}\in\chi}\prod_{f\notin(\mathcal{R}_{\operatorname{cc}}\backslash\{e^{\prime}(\operatorname{cc})\})}(Q^{\ell(f)})_{y_{u}x_{v}}\sum_{\ast\ast\ast}\prod_{\operatorname{cc}\in\chi}\prod_{f\in(\mathcal{R}_{\operatorname{cc}}\backslash\{e^{\prime}(\operatorname{cc})\})}(Q^{\ell(f)})_{y_{u}x_{v}}$
[…]
$\sum_{\ast\ast\ast}\prod_{\operatorname{cc}\in\chi}\prod_{f\in(\mathcal{R}_{\operatorname{cc}}\backslash\{e^{\prime}(\operatorname{cc})\})}(Q^{\ell(f)})_{y_{u}x_{v}}\leq\binom{M-L+R-1}{R-1}\kappa^{M-L+R}\leq(2\kappa)^{M-L+R-1}\kappa.$ (33)
Implementing the previous bound in (32)
$\sum_{\gamma^{\prime}\sim\gamma}q(\gamma^{\prime})$ $\leq\frac{1}{2}\sum_{L=0}^{M}(2\kappa)^{M-L+R}\sum_{\ast\ast}\prod_{\operatorname{cc}\in\chi}\sum_{\begin{subarray}{c}y_{u},x_{v}\in[N]\\ (u,v)\in\mathcal{Y}_{\operatorname{cc}}\times\mathcal{X}_{\operatorname{cc}}\end{subarray}}\prod_{f\in\{e^{\prime}\}\sqcup\mathcal{T}_{\operatorname{cc}}}(Q^{\ell(f)})_{y_{u}x_{v}}$
$\leq\frac{1}{2}\sum_{L=0}^{M}(2\kappa)^{M-L+R}\sum_{\ast\ast}N^{s-p}\delta^{L_{\mathcal{F}}}\prod_{\operatorname{cc}\in\chi_{\mathrm{d}}}\operatorname{Tr}_{N}\left(Q^{\ell(e(\operatorname{cc}))}{Q^{\ast}}^{\ell(e^{\prime}(\operatorname{cc}))}\right)$
where we used the fact that all $\left(\mathcal{G}_{\operatorname{cc}}\right)_{\operatorname{cc}\in\chi}$ are disjoint graphs in the first line and we applied 7.7 to each $\operatorname{cc}\in\chi$ in the second line, setting $L_{\mathcal{F}}:=\sum_{\operatorname{cc}}L_{\mathcal{T}_{\operatorname{cc}}}$. Applying the same reasoning replacing $(Q^{\ell(f)})_{f}$ with ($(\{rQ\}^{\ell(f)})_{f}$ we obtain
$\sum_{\gamma^{\prime}\sim\gamma}q_{r}(\gamma^{\prime})$ $\leq\frac{1}{2}\sum_{L=0}^{M}(2r\kappa)^{M-L+R}\sum_{\ast\ast}N^{s-p}(r\delta)^{L_{\mathcal{F}}}(1-r)^{T}\prod_{\operatorname{cc}\in\chi_{\mathrm{d}}}\operatorname{Tr}_{N}\left(\{rQ\}^{\ell(e(\operatorname{cc}))}\{rQ^{\ast}\}^{\ell(e^{\prime}(\operatorname{cc}))}\right).$
We now sum over all $(\ell_{cc},\ell_{\operatorname{cc}}^{\prime})_{\operatorname{cc}\in\chi_{d}}$ possible and obtain
$\sum_{\gamma^{\prime}\sim\gamma}q(\gamma^{\prime})$ $\leq\frac{1}{2}N^{s-p}\sum_{L=0}^{M}(2\kappa)^{M-L+R}\sum_{L_{\mathrm{d}}=0}^{L}\delta^{L-L_{\mathrm{d}}}|\mathcal{L}^{\prime}(\mathcal{F},L-L_{\mathrm{d}})|\sum_{\begin{subarray}{c}(\ell_{\operatorname{cc}},\ell_{\operatorname{cc}}^{\prime})_{\operatorname{cc}\in\chi_{\mathrm{d}}}\\ \sum_{\operatorname{cc}}\ell_{\operatorname{cc}}+\ell_{cc}^{\prime}=L_{\mathrm{d}}\end{subarray}}\prod_{\operatorname{cc}\in\chi_{\mathrm{d}}}\operatorname{Tr}_{N}\left(Q^{\ell_{\operatorname{cc}}}{Q^{\ast}}^{\ell_{\operatorname{cc}}^{\prime}}\right)$
$\leq\frac{1}{2}N^{s-p}(2m(\ell+1))^{F}\sum_{L=0}^{M}\sum_{L_{\mathrm{d}}=0}^{L}(2\kappa)^{M-L+R}\delta^{L-L_{\mathrm{d}}}\sum_{\begin{subarray}{c}(\ell_{\operatorname{cc}},\ell_{\operatorname{cc}}^{\prime})_{\operatorname{cc}\in\chi_{\mathrm{d}}}\\ \sum_{\operatorname{cc}}\ell_{\operatorname{cc}}+\ell_{cc}^{\prime}=L_{\mathrm{d}}\end{subarray}}\prod_{\operatorname{cc}\in\chi_{\mathrm{d}}}\operatorname{Tr}_{N}\left(Q^{\ell_{\operatorname{cc}}}{Q^{\ast}}^{\ell_{\operatorname{cc}}^{\prime}}\right).$
For all $0\leq L_{\mathrm{d}}^{1}+L_{\mathrm{d}}^{2}=L_{\mathrm{d}}$, denoting $\mathrm{x}_{\mathrm{d}}=|\chi_{\mathrm{d}}|$ we have
$\sum_{\begin{subarray}{c}(\ell_{\operatorname{cc}})_{\operatorname{cc}}\in\mathcal{L}^{\prime}(\chi_{\mathrm{d}},L_{\mathrm{d}}^{1})\\ (\ell^{\prime}_{\operatorname{cc}})_{\operatorname{cc}}\in\mathcal{L}^{\prime}(\chi_{\mathrm{d}},L_{\mathrm{d}}^{2})\end{subarray}}\prod_{\operatorname{cc}\in\chi_{\mathrm{d}}}\operatorname{Tr}_{N}\left(Q^{\ell_{\operatorname{cc}}}{Q^{\ast}}^{\ell_{\operatorname{cc}}^{\prime}}\right)$ $\leq\sum_{\begin{subarray}{c}(\ell_{\operatorname{cc}})_{\operatorname{cc}}\in\mathcal{L}(\chi_{\mathrm{d}},L_{\mathrm{d}}^{1}+\mathrm{x}_{\mathrm{d}})\\ (\ell^{\prime}_{\operatorname{cc}})_{\operatorname{cc}}\in\mathcal{L}(\chi_{\mathrm{d}},L_{\mathrm{d}}^{2}+\mathrm{x}_{\mathrm{d}})\end{subarray}}\prod_{\operatorname{cc}\in\chi_{\mathrm{d}}}\operatorname{Tr}_{N}\left(Q^{\ell_{\operatorname{cc}}}{Q^{\ast}}^{\ell_{\operatorname{cc}}^{\prime}}\right)$
$\leq\sum_{T=0}^{L_{\mathrm{d}}^{1}\land L_{\mathrm{d}}^{2}+\mathrm{x}_{\mathrm{d}}}\sum_{\begin{subarray}{c}(\ell_{t})_{t}\in\mathcal{L}(T,L_{\mathrm{d}}^{1}+\mathrm{x}_{\mathrm{d}})\\ (\ell^{\prime}_{t})_{\operatorname{cc}}\in\mathcal{L}(T,L_{\mathrm{d}}^{2}+\mathrm{x}_{\mathrm{d}})\end{subarray}}\prod_{t=1}^{T}\operatorname{Tr}_{N}\left(Q^{\ell_{t}}{Q^{\ast}}^{\ell_{t}^{\prime}}\right)$
$\leq\Phi(a^{L_{\mathrm{d}}^{1}+\mathrm{x}_{\mathrm{d}}}{a^{\ast}}^{L_{\mathrm{d}}^{2}+\mathrm{x}_{\mathrm{d}}})\leq\mathrm{N}_{L_{\mathrm{d}}^{1}+\mathrm{x}_{\mathrm{d}}}^{L_{\mathrm{d}}^{1}+\mathrm{x}_{\mathrm{d}}}\mathrm{N}_{L_{\mathrm{d}+\mathrm{x}_{\mathrm{d}}}^{2}}^{L_{\mathrm{d}}^{2}+\mathrm{x}_{\mathrm{d}}}$
where for the last two inequalities we applied 4.5 and we have used that for $b,h\in\mathcal{A}$
$\Phi(bh^{*})\leq\sqrt{\tau(bb^{*})\tau(hh^{*})},\leavevmode\nobreak\ \leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ \leavevmode\nobreak\ \tau(b^{\ast}b)=\mathbb{E}[\|b^{\ast}b(e_{0})\|_{2}^{2}]\leq\|b\|^{2}.$
The integer $n$ being fixed, for all integer $L=\mathrm{q}n+\mathrm{r}\geq 1$ with $10\leq\mathrm{r}\leq n-$ 1, We set $c(n):=(\kappa\vee 1)(1+K)^{n}$ and we have:
$\mathrm{N}_{L}^{L}=\|(u+q)^{L}\|\leq\|(u+q)^{n}\|^{\mathrm{q}}\|(u+q)\|^{\mathrm{r}}\leq\mathrm{N}_{n}^{L-\mathrm{r}}(1+\|Q\|)^{n}\leq c(n)\mathrm{N}_{n}^{L}$
where we used that the operator norm is sub-multiplicative. For the convex combination case we have
$(1-r)^{2(\mathrm{x}_{\mathrm{d}}-1)}\sum_{(\ell_{\operatorname{cc}},\ell^{\prime}_{\operatorname{cc}})_{\operatorname{cc}}}\prod_{\operatorname{cc}\in\chi_{\mathrm{d}}}\operatorname{Tr}_{N}\left(\{rQ\}^{\ell_{\operatorname{cc}}}\{rQ^{\ast}\}^{\ell_{\operatorname{cc}}^{\prime}}\right)\leq\sum_{\ast\ast\ast\ast}(1-r)^{2(\mathrm{x}_{\mathrm{d}}-1)}\prod_{\operatorname{cc}\in\chi_{\mathrm{d}}}\operatorname{Tr}_{N}\left(\{rQ\}^{\ell_{\operatorname{cc}}}\{rQ^{\ast}\}^{\ell_{\operatorname{cc}}^{\prime}}\right)$
$\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leq\sum_{T=0}^{L_{\mathrm{d}}^{1}\land L_{\mathrm{d}}^{2}+\mathrm{x}_{\mathrm{d}}}\sum_{\begin{subarray}{c}(\ell_{t})_{t}\in\mathcal{L}(T,L_{\mathrm{d}}^{1}+\mathrm{x}_{\mathrm{d}})\\ (\ell^{\prime}_{t})_{\operatorname{cc}}\in\mathcal{L}(T,L_{\mathrm{d}}^{2}+\mathrm{x}_{\mathrm{d}})\end{subarray}}(1-r)^{2(T-1)}\prod_{t=1}^{T}\operatorname{Tr}_{N}\left(\{rQ\}^{\ell_{t}}\{rQ^{\ast}\}^{\ell_{t}^{\prime}}\right)$
$\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leq\Phi(p^{L_{\mathrm{d}}^{1}+\mathrm{x}_{\mathrm{d}}}{p^{\ast}}^{L_{\mathrm{d}}^{2}+\mathrm{x}_{\mathrm{d}}})\leq\mathrm{N}_{r,L_{\mathrm{d}}^{1}+\mathrm{x}_{\mathrm{d}}}^{L_{\mathrm{d}}^{1}+\mathrm{x}_{\mathrm{d}}}\mathrm{N}_{r,L_{\mathrm{d}+\mathrm{x}_{\mathrm{d}}}^{2}}^{L_{\mathrm{d}}^{2}+\mathrm{x}_{\mathrm{d}}}.$
We obtain
$\sum_{\gamma^{\prime}\sim\gamma}q(\gamma^{\prime})$ $\leq N^{s-p}(2m(\ell+1))^{F}2m(\ell+1)c(n)^{2}\sum_{L=0}^{M}\sum_{L_{\mathrm{d}}=0}^{L}(2\kappa)^{M-L+R}\delta^{L-L_{\mathrm{d}}}\mathrm{N}_{n}^{L_{\mathrm{d}}+\mathrm{x}_{\mathrm{d}}}$
$\leq N^{s-p}(2m(\ell+1))^{F}(2m(\ell+1))^{2}c(n)^{2}\rho_{n}^{2m(\ell+1)}$
where we used the fact that owing to 4.7 we have $(2\kappa)\vee\delta\vee\mathrm{N}_{n}\geq 1$ and the fact that $M+R+\mathrm{x}_{\mathrm{d}}\leq 2m(\ell+1)$. It remains to apply 7.8 and adapt the constant $c(n)>0$. ∎
We now give the proof of 7.8.
Proof of 7.8.
We have
$F:=\sum_{\operatorname{cc}\in\chi_{\mathrm{d}}}(|\mathcal{X}_{\operatorname{cc}}|+|\mathcal{Y}_{\operatorname{cc}}|-2)+\sum_{\operatorname{cc}\notin\chi_{\mathrm{d}}}(|\mathcal{X}_{\operatorname{cc}}|+|\mathcal{Y}_{\operatorname{cc}}|-1).$
We introduce $\chi_{\mathrm{d}}^{\prime}$ the set of $\operatorname{cc}\in\chi_{\mathrm{d}}$ such that the connected component verifies $|\mathcal{X}_{\operatorname{cc}}|=1$ and there is only one incoming edge $e\in\mathrm{A}_{\gamma}$ of multiplicity $m(e)\geq 2$. For all $\operatorname{cc}\in\chi_{\mathrm{d}}^{\prime}$, we have $\mathcal{T}_{\operatorname{cc}}\backslash\{e(\operatorname{cc})\}=\emptyset$, therefore:
$F\leq\sum_{\operatorname{cc}\notin\chi_{\mathrm{d}}^{\prime}}(|\mathcal{X}_{\operatorname{cc}}|+|\mathcal{Y}_{\operatorname{cc}}|-1)\leq p+\sum_{\operatorname{cc}\notin\chi_{\mathrm{d}}^{\prime}}|\mathcal{Y}_{\operatorname{cc}}|$
where we used that $\sum_{\operatorname{cc}\in\chi}(|\mathcal{X}_{\operatorname{cc}}|-1)=p$. Now we re-write $Y:=\sum_{\operatorname{cc}\notin\chi_{\mathrm{d}}^{\prime}}|\mathcal{Y}_{\operatorname{cc}}|$ noticing that we have $\mathcal{Y}:=\bigcup_{\operatorname{cc}\notin\chi_{\mathrm{d}}^{\prime}}\mathcal{Y}_{\operatorname{cc}}=\mathcal{Y}_{1}\sqcup\mathcal{Y}_{2}\sqcup\mathcal{Y}_{3}$ where $\mathcal{Y}_{1}$ is the set of $y\in\mathcal{Y}$ seen only one time, i.e. there exist a unique $(i,t)\in\boldsymbol{T}$ such that $y_{i,t}=y$, $\mathcal{Y}_{2}$ is the set of $y\in\mathcal{Y}$ such that there exists an edge $(x,y)\in\mathrm{A}_{\gamma}$ of multiplicity at least two and such that $|\mathcal{X}_{\operatorname{cc}}|\geq 2$ and finally $\mathcal{Y}_{3}$ is the set of $y\in\mathcal{Y}$ seen by at least two different edges of $\mathrm{A}_{\gamma}$. Recall that the edges $e\in\mathrm{A}_{\gamma}$ are not multiple, but have a multiplicity instead. Precisely, recall that for any $e=(x,y)\in\mathrm{A}_{\gamma}$ there exists $(i,t)\in\boldsymbol{T}$ such that we have
$(x_{i,t},y_{i,t})=(x,y)\text{ and }(\operatorname{cc}^{\prime},\operatorname{cc})=(\operatorname{cc}(x_{i,t}),\operatorname{cc}(x_{i,t+1}))$
and the multiplicity of an edge $e\in\mathrm{A}_{\gamma}$ is the number of $(i,t)$ verifying the previous equation. Then the set $\mathcal{Y}_{1}$ is in bijection with the corresponding incoming edges and these edges are inconsistent and of multiplicity one (see 6.5). Therefore we have $|\mathcal{Y}_{1}|\leq a_{1}$. If $y$ is seen by different edges, at least one of them is either discovered at a merging time or is an excess edge and therefore $|\mathcal{Y}_{3}|\leq g+m$. Finally the number of $y$ seen several times by the same edge and such that $|\mathcal{X}_{\operatorname{cc}}|\geq 2$ is bounded by the number of corresponding edges. We denote by $\mathrm{E}_{22}$ the number of edges $e:=(x,y)\in\mathrm{A}_{\gamma}$ linking $\operatorname{cc}^{\prime}$ to $\operatorname{cc}$ for some $\operatorname{cc}^{\prime},\operatorname{cc}\in\chi$ and such that $m(e)\geq 2$ and $|\mathcal{X}_{\operatorname{cc}}|\geq 2$. We proceed as in the proof of [BQZ18, Lemma 5] to bound $\mathrm{E}_{22}$. We call in-degree of a connected component $\operatorname{cc}\in\chi$ the number of edges $e\in\mathrm{A}_{\gamma}$ that links some $\operatorname{cc}^{\prime}\in\chi$ to $\operatorname{cc}$. Denoting $c_{k}$ the number of $\operatorname{cc}\in\chi$ with in-degree equal to $k$ and $c_{\geq 2}:=\sum_{k\geq 2}c_{k}$ we have:
$c_{0}+c_{1}+c_{\geq 2}=s-p\leavevmode\nobreak\ \leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ \leavevmode\nobreak\ c_{1}+2c_{\geq 2}\leq\sum_{k}kc_{k}=a.$
Subtracting to the right-hand side, twice the left hand side,
$c_{1}\geq 2(s-p)-a-2c_{0}\geq a-2g-2m+2.$
The bound $c_{0}\leq 2m$ follows from the observation that only a component $\operatorname{cc}\in\chi$ such that $\operatorname{cc}=\operatorname{cc}(x_{j,1})$ for some $1\leq j\leq 2m$ can be of in-degree 00. Besides, denoting $\mathrm{E}_{12}$ the number of edges of $\mathrm{A}_{\gamma}$ that are of multiplicity at least $2$ and such that the outgoing connected component $\operatorname{cc}\in\chi$ verifies $|\mathcal{X}_{\operatorname{cc}}|=1$, we observe that $c_{1}\leq a_{1}+\mathrm{E}_{12}$. Indeed we can associate each connected component $\operatorname{cc}$ of in-coming degree equal to one to its only in-coming edge. In particular we have
$|\mathcal{Y}_{3}|\leq\mathrm{E}_{22}=a-a_{1}-\mathrm{E}_{12}\leq a-c_{1}\leq 2g+2m-2\leavevmode\nobreak\ \leavevmode\nobreak\ \text{and therefore}\leavevmode\nobreak\ \leavevmode\nobreak\ F\leq 3(m+g)-2+a_{1}+p.$
The last lemma of this section will estimate the contribution $\mu(\gamma)$ defined in (28) using 6.5.
Lemma 7.9.
There is a constant $c>0$ such that, if $\gamma\in\mathrm{W}_{\ell,\vec{T}}(s,a,p)$, $g=a-s+p+1$ and $a_{1}$ is the number of arcs in $\mathrm{A}_{\gamma}$ visited exactly once by $\gamma$ then
$\left|\mu(\gamma)\right|\leq c^{m+g}N^{-a}\left(\frac{6\ell m}{\sqrt{N}}\right)^{(a_{1}-4g-2m+2p)_{+}}.$
Moreover we have $a_{1}\geq 2(a-(\ell+1)m)$.
Proof.
Let $\mathrm{A}_{1}\subset\mathrm{A}_{\gamma}$ be the set of $e=(x,y)$ which are visited exactly once in $\gamma$, i.e. we have
$\sum_{i=1}^{2m}\sum_{t=1}^{T_{i}}\mathds{1}(e_{i,t}=e)=1.$
Let $\mathrm{A}_{\ast}$ be the set of inconsistent arcs (see definitions above 6.5) and we denote $b$ its cardinal. We consider $\mathrm{A}_{1}^{\prime}\subset\mathrm{A}_{1}$ the subset of consistent arcs of multiplicity one. We have
$|\mathrm{A}_{1}^{\prime}|+|\mathrm{A}_{\ast}|\geq|\mathrm{A}_{1}|$
We set $a_{1}^{\prime}:=|\mathrm{A}_{1}^{\prime}|$ and $a_{\geq 2}:=|\mathrm{A}_{\gamma}\backslash\mathrm{A}_{1}|$. We have
$a_{1}+a_{\geq 2}=a\leavevmode\nobreak\ \leavevmode\nobreak\ \text{ and }\leavevmode\nobreak\ \leavevmode\nobreak\ a_{1}+2a_{\geq 2}\leq\sum_{k}ka_{k}=T\leq 2m(\ell+1).$
Therefore:
$a_{1}\geq 2(a-m(\ell+1)).$
It gives the second claim of the lemma. We now count the inconsistent edges using the vocabulary of the the proof of 7.2. Exploring $\gamma$ chronologically, seeing a new inconsistent edge means that we discover an edge with either one of its ending vertices that had been seen before. That can happen either at the beginning of a sequence of times that goes out of the existing forest so far, at a merging time or at a first visit of an excess edge. Every time $(i,t)$ where we start a sequence of times going out of the existing forest and such that we discover a new inconsistent edge is preceded by a first visit of an excess edge. Every such step can create $2$ inconsistent edge: the one we see at this time and the one seen before with at least one of its ending point in common. Therefore we have
$b\leq 4g+2m.$
However the bound on $b$ can be slightly improved. Indeed each time we discover a new element of an already seen connected component it means that either we have taken an excess edge or we are at a merging time and the corresponding edge $(x,y)$ is such that $x$ has never been seen before and therefore no possibly inconsistent outgoing edge was discovered before. Also either the $y$ in this edge has never been seen before, and therefore no new inconsistent edge here, or it has been seen before and we are in presence of an excess edge. In conclusion each time we discover a new element of a connected component $\operatorname{cc}$, we have counted $2$ new inconsistent edges that cannot be. Therefore we have in fact
$b\leq 4g+2m-2p$
and
$a^{\prime}_{1}\geq a_{1}-4g-2m+2p.$
It remains to apply 6.5. ∎
We have all the ingredients to now prove 7.1.
Proof of 7.1.
We prove the case #=#\#$\emptyset$# = ∅ and all the computations are exactly the same for #$\#=r$. We set
$m=\lceil\ln(N)^{\alpha}\rceil.$
We have
$\mathbb{P}\left(\|\underline{A}^{(\ell)}\|\geq e^{C\ln(N)^{1-\alpha}}\rho_{n}^{\ell}\right)\leq\mathbb{P}\left(\|\underline{A}^{(\ell)}\|^{2m}\geq N^{C}\rho_{n}^{2m\ell}\right).$
Applying Markov inequality to the previous probability and owing to (30), proving our proposition boils down to proving that there exists a constant $\mathfrak{c}=\mathfrak{c}(n)>0$ such that
$S:=\sum_{s,a,p,\vec{T}}\left|\mathcal{W}_{\ell,\vec{T}}(s,a,p)\right|\underset{\gamma\in\mathrm{W}_{\ell,\vec{T}}(s,a,p)}{\max}\left\{|\mu(\gamma)|\sum_{\gamma^{\prime}}q(\gamma^{\prime})\right\}\leq N^{\mathfrak{c}}\rho_{n}^{2m\ell}.$ (34)
Let $\gamma\in\mathrm{W}_{\ell,\vec{T}}(s,a,p)$ and $a_{1}$ be the number of $(x,y)\in\mathrm{A}_{\gamma}$ of multiplicity one. Applying 7.5 and 7.9 we have
$|\mu(\gamma)|\sum_{\gamma^{\prime}\sim\gamma}q(\gamma^{\prime})\leq N^{s-p}(cm\ell)^{a_{1}+3(m+g)+p}c(n)\rho_{n}^{2m(\ell+1)}c^{m+g}N^{-a}\left(\frac{6\ell m}{\sqrt{N}}\right)^{(a_{1}-4g-2m+2p)_{+}}.$
In regard of the conditions on $Q$ in 2.2 and the definition of $\rho_{n}$, $\ell$ and $m$, we find $c>1$ such that $\rho_{n}\leq c\ell m$. Since $cm\ell\geq 1$ we have $(cm\ell)^{a_{1}+p}\leq(cm\ell)^{(a_{1}-4g-2m+2p)_{+}}(2m\ell)^{4g+2m}$ and using that $a_{1}\geq 2(a-m(\ell+1))$ we deduce:
$|\mu(\gamma)|\sum_{\gamma^{\prime}\sim\gamma}q(\gamma^{\prime})\leq c(n)(cm\ell)^{8g+8m}N^{-g+1}\rho_{n}^{2m\ell}\left(\frac{(cm\ell)^{4}}{N}\right)^{(a-(\ell+2)m-2g+p)_{+}},$
where $c>0$ is adapted if necessary. To ease notation we set
$\epsilon:=\frac{(cm\ell)^{4}}{N}.$
Considering any $\ell=\mathcal{O}(\ln(N))$, we have $\epsilon=o(1)$. We apply 7.2 to $S$ defined by (34) and sum over all $\vec{T}$ possible. We have at most $\ell^{2m}$ possibilities for $\vec{T}$ and considering the fact that for $c>1$ large enough we have $a,s$ and $\ell$ bounded by $cm\ell$ we obtain
$S^{\prime}$ $:=\frac{S}{\rho_{n}^{2m\ell}}\leq c(n)N\sum_{\begin{subarray}{c}s,a,p\\ g(s,a,p)\geq 0,\leavevmode\nobreak\ p\leq 2(m+g)\end{subarray}}\ell^{2m}(2m\ell)^{12m(g+5)}(cm\ell)^{8g+8m}N^{-g}\epsilon^{(a-(\ell+2)m-2g+p)_{+}}$
$\leq cN(cm\ell)^{80m}\sum_{\begin{subarray}{c}s,g,p\\ g(s,a,p)\geq 0,\leavevmode\nobreak\ p\leq 2(m+g)\end{subarray}}(cm\ell)^{20mg}N^{-g}\epsilon^{(s-\ell^{\prime}m-g)_{+}},$
where we performed the change of variable $g=a-s+p+1$ and where we consider $\ell^{\prime}:=\ell+2+1/m$. Now summing over $p\leq 2(m+g)$ we have in the sum an additional term $2(m+g)$ which we bound with $cm\ell$. Considering $\mathfrak{c}>0$ such that $c(n)(cm\ell)^{81m}\leq N^{\mathfrak{c}}$:
$S^{\prime}\leq N^{\mathfrak{c}}\sum_{s,g}\left(\frac{L}{N}\right)^{g}\epsilon^{(s-\ell^{\prime}m-g)_{+}}$
where we set $L:=(cm\ell)^{20m}.$ Decomposing the sum on the right hand side above we have
$S^{\prime}\leq S_{1}+S_{2}+S_{3}$
[…]
$\frac{L}{N}=\frac{e^{c\ln(\ln(N))\ln(N)^{\alpha}}}{N}\leq\frac{1}{2}.$
and therefore
$S_{1}\leq N^{\mathfrak{c}}\sum_{s=1}^{\ell^{\prime}m}2\leq N^{\mathfrak{c}^{\prime}}.$
Similarly, using that $\frac{L}{N\epsilon}\geq 2$ for $N$ large enough and adapting the constants we have
$S_{2}$ $=N^{\mathfrak{c}}\sum_{s=\ell^{\prime}m+1}^{+\infty}\sum_{g=0}^{s-\ell^{\prime}m}\left(\frac{L}{N\epsilon}\right)^{g}\epsilon^{s-\ell^{\prime}m}$
$\leq N^{\mathfrak{c}^{\prime}}\sum_{s=\ell^{\prime}m+1}^{+\infty}\left(\frac{L}{N}\right)^{s-\ell^{\prime}m}\leq N^{\mathfrak{c}^{\prime}}\sum_{k=0}^{+\infty}\left(\frac{L}{N}\right)^{k}.$
Once again considering $N$ large enough the geometric sum converges and updating the constant we have the wanted bound. Finally for $N$ large enough we have
$S_{3}$ $=N^{\mathfrak{c}}\sum_{s=\ell^{\prime}m+1}^{+\infty}\sum_{g=s-\ell^{\prime}m+1}^{+\infty}\left(\frac{L}{N}\right)^{g}\leq N^{\mathfrak{c}}\sum_{s=\ell^{\prime}m+1}^{+\infty}2\left(\frac{L}{N}\right)^{s-\ell^{\prime}m+1}\leq N^{\mathfrak{c}^{\prime}}\sum_{k=0}^{+\infty}\left(\frac{L}{N}\right)^{k}.$
It remains to consider $N$ large enough and $\mathfrak{c}^{\prime\prime}>0$ large enough so (34) holds. ∎
7.4 Operator norm of $R^{(\ell)}_{k,\#}$
We will now adapt what have been doing to bound the operator norm $A_{\#}^{(\ell)}$ to upper bound the operator norm of $R^{(\ell)}_{k,\#}$ defined by (24) for #=#\#$\emptyset$# = ∅ and (25) for #$\#=r$. We set
$\Delta:=1+\kappa+\delta\leavevmode\nobreak\ \leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ \leavevmode\nobreak\ \Delta_{r}:=1+r\kappa+r\delta$
where $\kappa$ is defined by (2) and $\delta>0$ comes from (H:1).
Proposition 7.10.
Let $(Q,M),\leavevmode\nobreak\ n,\leavevmode\nobreak\ \rho_{n}$ and $A$ be as defined in 2.2. For any $c>0$, $10<\alpha<$ 1 and $1\leq k\leq\ell\leq\frac{\ln(N)}{7\ln(d)}$, there exists $C=C(c,\alpha,n,\delta)>0$ such that with probability at least $1-N^{-c}$,
$\|R_{k,\#}^{(\ell)}\|\leq e^{C\ln(N)^{1-\alpha}}\Delta_{\#}^{\ell}.$
In all this section the notations are the one introduced in Section 7.1. We also give the proof of the previous proposition only in the case #=#\#$\emptyset$# = ∅. We again use the inequality
$\|R^{(\ell)}_{k}\|^{2m}\leq\operatorname{Tr}\left\{\left(R_{k}^{(\ell)}{R_{k}^{(\ell)}}^{\top}\right)^{m}\right\}=:\operatorname{Tr}[k,\ell,m].$ (35)
Let $\vec{T}\in[\ell]^{2m}$ be fixed. We denote by $\mathrm{W}^{\prime}_{\ell,\vec{T}}$ the set of $\gamma:=(\gamma_{1},...,\gamma_{2m})$ such that for all $i\in[2m]$
  • $T(\gamma_{i}):=T_{i}$ and $\gamma_{i}=(x_{i,1},y_{i,1},\ell_{i,1},x_{i,2},...,x_{i,T_{i}},y_{i,T_{i}},\ell_{i,T_{i}},x_{i,T_{i}+1})\in\mathrm{T}^{k,\ell}$,
  • the condition (26) is verified.
With this notations, expending the trace in (35) gives
$\operatorname{Tr}[k,\ell,m]=\sum_{\vec{T}\in[\ell]^{2m}}\sum_{\gamma\in\mathrm{W}^{\prime}_{\ell,\vec{T}}}\prod_{i=1}^{2m}\prod_{t=1}^{k-1}\underline{M}_{x_{i,t}y_{i,t}}(Q^{\ell_{i,t}})_{y_{i,t}x_{i,t+1}}\cdot(Q^{\ell_{i,k}})_{y_{i,k}x_{i,k+1}}\cdot\prod_{t=k+1}^{T_{i}}M_{x_{i,t}y_{i,t}}(Q^{\ell_{i,t}})_{y_{i,t}x_{i,t+1}}.$ (36)
In regard of the definition of $\mathrm{T}^{k,\ell}$, if $\vec{T}\in[\ell]^{2m}$ and there exists $i$ such that $T_{i}<k$ then $\mathrm{W}^{\prime}_{\ell,\vec{T}}=\emptyset$. For $i\in[2m]$ we set
$\gamma_{i}^{\prime}=(x_{i,1},y_{i,1},\ell_{i,1},x_{i,2},...,x_{i,k-1},y_{i,k-1},\ell_{i,k-1},x_{i,k})$
and
$\gamma_{i}^{\prime\prime}=(x_{i,k+1},y_{i,k+1},\ell_{i,k+1},...,x_{i,T_{i}},y_{i,T_{i}},\ell_{i,T_{i}},x_{i,T_{i}+1}).$
As in Section 7.1 we define the graph $K_{\gamma}$ for a fixed $\gamma\in\mathrm{W}^{\prime}_{\ell,\vec{T}}$. We consider $X_{\gamma}:=\{x_{i,t}:i\in[2m],t\in[T_{i}+1]\}$ and $Y_{\gamma}:=\{y_{i,t}:i\in[2m],t\in[T_{i}]\}$. As before we set $K_{\gamma}=(X_{\gamma},E(K_{\gamma}))$ and $\{x,x^{\prime}\}$ is a vertex if there exists $(i,t),(j,s)\in\boldsymbol{T}$ such that
$x_{i,t+1}=x,\leavevmode\nobreak\ x_{j,s+1}=x^{\prime}\leavevmode\nobreak\ \leavevmode\nobreak\ \text{ and }\leavevmode\nobreak\ \leavevmode\nobreak\ y_{i,t}=y_{j,s}.$
For $x\in X_{\gamma}$ we consider again $\operatorname{cc}(x)$ its connected component in $K_{\gamma}$ and $\chi_{\gamma}=X_{\gamma}/\sim$ the set of connected component of $K_{\gamma}$. The arcs of $\gamma$ are the pairs $(x_{i,t},y_{i,t})$ for $t\neq k$ and we denote by $\mathrm{A}^{\prime}_{\gamma}$ the set of arcs of $\gamma$. For $s,a,p$ integers, we denote by $\mathrm{W}^{\prime}_{\ell,\vec{T}}(s,a,p)$ the set of $\gamma\in\mathrm{W}^{\prime}_{\ell,\vec{T}}$ verifying $|X_{\gamma}|=s$, $|X_{\gamma}/\sim|=s-p$ and $|\mathrm{A}_{\gamma}^{\prime}|=a$. Adapting the computations for the case #$\#=r$ and taking the expectation of (36) we have
$\mathbb{E}\|R_{k,\#}^{(\ell)}\|^{2m}\leq\mathbb{E}(\operatorname{Tr}[k,\ell,m])=\sum_{\vec{T},s,a,p}\leavevmode\nobreak\ \leavevmode\nobreak\ \sum_{\gamma\in\mathrm{W}^{\prime}_{\ell,\vec{T}}(s,a,p)}\mu^{\prime}(\gamma)q_{\#}(\gamma)$
where for $\gamma\in\mathrm{W}^{\prime}_{\ell,\vec{T}}(s,a,p)$ we define
$\mu^{\prime}(\gamma):=\mathbb{E}\prod_{i=1}^{2m}\prod_{t=1}^{k-1}\underline{M}_{x_{i,t}y_{i,t}}\prod_{t=k+1}^{T_{i}}M_{x_{i,t}y_{i,t}}$ (37)
and
$q(\gamma):=\prod_{i=1}^{2m}\prod_{t=1}^{T_{i}}(Q^{\ell_{i,t}})_{y_{i,t}x_{i,t+1}}\leavevmode\nobreak\ \leavevmode\nobreak\ \text{ and }\leavevmode\nobreak\ \leavevmode\nobreak\ q_{r}(\gamma):=(1-r)^{T-1}\prod_{i=1}^{2m}\prod_{t=1}^{T_{i}}(\{rQ\}^{\ell_{i,t}})_{y_{i,t}x_{i,t+1}}$ (38)
where for $\vec{T}=(T_{1},...,T_{2m})\in[\ell]^{2m}$ we set $T=\sum_{i}T_{i}$. We decompose $\mathrm{W}^{\prime}_{\ell,\vec{T}}(s,a,p)$ into equivalent classes. We say that $\gamma\sim\gamma^{\prime}$ if there exists $\alpha,\beta$ permutations of $[N]$ such that for all $i\in[2m]$ and all $t\in[T_{i}+1]$ we have $\alpha(x_{i,t})=x^{\prime}_{i,t}$ and $\beta(y_{i,t})=y^{\prime}_{i,t}$, where $\gamma^{\prime}:=(\gamma_{1}^{\prime},...,\gamma_{2m}^{\prime})$ and $\gamma_{i}^{\prime}:=(x_{i,1}^{\prime},y_{i,1}^{\prime},\ell_{i,1}^{\prime},....,x^{\prime}_{i,T_{i}},y^{\prime}_{i,T_{i}},\ell_{i,T_{i}}^{\prime},x^{\prime}_{i,T_{i}+1})$ for $i\in[2m]$. We denote by $\mathcal{W}^{\prime}_{\ell,\vec{T}}(s,a,p):=\mathrm{W}^{\prime}_{\ell,\vec{T}}(s,a,p)/\sim$ the set of equivalent classes. We have
$\mathbb{E}\|R_{k,\#}^{(\ell)}\|^{2m}\leq\sum_{\vec{T},s,a,p}|\mathcal{W}^{\prime}_{\ell,\vec{T}}(s,a,p)|\underset{\gamma\in\mathrm{W}^{\prime}_{\ell,\vec{T}}(s,a,p)}{\max}\left(|\mu^{\prime}(\gamma)|\sum_{\gamma^{\prime}\sim\gamma}q_{\#}(\gamma^{\prime})\right).$ (39)
As before we first bound the cardinality of $\mathcal{W}^{\prime}_{\ell,\vec{T}}(s,a,p)$, independently of the choice of $\vec{T}$.
Lemma 7.11.
If $g^{\prime}:=a+p-s<0$ or $2g^{\prime}+10m>p$ then $\mathrm{W}^{\prime}_{\ell,\vec{T}}(s,a,p)$ is empty. Otherwise we have:
$|\mathcal{W}^{\prime}_{\ell,\vec{T}}(s,a,p)|\leq(\ell(a+2m)^{2}(s-p)^{2}s^{2})^{4m(g^{\prime}+7)}.$
Proof.
We repeat the construction given in Section 7.2. Let $\ell\geq 1$, $\vec{T}\in[\ell]^{2m}$ and $\gamma\in\mathrm{W}^{\ell,\vec{T}}(s,a,p)$ be fixed. We set $\gamma_{i,t}:=(x_{i,t},y_{i,t},x_{i,t+1})$ for $i\in[2m]$ and $t\in[T_{i}]$. We explore $(\gamma_{i,t})_{(i,t)\in\boldsymbol{T}}$ chronologically where $\boldsymbol{T}:=\{(i,t):i\in[2m],t\in[T_{i}]\backslash\{k\}\}$ is ordered as in Section 7.2. Recall in particular that for $i\leq i^{\prime}$, for all $t\in[T_{i}]\backslash\{k\}$ and $t^{\prime}\in[T_{i^{\prime}}]\backslash\{k\}$ we have $(i,t)\preceq(i^{\prime},t^{\prime})$, and for all $t\leq t^{\prime}\in[T_{i}]\backslash\{k\}$ we have $(i,t)\preceq(i,t^{\prime})$.
For $y\in Y_{\gamma}$, we denote by $\bar{y}$ the order of apparition of $y$ in the ordered sequence $(y_{i,t})_{(i,t)\in\boldsymbol{T}}$. The notation $\bar{x}$ and $\bar{\operatorname{cc}}(x)$ are defined similarly in regard of the apparition order of $x\in X_{\gamma}$ or $\operatorname{cc}(x)\in K_{\gamma}/\sim$. Finally for $x^{\prime}\in X_{\gamma}$, we set the connected component color of $x$ as the function $\bar{c}(\bar{x})$ that return the first element ever seen of the connected component of $x$, i.e.
$\bar{c}(\bar{x}):=\min\{\bar{x}^{\prime}\leq\bar{x}:x^{\prime}\in\operatorname{cc}(x)\}.$
Finally we define the connected component mark of $x\in X_{\gamma}$ by $\vec{x}:=(\bar{c}(\bar{x}),\bar{x})$ and we set $\bar{\gamma}_{i,t}:=(\vec{x}_{i,t},\bar{y}_{i,t},\vec{x}_{i,t+1})$ for all $(i,t)\in\boldsymbol{T}$. By construction, knowing $(\bar{\gamma}_{i,t})_{(i,t)\in\boldsymbol{T}}$ is equivalent to knowing the equivalent class of $\gamma$. We will now encode the sequence $(\bar{\gamma}_{i,t})_{(i,t)\in\boldsymbol{T}}$.
We set $V_{\gamma}:=[s-p]$ and consider the colored directed graph $G_{\gamma}^{\prime}:=(V_{\gamma},E_{\gamma}^{\prime})$ defined as follows. For each time $(i,t)\in\boldsymbol{T}$ such that $t\neq k$, we put a directed edge $e_{i,t}:=(\bar{\operatorname{cc}}(x_{i,t}),\bar{\operatorname{cc}}(x_{i,t+1}))\in E_{\gamma}^{\prime}$ and colorize it with $(\bar{x}_{i,t},\bar{y}_{i,t})$. By definition we will have $|\mathrm{A}_{\gamma}^{\prime}|=|E_{\gamma}^{\prime}|=a$. We denote by $\bar{G}^{\prime}_{\gamma}$ the associate undirected graph. We observe that each connected component of $\bar{G}^{\prime}_{\gamma}$ has at least one cycle. Indeed let $\mathcal{C}\subset\bar{G}^{\prime}_{\gamma}$ be a connected component and $i\in[2m]$ such that $\operatorname{Im}(\gamma_{i})\cap\mathcal{C}\neq\emptyset$. Then the path $\gamma_{i}$ being tangled we have at least two cycles in the image of $\gamma_{i}$. Also for all $\epsilon\in\{\prime,\prime\prime\}$, by definition $\operatorname{Im}(\gamma^{\epsilon}_{i})$ is contained in one connected component of $\bar{G}^{\prime}_{\gamma}$. Therefore we have only two possibilities. Either $\operatorname{Im}(\gamma_{i}^{\prime})\cap\operatorname{Im}(\gamma_{i}^{\prime\prime})\neq\emptyset$ and therefore $\operatorname{Im}(\gamma_{i})\subset\mathcal{C}$, then both cycles (and the others) of $\gamma_{i}$ are contained in $\mathcal{C}$. One could argue that one of the edges of one of the cycle might be $(x_{i,k},y_{i,k})$, which is not an element of $E_{\gamma}^{\prime}$. In this case, the fact that we have two cycles in $\gamma_{i}$ implies that removing one edge leaves at least one cycle. Otherwise, $\gamma_{i}$ being tangled and the images of $\gamma_{i}^{\epsilon}$ being distinct implies that both $\gamma_{i}^{\prime}$ and $\gamma_{i}^{\prime\prime}$ have a cycle and therefore at least one is in $\mathcal{C}$. We deduce from this fact that in each connected component we have at least as much edges than we have vertices and therefore, summing over all connected component
$0\leq g^{\prime}=|E_{\gamma}^{\prime}|-|V_{\gamma}|=a-s+p.$
This is the first claim of the lemma.
For every time $(i,t)\in\boldsymbol{T}$ we define $G_{i,t}^{\prime}$ as the subgraph of $G_{\gamma}^{\prime}$ spanned by the edges $e_{j,s}$ for all $(j,s)\preceq(i,t)$. We have $G_{2m,T_{2m}}^{\prime}=G_{\gamma}^{\prime}$ and we again define a spanning forest $\Gamma_{i,t}$ of $G_{i,t}^{\prime}$ inductively. We set $\Gamma_{1,0}=(\{1\},\emptyset)$. We say that $(i,t)\in\boldsymbol{T}$ is a first time if adding $e_{i,t}$ to $\Gamma_{(i,t)^{-}}$ does not create a weak cycle. If $(i,t)\in\boldsymbol{T}$ is a first time then $\Gamma_{i,t}:=\operatorname{Span}\{\Gamma_{(i,t)^{-}},e_{i,t}\}$, otherwise $\Gamma_{i,t}=\Gamma_{(i,t)^{-}}$. Finally we set $\Gamma:=\Gamma_{2m,T_{2m}}$.
For each even era $i\in[2m]$, we define the first merging time $(i,t^{\prime}_{i})$ as the smallest time $(i,t)$ with $11\leq t\leq k-$ 1, if it exists, such that $\Gamma_{i,t}$ and $\Gamma_{(i,1)^{-}}$ have the same number of connected component. If this time does not exists exist we take as convention $t^{\prime}_{i}=k$. Similarly, for each $i$, we define the second merging time $(i,t_{i}^{\operatorname{\prime\prime}})$ as the smallest time $(i,t)$ with $k\leq t\leq T_{i}$ such that $\Gamma_{i,t}$ has the same number of connected components than $\Gamma_{(i,k)^{-}}$. If this time does not exists we set $t^{\operatorname{\prime\prime}}_{i}=T_{i}+1$. In regard of (26), we have $t^{\operatorname{\prime\prime}}_{i}\leq T_{i}$ for $i$ even. Also notice that if $t_{i}^{\epsilon}\geq 2$ then the merging time is a first time.
The edges of $G_{\gamma}^{\prime}\backslash\Gamma$ are called excess edges. The times $(i,t)\in\boldsymbol{T}$ such that $e_{i,t}$ is an excess edge are called important time. Denoting $c_{\gamma}$ the number of connected component of $G_{\gamma}^{\prime}$, the total number excess edges is
$|G_{\gamma}^{\prime}\backslash\Gamma|=|E_{\gamma}^{\prime}|-|V_{\gamma}|+c_{\gamma}=g^{\prime}+c_{\gamma}.$
To recover the previous equality, we sum the number of excess edges in each connected component. Owing to (26), there is at most one new connected component appearing for each era and therefore we have $1\leq c_{\gamma}\leq 2m$. Besides there is at least one cycle in each connected component therefore there are at most $g^{\prime}+1$ excess edges in each connected component of $G_{\gamma}^{\prime}$.
By construction the path $\gamma_{i}^{\prime}$ and $\gamma^{\operatorname{\prime\prime}}_{i}$ can be decomposed by a successive repetition of the pattern (1)(2)(3) given by
  1. (1)
    a sequence of first times (possibly empty)
  2. (2)
    an important time or a merging time
  3. (3)
    a path using the colored edges of the forest defined so far (possibly empty).
We encode the times as follows. If $(i,t)\in\boldsymbol{T}$ is an important time we mark $(i,t)$ by the vector $(\bar{y}_{i,t},\vec{x}_{i,t+1},\vec{x}_{i,\tau})$ where $(i,\tau)$ is the next step outside $\Gamma_{i,t}$ (by convention if $\gamma_{i}$ stays in the forest for the remaining time of the era $i$ we set $\tau=T_{i}+1$). By construction $(i,\tau)$ is the next first, important or merging time. Similarly if $(i,t)$ is a first merging time, we mark the time $(i,t)$ by the first merging time mark $(\bar{y}_{i,t},\vec{x}_{i,t+1},\vec{x}_{i,t})$ where $(i,\tau)$ is the next step outside $\Gamma_{i,t}$. For $(i,t)$ a second merging time we mark it with the second merging time mark $(\bar{y}_{i,k},\bar{y}_{i,t},\vec{x}_{i,t+1},\vec{x}_{i,\tau})$ where $(i,\tau)$ is first time outside $\Gamma_{i,t}$. The positions of the important and merging times and their marks encodes $(\bar{\gamma}_{i,t})_{(i,t)\in\boldsymbol{T}}$. Besides denoting $\operatorname{cc}_{k}$ the set of vertices of the $k$-th connected component of $K_{\gamma}$ we have that $p=\sum_{k=1}^{s-p}(|\operatorname{cc}_{k}|-1)$. Hence $p$ is equal to the number of times we discover a new element $x_{i,t}$ in a connected component that we already have seen. The number of such times is upper bounded by twice the number of excess edges plus the number of merging times
$p\leq 2(g^{\prime}+N_{\gamma}+3m)\leq 2g^{\prime}+10m.$
It proves the second statement of the lemma.
As in Section 7.2, the number of important times encoded is to important for our purposes, therefore we improve the way we encode the paths taking into account the tangle-freeness of $\gamma^{\epsilon}_{i}$. We partition the encoding era wise, and partition the important times in each era into three categories: short cycling times, long cycling times and superfluous times. Let $i\in[2m]$ and $\epsilon\in\{\prime,\operatorname{\prime\prime}\}$ be fixed. If it exists, let $(i,t_{0})$ be the smallest time such that $\operatorname{cc}(x_{i,t_{0}+1})\in\{\operatorname{cc}(x_{i,\ast}),\operatorname{cc}(x_{i,\ast+1}),...,\operatorname{cc}(x_{i,t_{0}})\}$ where =1$\ast$=1∗ = 1 if $\epsilon=\prime$ and $\ast=k+1$ otherwise. Let $\ast\leq\sigma\leq t_{0}$ be such that $\operatorname{cc}(x_{i,t_{0}+1})=\operatorname{cc}(x_{i,\sigma})$. By assumption, $\mathrm{C}_{i}=(\operatorname{cc}(x_{i,\sigma}),...,\operatorname{cc}(x_{i,t_{0}+1}))$ will be the only cycle of $G_{\gamma}^{\prime}$ seen by $\gamma_{i}^{\epsilon}$. The last important time $(i,t)\preceq(i,t_{0})$ is called the short cycling time. We denote by $(i,\hat{t})$ the smallest time $(i,\hat{t})\succeq(i,\sigma)$ such that $\operatorname{cc}(x_{i,\hat{t}+1})\notin\mathrm{C}_{i}$. If $\gamma_{i}^{\epsilon}$ stays in $\Gamma_{i,t}$ for its remaining time, by convention we set $\hat{t}=k$ (resp. $\hat{t}=T_{i}+1$) for $\epsilon=\prime$ (resp. for $\epsilon=\prime\prime$). We replace the mark of the short cycling time by $(\bar{y}_{i,t},\vec{x}_{i,t+1},\sigma,\hat{t},\vec{x}_{i,\tau})$ where once again $(i,\tau)$ is such that $\tau\geq t$ and $(i,\tau)$ is the next step outside of $\Gamma_{i,t}$. The time $(i,\tau)$ is the next first, important or merging time after $(i,\hat{t})$. If $\gamma_{i}^{\epsilon}$ stays in $\Gamma_{i,t}$ we set $\tau=k$ (resp. $\tau=T_{i}+1$) for $\epsilon=\prime$ (resp. for $\epsilon=\operatorname{\prime\prime}$). Important times $(i,t^{\prime})$ with $1\leq t^{\prime}\leq t$ or $\tau\leq t^{\prime}\leq k$ (or in the case $\epsilon=\operatorname{\prime\prime}$, $k+1\leq t^{\prime}\leq t$ or $\tau\leq t^{\prime}\leq T_{i}$) are called long cycling times. The other important times with $t<t^{\prime}<\tau$ are called superfluous.
We have our second encoding and we can recover the sequence $(\bar{\gamma}_{i,t})_{(i,t)\in\boldsymbol{T}}$ from the positions of the merging times, the long cycling times, the short cycling times and their respective marks. For each $i\in[2m]$ and for each $\epsilon\in\{\prime,\operatorname{\prime\prime}\}$, there is at most one short cycling time, one merging time, and as argued in the proof of 7.2 $g^{\prime}+1$ long cycling times. For each time we have $T_{i}\leq\ell$ or $k\leq\ell$ ways to position them. Therefore we have a total of $\ell^{4m(g^{\prime}+3)}$ ways to positions our marked times. We have $|Y_{\gamma}|\leq a+2m=:a^{\prime}$, where the term $2m$ coming from the element $y_{i,k}$ for $i\in[2m]$. We have at most ${a^{\prime}}(s-p)^{2}s^{2}$ ways to mark a long cycling time, at most $a^{\prime}(s-p)^{2}s^{2}\ell^{2}$ ways to mark a short cycling time and ${a^{\prime}}^{2}(s-p)^{2}s^{2}$ ways to mark a merging time. We deduce from that
$|\mathcal{W}^{\prime}_{\ell,\vec{T}}(s,a,p)|$ $\leq\ell^{4m(g^{\prime}+3)}({a^{\prime}}(s-p)^{2}s^{2})^{4m(g^{\prime}+1)}(a^{\prime}(s-p)^{2}s^{2}\ell^{2})^{4m}({a^{\prime}}^{2}(s-p)^{2}s^{2})^{4m}$
$\leq(\ell(a+2m)^{2}(s-p)^{2}s^{2})^{4m(g^{\prime}+7)}.$
Lemma 7.12.
The exists $c>0$ such that for all $\gamma\in\mathrm{W}_{\ell,\vec{T}}^{\prime}(s,a,p)$,
$\sum_{\gamma^{\prime}\sim\gamma}q_{\#}(\gamma^{\prime})\leq N^{s-p}cm\ell\Delta_{\#}^{2m(\ell+1)}.$
For $\gamma\in\mathrm{W}_{\ell,\vec{T}}^{\prime}(s,a,p)$ we consider for each $\operatorname{cc}\in\chi_{\gamma}$ the connected and bi-partite graph in the sense 7.6 $\mathcal{G}_{\operatorname{cc}}=(\mathcal{X}_{\operatorname{cc}}\bigsqcup\mathcal{Y}_{\operatorname{cc}},\mathcal{E}_{\operatorname{cc}})$ and its underlying covering tree $\mathcal{T}_{\operatorname{cc}}$ constructed as we did in Section 7.3. We consider $\mathcal{G}$ the union graph as in Section 7.3 and
$\mathcal{F}=\bigsqcup_{\operatorname{cc}\in\chi}\mathcal{T}_{\operatorname{cc}},\leavevmode\nobreak\ \leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ \leavevmode\nobreak\ \mathcal{R}=\bigsqcup_{\operatorname{cc}\in\chi}\mathcal{R}_{\operatorname{cc}}.$
Proof.
We give the detail of the proof for the case #=#\#$\emptyset$# = ∅. We start again from (32), valid again here. Using the same notations then in the proof of 7.5, for a fixed choice of $\{(y_{u},x_{v}):\leavevmode\nobreak\ (u,v)\in\mathcal{Y}\times\mathcal{X}\}$, we rewrite and bound the second sum in (32) as follows
$\sum_{\ast}\prod_{\operatorname{cc}\in\chi}\prod_{f\in\mathcal{E}_{\operatorname{cc}}}(Q^{\ell(f)})_{y_{u}x_{v}}$ $=\sum_{L=0}^{M}\sum_{\begin{subarray}{c}\ast\ast\end{subarray}}\prod_{\operatorname{cc}\in\chi}\prod_{f\in\mathcal{T}_{\operatorname{cc}}}(Q^{\ell(f)})_{y_{u}x_{v}}\sum_{\ast\ast\ast}\prod_{\operatorname{cc}\in\chi}\prod_{f\in\mathcal{R}_{\operatorname{cc}}}(Q^{\ell(f)})_{y_{u}x_{v}}$
[…]
$\sum_{\gamma^{\prime}\sim\gamma}q(\gamma^{\prime})$ $\leq\sum_{L=0}^{M}(\kappa+1)^{M-L+R}\sum_{\ast\ast}\prod_{\operatorname{cc}\in\chi}\sum_{\begin{subarray}{c}y_{u},x_{v}\in[N]\\ (u,v)\in\mathcal{Y}_{\operatorname{cc}}\times\mathcal{X}_{\operatorname{cc}}\end{subarray}}\prod_{f\in\{e^{\prime}\}\sqcup\mathcal{T}_{\operatorname{cc}}}(Q^{\ell(f)})_{y_{u}x_{v}}$
$\leq\sum_{L=0}^{M}(\kappa+1)^{M-L+R}\sum_{\ast\ast}N^{s-p}\delta^{L}\leq N^{s-p}cm\ell(\kappa+\delta+1)^{2m(\ell+1)}$
where we applied 7.7 between the first and second line and bounded $M\leq cm\ell$ and $\left|\mathcal{L}^{\prime}(\mathcal{F},L)\right|\leq\binom{L+F-1}{F-1}$. We also used the fact that $M+R+F=2m(\ell+1)$. ∎
Lemma 7.13.
There exists a constant $c>0$ such that, if $\gamma\in\mathrm{W}_{\ell,\vec{T}}^{\prime}(s,a,p)$, denoting $g^{\prime}=a+p-s$ we have
$|\mu^{\prime}(\gamma)|\leq c^{m+g^{\prime}}N^{-a}.$
Proof.
We use the terminology used in the proof of 7.11 above. Let $A_{\ast}$ be the set of inconsistent arcs of $A_{\gamma}^{\prime}$ (as defined above in 6.5). As argued in the proof of 7.9,$|A_{\ast}|$ is bounded by four times the number of excess edges plus twice the number of merging times. There are at most $g^{\prime}+2m$ excess edges and $3m$ merging times, hence:
$|A_{\ast}|\leq 4(g^{\prime}+2m)+6m.$
Applying 6.5 we obtain the wanted result. ∎
We can now prove 7.10.
Proof of 7.10.
We set
$m=\lceil\ln(N)^{\alpha}\rceil.$
From (39) and Markov inequality, it suffices to prove that there exists $\mathfrak{c}>0$ such that
$S=\sum_{\vec{T},s,a,p}|\mathcal{W}^{\prime}_{\ell,\vec{T}}(s,a,p)|\underset{\gamma\in\mathrm{W}^{\prime}_{\ell,\vec{T}}(s,a,p)}{\max}\left(|\mu^{\prime}(\gamma)|\sum_{\gamma^{\prime}\sim\gamma}q_{\#}(\gamma^{\prime})\right)\leq N^{\mathfrak{c}}\Delta_{\#}^{2m\ell}.$ (40)
We fix $\ell\geq 0$ and $\vec{T}\in[\ell]^{2m}$ and $\gamma\in\mathrm{W}^{\prime}_{\ell,\vec{T}}(s,a,p)$. We set $g^{\prime}=g^{\prime}(s,a,p)=a-s+p$. Owing to 7.13 and 7.12 we have
$|\mu^{\prime}(\gamma)|\sum_{\gamma^{\prime}\sim\gamma}q(\gamma^{\prime})\leq c^{m+g^{\prime}}cm\ell\Delta_{\#}^{2m(\ell+1)}N^{-g^{\prime}}.$
Besides we find $c>0$ such that $\ell$, $a+2m$, $s$ and $(s-p)$ are bounded by $cm\ell$. Then owing to 7.11
$|\mathcal{W}^{\prime}_{\ell,\vec{T}}(s,a,p)|\leq(cm\ell)^{28m(g^{\prime}+7)}.$
Therefore summing over all $\vec{T}$ possible, i.e. at most $\ell^{2m}\leq(cm\ell)^{2m}$ we have
$S$ $\leq\sum_{\begin{subarray}{c}s,a,p;\\ g^{\prime}(s,a,p)\geq 0,\leavevmode\nobreak\ p\leq 2g^{\prime}+10\end{subarray}}\Delta_{\#}^{2m(\ell+1)}(cm\ell)^{28m(g^{\prime}+8)+1}N^{-g^{\prime}}c^{m+g^{\prime}}$
$\leq(c\ell m)^{227m}\Delta_{\#}^{2m(\ell+1)}\sum_{g^{\prime}}(c\ell m)^{28mg^{\prime}}N^{-g^{\prime}}$
where at the last line we performed the change of variable $a\to g^{\prime}=a+p-s$ and we summed over the variable $s\leq cm\ell$ and $p\leq 2g^{\prime}+10m\leq cm\ell$. Up to updating the constant $c>0$ we have
$S\leq N^{\mathfrak{c}}\Delta_{\#}^{2m\ell}\sum_{g^{\prime}}\left(\frac{L}{N}\right)^{g^{\prime}}$
with $L=(c\ell m)^{28m}\leq N/2$ for $N$ large enough. ∎
8 Proof of 5.1
All ingredients are now gathered to prove 5.1. Let $10<\alpha,r<$ 1, $10<c<c^{\prime}<$ 1 and $n\geq 0$ be fixed. We can assume for all that follows that $\ell\leq\frac{\ln(N)}{\ln(d)}$ since otherwise the bound of 6.4 would be larger than $1$. Notice that for $N$ large enough, $c\frac{\ell^{4}(4d)^{3\ell}}{N}\leq c\frac{(4d)^{6\ell}}{N}$. Therefore denoting $\Omega$ the event $(M,Q)$ is $\ell$-tangle free and applying 6.4 and 6.3, for any $C>0$ we have
$\mathbb{P}\left(\|A^{\ell}_{\#|_{\boldsymbol{1}^{\perp}}}\|\geq e^{C\ln(N)^{1-\alpha}}\rho_{n}^{\ell}\right)$ $\leq\mathbb{P}\left(\|A^{\ell}_{\#|_{\boldsymbol{1}^{\perp}}}\|\geq e^{C\ln(N)^{1-\alpha}}\rho_{n}^{\ell}\cap\Omega\right)+\frac{(4d)^{6\ell}}{N}$
$\leq\mathbb{P}\left(J\geq e^{C\ln(N)^{1-\alpha}}\rho_{n}^{\ell}\right)+\frac{(4d)^{6\ell}}{N},$
where
$J:=\|\underline{A}_{\#}^{(\ell)}\|+\frac{1}{N}\sum_{k=1}^{\ell}\|R_{k,\#}^{(\ell)}\|.$
Owing to 7.1 and 7.10 we find $C=C(n)>0$ such that with probability at least $1-2N^{-c^{\prime}}$
$J$ $\leq e^{C\ln(N)^{1-\alpha}}\rho_{n,\#}^{\ell}+\frac{1}{N}\sum_{k=1}^{\ell}e^{C\ln(N)^{1-\alpha}}\Delta_{\#}^{\ell}\leq e^{C\ln(N)^{1-\alpha}}\rho_{n}^{\ell}\left(1+\frac{\ell}{N}\left(\frac{\Delta_{\#}}{\rho_{n,\#}}\right)^{\ell}\right)$
$\leq e^{C\ln(N)^{1-\alpha}}\rho_{n,\#}^{\ell}\left(1+\frac{\ell}{\sqrt{N}}\right)$
where in the second line we used that owing to 4.7 we have $\rho_{n}\geq 1$ and $\rho_{n,r}\geq 1-r$ and we considered $\ell\geq\frac{(1-c^{\prime})\ln(N)}{7K^{\prime}_{\#}\ln(d)}$. Adjusting the constant $C$ and for $N$ large enough we have the wanted result.